A Complete Bibliography of Publications in American Mathematical Monthly: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 May 2024 Version 1.27

Title word cross-reference

$ [465]. (0, n] [1467]. (1 + 1/x)^{x} [582, 801]. (n, m] [1467]. (x^{3} − n)(x^{2} + 3) [586]. 0.(1)(4)(9)(16)(25) [833]. 1 [377, 901, 782]. 1 + 2 + ⋯ + n = n(n + 1)/2 [1451]. 1/2 [377]. 1/n [377]. $109 [434]. $111.00 [409]. $127.00 [519]. $139.00 [465]. 14 [28]. $14.95 [465]. $149.00 [731]. $189.00 [731]. $19.95 [1221]. 2/3 [966]. 23/67 [510]. 24 [379]. $24.00 [1221]. $26.99 [607]. $27.95 [660, 370]. $29.60 [396]. $29.95 [287, 645]. 3 [561, 835, 1247]. $30.00 [810]. $32.00 [344]. 33/97 [510]. $34.95 [420]. 35 [820]. $35.00 [507]. $39.95 [788]. 3x + 1 [1003]. $40.99 [330]. $45 [574]. $47.00 [688]. $49.00 [713]. $49.60 [358]. $49.95 [1192, 382, 450]. $50 [1047, 536]. $50.00 [465]. $52.95 [314]. $53.00 [493, 627]. $58 [555]. $61 [419]. $62 [358]. $65 [148, 302]. $69.95 [479]. $74.00 [554]. 759250125\sqrt{2} [93]. $79.95 [590]. $85.00 [465]. 99 ⋯ 900 ⋯ 0 [611]. A [1299]. AB = C [1299]. ℵ_{0} [864]. ax + by = gcd(a, b) [461]. ax^{2} + by^{2} + cz^{2} + dt^{2} [1303]. B [1299, 1455]. C [1299]. C(Ω) [534]. cos(2π/n) [1060]. cos(x) [1412]. d [513]. e [1387, 19]. e < (1 + \frac{1}{n})^{n+0.5} [46]. e^{1/k} [1071]. E_{2} [200]. ℓ^{2} [122]. ℓ^{p} [55]. f [756]. \frac{p^{3}−p^{2}}{2} [486]. Γ′(1) = −γ [9]. GL_{2}(R_{≥0}) [841]. GL_{n}(C) [915]. H [239]. k [1146, 1253, 1397]. kn − 1 [704]. L [1485, 782]. L^{p} [1295, 950]. L^{p}(R) + L^{q}(R) [390]. L^{r}(R) [390]. L_{2}(R) [567]. L_{p} [81]. ln k [762]. m [858]. N [426]. Q [1394]. R [145, 805, 463]. R^{2} [341]. R^{n} [742, 746]. Z/pZ [458]. Z[x] [60]. N [388, 159, 1279, 425, 832, 1174, 1386, 378, 901, 1264, 16, 29, 720, 238]. n^{th} [795]. n × n [1024]. P [1171, 55]. π [85, 1258, 19]. π^{2} [1391]. R [961]. R^{2} [965]. s [1386]. S_{n} [300]. sin(n) [1218]. sin x [470]. SO(3) [1376]. \sqrt{2} [693, 1056]. \sqrt{2 + \sqrt{2}} [19]. ∑_{k=n}^{∞} k^{−s} [724]. √2 [601, 480]. ℘ [51]. x^{2} + m̄y^{2} = k̄ [458]. x^{3} = x [442]. x^{4} ± y^{4} = iz^{2} [96]. x_{k} ≡ 1(modn) [1355]. ζ(2) = \frac{π^{2}}{6} [473]. ζ(2) = π^{2}/6 [56, 223]. ζ(2k) [773, 895]. ζ(2m) [996]. ζ(2n) [406].

-A4 [820]. -ary [858]. -Cell [379]. -Constructible [513]. -Cube [1264]. -Dimensional [1279, 425, 832, 378]. -Free [55]. -Function [51]. -Functions [782, 1485]. -gon [159]. -in-a-Line [561]. -Orderings [1171]. -Person [388]. -points [239]. -Property [966]. -Ramanujan [1146]. -Resolvability [864]. -Set [28]. -Simplex [1174, 720]. -sums [1386]. -th [1397]. -th-order [1253]. -Tilings [756].

//maa.pinnaclecart.com [507].

0-387-74502-5 [434]. 0-471-24195-4 [409]. 0-8218-3748-6 [344]. 0486492370.html [1221].

1 [507]. 100-year-old [414].

2014 [558]. 2016 [926]. 2018 [1418]. 215pp [450]. 2nd [590].

3 [1192]. 3-Periodic [1403]. 3rd [434].

4th [465].

5th [1136].

978 [1192, 507]. 978-0-262-01935-4 [645]. 978-0-262-19587-4 [302]. 978-0-387-74640-1 [479]. 978-0-465-03292-1 [607]. 978-0-486-49237-7 [1221]. 978-0-521-73794-4 [330]. 978-0-674-72500-3 [788]. 978-0-691-14247-0 [287]. 978-0-691-15270-7 [370]. 978-0-7864-4978-1 [574]. 978-0-8218-3750-4 [465]. 978-0-8218-4316-1 [419]. 978-0-8218-4794-7 [358]. 978-0-8218-5242-2 [396]. 978-0-8218-5368-9 [688]. 978-0-8218-6919-2 [493]. 978-0-8218-6932-1 [554]. 978-0-8218-8321-1 [713]. 978-0-8218-8795-0 [555]. 978-0-8229-4418-8 [660]. 978-0-88385-043-5 [314]. 978-0-88385-574-4 [810].

978-0-88385-581-2 [536]. 978-0-88385-756-4 [465]. 978-0-88385-777-9 [627]. 978-0-914098-91-1 [465]. 978-0-9838837-0-8 [465]. 978-1-107-00202-9 [519]. 978-1-4214-1918-3 [1192]. 978-1-4614-0298-5 [450]. 978-1-4614-5724-4 [731]. 978-1-4614-5725-1 [731]. 978-1-4704-1495-5 [1047]. 978-1-56881-271-7 [148]. 978-1-58488-517-7 [590]. 978-1-6144-610-1 [507]. 978-1442180956 [465]. 978-3-642-19532-7 [420]. 978-3-642-22463-8 [382]. 978-981-4725-72-9 [1221].

A4 [820]. Abacus [1385]. Abelian [1007]. Abscissas [1128]. Absent [1457]. Absent-Minded [1457]. Absolute [225, 1184, 679, 757]. Abused [607, 1497]. Academic [382, 465, 1480, 1488]. Accountant [244]. Accounts [1105]. Achievement [207]. Achieving [1332]. Action [1356, 148]. Actions [1129]. Actually [1125]. Acute [1460]. AdaM [1229]. Add [156]. Addenda [845]. Adding [931]. Addition [976]. Additive [143, 649, 1069, 525, 965]. Adjoint [711]. Admissions [439]. Ado [1097]. Advances [1176]. Adventitious [1008]. Adventures [1385]. Advice [1105]. Afraid [1303]. Africa [1152]. After [946]. Again [559, 585, 297, 480]. AGM [859]. Agnesi [1149]. Ago [1198, 1393, 691, 692]. Agreeable [3]. Agreement [4]. AI [1041]. Aiyar [244]. akademischen [1480]. Alaoglu [552]. Alberto [660]. Alcuin [269]. Alexander [479]. Algebra [1156, 258, 511, 262, 361, 172, 72, 532, 1325, 1136, 362]. Algebraic [143, 631, 722, 361, 43, 745, 336, 280, 1419]. Algebras [1189]. Algorithm [109, 1282, 43, 1431, 903, 461, 857, 192]. Almost [307, 1282, 1410, 1204]. Almost-Near [307]. Along [319]. Alternate [1031, 616, 795, 1010]. Alternating [1424, 1271, 69, 1228, 921, 654, 622, 1244]. Alternative [959, 807]. Alvarez [421]. Álvaro [396]. AM [1100]. America [465, 810, 507, 536, 627, 314]. American [1198, 554, 465, 344, 396, 493, 555, 1047, 1393, 713, 358, 688, 419, 317, 735]. Amity [98]. Among [979]. AMS [358]. Analog [1405, 516, 976]. Analogue [182]. Analogues [487]. Analysis [928, 465, 511, 247, 512, 519, 119, 267, 716, 440, 285, 352, 1421, 1482, 554]. Analytic [1011, 319, 1311]. Analytical [579]. Ancient [349, 31]. Angle [786, 320, 471, 1174, 1008]. Angles [637, 964, 720, 1008]. Angular [393]. Anonymity [1004]. Answer [738, 208]. Anthology [735]. Anticipation [51]. Antiprisms [1444]. Any [771, 597, 586]. Anything [1240]. Apollonian [180]. Appearances [574, 1493]. Appending [597, 157]. Application [1094, 743, 596, 507, 71, 490]. Applications [1394, 1054, 894]. Applied [425, 358]. Apportionment [619]. Approach [320, 1238, 284, 1435, 959, 905, 319, 272, 665, 1227, 1353, 1440, 1482, 1136]. Approval [1199]. Approximate [928]. Approximately [465]. Approximation [474, 583, 855, 798, 1124, 155]. Approximations [68]. Arbelos [516]. Arbitrary [1409]. Arc [650]. Archimedean [1263, 548, 1217, 903]. Archimedes [966, 1232, 132]. Arcology [973]. Arctangent [1423]. Area [832, 1075, 874, 1290, 741, 1183]. Argument [618].

Arguments [1132]. Arising [1373]. Arithmetic [1331, 740, 356, 683, 1410, 1284, 1186, 1381, 704, 1083]. Arithmetical [413]. Arnol’d [131]. Aronszajn [311]. Arrangement [159]. Arrangements [816, 317]. Arrays [67]. Art [117, 1006]. Articles [735]. Artifacts [183]. ary [858]. Arzelà [20]. Ask [1303]. Asked [945]. Associated [1403, 1348, 500]. Association [465, 810, 507, 536, 627, 314]. Associativity [1334]. Asymptotic [582, 60, 68, 801]. Asymptotically [1465]. Asymptotics [855]. Atmospheric [1094]. Aurélien [421]. Author [384, 521]. Automorphism [1439]. Automorphisms [199]. Average [1074]. Averages [8]. Averaging [991]. Avoiding [740]. Award [1080, 1222, 39, 167, 1359, 926, 558, 279, 424, 734]. Awards [1418]. Axiom [707]. Axioms [156, 399].

B [810]. Bacharach [854]. Back [12, 25, 37, 49, 63, 76, 88, 101, 114, 127, 138, 151, 165, 178, 190, 203, 215, 228, 242, 254, 265, 277, 288, 303, 315, 331, 345, 359, 371, 385, 397, 410, 422, 435, 452, 466, 481, 494, 508, 522, 537, 556, 575, 591, 609, 628, 646, 661, 672, 673, 694, 695, 714, 732, 750, 768, 790, 811, 828, 849, 869, 890, 1049, 909, 924, 938, 953, 971, 986, 1000, 1016, 1034, 1195, 1065, 1078, 1091, 1106, 1121, 1137, 1153, 1167, 1179]. Backward [716]. Baggett [536]. Bailey [148]. Balance [1294]. Balanced [1039]. Balancing [425]. Ball [1187, 1257, 1349, 967]. Balls [353]. Baltimore [1192]. Banach [448, 552, 603, 250]. Barbalat [1012]. Barry [1261]. Base [1455, 758]. Base- [1455]. Based [392, 582, 196]. Basel [1272]. Basement [1093]. Bases [1148]. Basic [607, 362]. Basis [567, 718]. Be [388, 1377, 1460, 130, 1469, 1283]. Beatty [348, 930]. Beautiful [541, 251, 1258]. Beauty [1139]. Becomes [1308]. Beginner [1221]. Beginners [247]. Beginning [1133]. Behavior [1127, 34]. Behind [366]. Beloch [181]. Benford [217]. Beppo [797]. Bergmann [382]. Berlin [382]. Bernoulli [677, 1273, 1294]. Bertrand [476]. Berzsenyi [926]. Besicovitch [833]. Best [173, 282]. Best-Fitting [282]. Bets [2]. Better [105]. Betting [978]. Between [1387, 118, 281, 1284, 656, 706, 779, 602, 617, 666]. Beukers [1380]. Beyond [668, 896]. Bézout [43]. Bhargava [1171]. Bicentennial [794]. Bicycle [414]. Big [1409]. Bijection [122]. Bijections [1360]. Billiards [1403]. Binary [683, 1298, 93]. Binet [1028]. Binomial [1245, 447, 1361, 338, 658, 583, 821, 1031, 211, 1099, 460, 814, 1227, 1150, 549, 622, 917, 820, 1266]. Binomial-like [338]. Biology [1178, 671]. Bipartite [45]. Birgit [382]. Birthday [474, 230]. Bisection [833]. Bisectors [471]. Bit [630]. Bite [334]. Bits [1298]. Blichfeldt [239]. Blichfeldt-type [239]. Blind [633]. Blind-friendly [633]. Block [143, 709]. Bluffing [831]. Boca [590]. Bodies [861]. Body [783, 752, 668]. Bold [892]. Bonnet [1088]. Book [287, 382, 969, 1192, 241, 993, 645, 554, 465, 113, 344, 810, 1105, 189, 264, 276, 731, 788, 1033, 952, 907, 396, 87, 887, 214, 866, 519, 867, 749, 1119, 985, 937, 1152, 202, 1014, 493, 574, 1090, 555, 177, 660, 479, 827, 1047, 507, 420, 480, 409, 48, 11, 24, 36, 62, 74, 100, 126, 164, 590, 923, 1136, 148, 421, 330, 253, 999, 536, 713, 358, 302, 434, 627, 767, 847, 607, 149, 370, 314, 688, 137, 227, 450, 419, 479, 827]. Books [607]. Bordered [90]. Borel [530, 729]. Borsuk [428]. Borwein [810, 148, 1018]. Both [981]. Bottazzini [731]. Bound [780, 1164, 364, 1466, 1353, 342]. Boundaries [246]. Boundary [1140, 1094]. Bounded [1432, 20]. Boundedness [198]. Bounds [1029, 724, 1461, 1133, 605]. Bouquet [247]. Box [1392, 195]. Bradley [764]. Bressoud [1222]. Brian [465]. Brief [735]. Bring [1070]. Brocard [786]. Brouwer [596, 1366, 934]. Bruijn [109, 525]. Brunn [1183]. Buck [1385]. Bug [763]. Build [1093]. Built [29]. Bulb [1166]. Bullard [392]. Buoyancy [462]. Buoyancy-Driven [462]. Butterfly [777].

C [507]. CA [465]. Cake [388, 195]. Cake-Cutting [388]. Calcul [465]. Calculation [353]. Calculus [465, 1296, 1133, 1047, 981, 1136, 193, 699, 900, 124, 1499, 153, 465, 465, 507]. Calkin [148, 1443]. Cambridge [645, 788, 519, 330, 302]. Came [470]. Can [1188, 130, 912]. Cancer [670]. Candy [1126]. Cannot [1377, 1381, 1498]. Cantelli [729]. Cantor [1331, 642, 168, 95, 1040]. Cantorvals [856]. Capacitance [490]. Caps [1307, 342]. Card [367, 294]. Cardioid [721]. Cardioids [1374]. Cards [105]. Carlitz [935]. Carlo [330, 65]. Carries [649]. Cars [1044]. Carus [314]. Case [685, 927, 270, 197]. Casual [1125]. Catalan [310, 19]. Catalan-Type [19]. Catch [1302]. Catch-Up [1302]. Catenary [540, 991]. Catenoids [832]. Cates [465]. Catherine [330]. Cauchy [465, 1082, 1197, 54, 284, 1308, 1028, 532, 405, 1015, 824, 804, 1088, 197, 837, 1183, 445]. Caustic [775]. Cavalieri [353]. Cayley [993, 299, 684, 854, 51, 1202]. Cech [552]. Cell [379]. Centennial [697]. Central [583, 821, 622]. Centrosymmetry [745]. Century [1232]. Certain [354, 739, 979, 1239, 501, 744, 500, 58]. Chain [65]. Chains [990, 361, 1144]. Challenging [287, 1484]. Chandler [11]. Change [1045, 1476, 844, 344]. Changing [309]. Chaos [835, 590, 1490]. Chaotic [144, 871]. Chapman [590]. Chapter [258]. Character [285]. Characteristic [991]. Characterization [1394, 394, 1339, 160, 640, 1217, 1310, 1273]. Characterizations [1349]. Characterizing [858, 1069]. Characters [782]. Charles [1080, 1222, 39, 167, 1359, 734, 926, 558, 279, 424]. Charleston [465]. Chase [1278]. Chaundy [392]. Chebyshev [1304, 1114, 1060, 1461, 1388, 273]. Chemistry [1105]. Chernoff [1461]. Chess [897]. Chief [244]. Chinese [485]. Choice [707]. Chord [319]. Chords [1009]. Chow [512]. Christiane [496]. Chromatic [364]. Circle [1209, 572, 964, 319, 682]. Circles [775, 992, 66, 1084, 155]. Circular [4]. Circumsolids [425]. City [396]. Clairaut [551]. Clark [344]. Class [1375, 958, 260, 446, 1228, 1243]. Classes [739]. Classic [419, 1478]. Classical [554, 555, 429]. ClassicalRealAnalysis.com [465]. Classification [707]. Claytor [927]. Close [1070, 1391]. Closed [1082, 1148, 857]. Closure [1405]. Closure-Complement [1405]. Clovers [565]. Cluster [1256]. Codes [1238]. coding [778]. Coefficient [739, 821, 211, 814, 1227].

Coefficients [944, 943, 658, 739, 583, 1259, 1099, 622, 1266]. Cohen [1395]. Coin [1381, 624]. Coincidence [1070]. Coins [298]. Collapse [792]. Collapsed [1377]. Collapsing [274]. Collection [1166, 1448]. Collections [290]. College [358, 1475, 439]. Colmez [607]. Color [524]. Colored [913, 1022]. Coloring [487, 615, 798, 780, 479, 827]. Colorings [634]. Combinatorial [109, 1141, 162, 921, 121, 622]. Combinatorics [325, 649, 1481, 706, 330]. Comments [1370]. Committees [245]. Common [1113, 852, 1099, 1466]. Communication [291]. Commutative [442, 1188]. Commute [1042]. Commuting [472]. Compact [803, 355, 878]. Compact-Open [355]. Compactification [552]. Compactness [813]. Companion [1208]. Comparison [863]. Competition [107, 235, 347, 483, 648, 830, 1002, 1155, 1289, 1429]. Competition* [1055]. Competitions [420, 1500]. Competitive [1302]. Complement [1405]. Complete [246, 45, 399]. Completeness [399]. Complex [511, 731, 448, 501, 1388, 682, 767, 1421, 1495]. Complexity [1164]. Components [746]. Composed [1299]. Composite [1436, 597, 157]. Composition [617, 186]. Compositions [740, 1444]. Compound [1087]. Comprehensive [726]. Computation [499, 773, 1489]. Computations [1381]. Computer [512]. Computing [668, 996, 860]. Concatenation [1434]. Concerning [1271, 824, 158]. Concise [20]. Concurrent [1335]. Conditions [282]. Conductor [490]. Cones [353]. Conewise [943]. Confidence [1432]. Configuration [514]. Configurations [1335]. Confluent [834]. Confocal [123]. Congressional [619]. Congruence [42, 549, 1131]. Congruences [1141, 211]. Congruent [66, 1130, 238]. Conic [916]. Conics [526]. Conjecture [975, 1459, 1341, 1103, 414, 60, 896, 1305, 210, 1430]. Connected [746]. Connection [602]. Connections [1128, 656, 1192]. Connes [123]. Consecutive [1424, 1339, 1434]. Consensus [1432]. Consequence [680]. Constant [957, 175, 341, 1338, 1204, 903, 8]. Constants [983, 389]. Construct [1262]. Constructible [513]. Construction [1224, 822, 1346, 1365]. Constructive [1189]. Contained [390]. Contains [159]. Content [1166, 11]. Continuation [212, 79]. Continued [594, 929, 260, 446, 161, 1258, 155]. Continuity [803, 757, 160, 711, 878]. Continuous [1111, 1409, 122, 803, 642, 1321, 948, 270, 463, 1227, 718, 699, 878]. Contour [604, 1433]. Controversy [1139]. Convergence [175, 679, 1102, 930, 1352, 83, 20]. Convergent [1094]. Converse [623]. Conversion [758]. Converting [246]. Convex [1140, 874, 783, 852, 1009, 1310, 267, 861]. Convexity [919, 982, 1295]. Convolution [921, 622]. Conway [210]. Cook [370]. Copies [783]. Coralie [607]. Core [1176]. Cores [1176]. Corput [719]. Correcting [928]. Correction [853]. Correctional [1237]. Corrections [1250]. Correlation [1024]. Corresponding [1024]. Corrigendum [1045, 962]. Cos [545]. Coset [681]. Cosine [566]. Cosines [684, 655, 1298]. cost [421]. Could [945].

Count [1438]. Countable [341]. Counterexample [355, 209]. Counterexamples [220]. Counting [945, 955, 747, 1177, 1144, 32, 1101, 1266]. Course [465, 1047, 1499]. Courtroom [1316, 1497, 607]. Cover [1170, 409]. Covered [639]. Covering [262, 796]. Coverings [66]. Covers [194]. Coxeter [459]. Cozzens [713]. Cramér [145]. Craps [912]. CRC [590]. Creases [181]. Creative [11]. Creativity [810]. Creek [465]. Criteria [1054]. Criterion [311, 475, 1213, 1081, 129, 1259]. Critical [173]. Crossword [621, 1306, 1315]. Cube [59, 378, 1181, 1264, 1425]. Cube-Free [59]. Cubes [1181]. Cubic [182, 1348, 1374, 656, 913, 1415, 1311, 1333]. Cubics [181, 120]. Cult [660, 1492]. Culture [574, 1480, 1488, 1493, 382]. Cuoco [514]. Curious [1394]. Curvature [957, 1236, 754, 324]. Curve [1082, 918, 17, 997]. Curves [1160, 1140, 560, 874, 1318, 415, 416, 396, 43, 1009, 29, 336, 282, 513, 1421, 1485]. Cutting [388]. cycle [1356]. Cycles [945, 548, 379]. Cyclic [763, 225, 611, 777, 197, 1131, 914, 744]. Cyclotomic [456, 677, 1354, 457]. Cylindrical [425]. Cylindroids [425].

D [148, 1247]. D.C [507]. Darboux [33]. Dark [536, 770]. Data [468, 1307, 282]. David [1222, 507, 434]. Davis [11]. Day [559]. DC [465, 810, 536, 627, 314]. Deal [105]. Decimal [1420]. Decoding [1238]. Decomposition [258, 1367, 1181, 965, 1026, 843]. Decompositions [387, 52]. Deconstructed [465]. Dedekind [7, 1243]. Deficient [1407]. Defined [1102, 1368]. Definition [1238]. Definitions [78]. Degrees [1437]. Dehn [736]. Deleting [818]. Demon [613]. Dennis [465]. Denny [590]. Denominators [1157]. Dense [471, 835, 1239, 1448]. Density [893]. Deranged [705]. Derangement [1323]. Derangements [15]. Derivation [1133, 795, 1010]. Derivations [212]. Derivative [463]. Derivatives [808, 1412, 503, 860]. Derived [1008]. Deriving [651]. Descartes [946, 682]. Descent [319]. Descriptive [1306, 1315]. Design [1041, 1247]. Designs [183]. Determinant [1390, 569]. Determinants [1252]. Determined [1386]. deutschsprachigen [1480]. Devaney [835]. Deviation [605]. Devious [930]. Devlin [574]. Diagrams [913]. Diameter [271, 897]. Diameters [948]. Diatomic [1114, 91]. Dice [1248, 912, 1276]. Dickson [158]. Dierk [420]. Difference [1426, 601, 855, 1127, 528]. Differences [225, 757, 1360]. Different [1044, 977]. Differentiable [1321, 463]. Differential [447, 172, 995, 79, 1273, 706, 1136, 1433]. Differentiate [785]. Differentiation [394]. Digit [1037, 597, 157]. Digits [93]. Dihedral [1174, 720]. Dilated [1042]. Dilog [259]. Dimension [1447, 111, 1280, 455]. Dimensional [685, 1279, 425, 832, 487, 378, 448]. Dimensions [615, 726, 340, 1305, 421]. Diophantine [218, 233, 855, 798, 96, 461]. Direct [625, 171, 532, 638, 489, 1117]. Directly [428]. Dirichlet [1209, 212]. Discontinuous [742, 122, 247, 800]. Discovered [129]. Discoveries [670, 116]. Discovering [85]. Discrete [1241, 911]. Discs [444]. Disks [562].

Dispersive [92]. Dissection [1130]. Dissections [246, 491]. Dissertation [825]. Distance [271, 1084]. Distances [1158]. Distinct [542, 901, 1299]. Distinguished [1080, 1222, 39, 167, 1359, 734, 926, 558, 279, 424]. Distribution [680, 376]. Divergence [930, 1312]. Diverges [1225]. Divide [944, 1190, 16]. Divided [757]. Divisibility [456, 323, 611, 873]. Divisible [1266]. Division [337, 388, 1282, 857]. Divisor [1209, 306, 1113, 1099, 70, 500]. Divisors [296, 814]. Dmitriev [786]. Dmitry [419]. Do [559, 717, 1058]. Doctoral [825]. Dodecahedron [1351]. Does [944, 919, 877, 157, 510]. Domain [1250, 1243]. Domains [577, 531]. Domino [1158]. Doodling [154]. Double [772, 1057, 1380, 473, 820]. Double- [772]. Dover [1221]. Down [759]. Down-Sets [759]. Dr [1080, 1222, 39, 167, 1359, 734, 558, 279, 424]. Dr. [926]. Dramatic [825]. Drawings [45]. Dream [1141]. Drilling [292]. Driven [1178, 462]. Drug [578]. Drug-Induced [578]. Ducci [1319]. Duren [554]. Dynamical [27, 1262, 911, 314, 1482]. Dynamics [407, 667, 871, 1483, 1486, 302]. Dynkin [786].

Early [1171, 831, 617]. Ease [913]. Easy [1164, 725]. eBook [731]. Echelon [643]. ed [434]. Edges [818]. Edited [1198, 810, 574, 420]. Edition [1136, 465, 590, 507]. Editor [524, 383, 451, 520, 608, 689, 789, 888, 970, 1120, 1193, 256, 437, 697, 1018, 1036, 1051, 1401, 1473, 1330, 1328, 75]. editors [382, 420, 664, 1169, 412]. Effective [361]. Efficiently [588]. Eigenproblems [1094]. Eigenvalue [546, 678]. Eigenvalues [781, 1464, 489]. Eigenvectors [489, 1411]. Eighth [1289]. Eisenstein [129, 1142]. Elasticity [426]. Elections [32]. Electric [990]. Electrical [490]. Electromagnetism [932]. Element [1439]. Elementary [377, 823, 958, 361, 724, 40, 355, 1133, 596, 603, 728, 822, 723, 1280, 1227, 713, 198, 1189, 1433, 376, 404, 20, 1496]. Elements [1019, 1268, 1337, 1477, 799, 1119, 409]. Elimination [1373]. Elizabeth [574]. Ellipse [1109]. Ellipses [123, 1403, 208, 1345, 1388]. Elliptic [1403, 1485, 678, 31, 396]. Embedding [1143]. Enable [670]. Encounters [1490, 590]. Encouraging [116]. Encryption [713, 1496]. End [970, 1472]. Endnotes [383, 451, 520, 608, 689, 789, 888, 1120, 1193, 1401, 1473, 1328, 75, 149]. England [645]. Enmity [98]. Entire [728, 1086, 881]. Entropy [1212]. Enumerating [740, 249]. Enumeration [705]. Envelope [305]. Envelopes [775]. Epple [382]. Equal [877, 1181, 510, 1020, 376, 308]. Equal-Tempered [1020]. Equality [808]. Equalization [327]. Equation [601, 338, 458, 1217, 96, 461, 1333, 1445, 1083, 1433, 219, 1294, 1010]. Equations [447, 21, 546, 739, 218, 233, 1464, 1348, 855, 1127, 79, 528, 1273, 1311, 280]. Equiangular [778, 504]. Equicevian [880]. Equidistant [526]. Equidistribution [1362]. Equilateral [1075, 1005, 1130]. Equilibria [1125, 1465]. Equimodular [170]. Equitable [1126]. Equivalent [428, 848]. Erdos [1003, 352, 1445, 1083]. Error [1244]. Errors [928]. Essays [574, 1493]. Essential [124]. Estimating [490]. Étienne [421]. Euclid [1119, 602, 185, 1027, 43, 510, 599, 1369]. Euclidean [1250, 311, 1030, 1143, 1081, 40, 43, 1349, 66, 1431, 1325, 321, 461, 1098]. Euler [1387, 175, 563, 823, 963, 328, 259, 862, 389, 895, 8, 725, 170, 1010]. Europe [1152]. Evaluate [406]. Evaluating [588]. Evaluation [354]. Evan [688]. Evasion [763]. Even [1132, 1309]. Eventually [817, 221]. Every [159, 1204, 902, 196]. Everyone [1208]. Everything [1303]. Everywhere [122, 1321, 1204, 463, 1448]. Evolutes [415]. Evolution [1082, 1112, 455, 65]. Evolutionary [302, 666, 1483, 1486]. Evolutoids [676]. Evolving [676]. Ewing [279]. Example [463]. Examples [818]. Exercise [349]. Exist [282]. Existence [552, 581, 79, 489]. Expanding [817, 221]. Expansion [1390, 823, 587]. Expect [497]. Expectation [838]. Expected [980]. Experimental [555, 1068, 148]. Explained [716]. Explaining [784]. Explicit [747, 965]. exploration [1494, 450]. Exploring [1192, 749]. Exponential [840, 501, 871]. Exponentially [897]. Exponents [134]. Expressions [983]. Extended [643, 583]. Extension [1211, 488, 1450]. Extensions [432, 640, 1446, 1025, 132, 914]. Extraordinary [337, 375]. Extrema [1111]. Extremal [1175]. Extreme [933, 806, 1117, 580]. Extremely [967].

Face [759]. Faces [15]. Facet [15]. Fact [838, 124]. Factor [916, 51]. Factorial [1249, 960]. Factorials [248]. Factoring [942, 1226]. Factorization [387, 1019, 426, 995, 1395]. Factorizations [143, 1039]. Fail [529]. Fair [912, 1276, 892]. Fair-Bold [892]. Fairview [465]. Falling [1053]. Familiar [297]. Family [817, 1030, 1406, 563]. Famous [1309]. Fancy [93]. Fano [1005]. Fantasies [731, 1495]. Far [1187]. Farey [7]. Fast [960]. Faulhaber [807]. Feedback [1285]. Ferguson [806]. Fermat [118, 407, 1131, 579]. Feuerbach [170]. Few [1268, 350, 366]. Fewer [1377]. Feynman [94]. Fiber [948]. Fibonacci [792, 1043, 941, 1342, 616, 1253, 1444, 819, 564, 1385]. Fickett [1341]. Fiction [574, 1493]. Field [1394]. Fields [679, 293, 1217, 872, 273]. Fifteen [142]. Fifth [830]. Filling [140]. Film [574, 1493]. Finding [2, 1019, 624]. Fine [753]. Finite [685, 21, 1190, 1268, 611, 448, 80, 293, 340, 273, 744, 796, 885]. Finite-Dimensional [685]. Finitely [1115]. Finiteness [294, 351]. Finsler [1059]. Firing [1385]. First [1426, 1146, 1156, 465, 129, 729, 172, 95, 334, 1047, 779, 699, 698, 1499, 235]. First-Year [465]. Fit [173]. Fitting [282]. Five [1055, 120, 1130, 793]. Fixed [596, 1366, 1205]. Fixes [1439]. Flag [317]. Flat [1469, 1283]. Flexagons [310]. Flip [974, 1158]. Flipping [1381]. Floor [1042]. Flourishing [1123]. Flowers [1172]. Flows [668]. Foci [1345]. Focus [1432]. Foldable [652]. Footnote [1231]. Ford [1418, 155]. Forecast [380]. Forensic [349]. Foreword [438, 574]. Forgot [599]. Forgotten [617]. Form [643, 683, 567, 581, 1449, 1058]. Formal [1297]. Formation [665]. Forming [245]. Forms [1304, 942, 82, 396, 836, 1136, 1485]. Formula [993, 582, 1219, 541, 488, 747, 543, 1308, 568, 834, 760, 60, 261, 72, 547, 807, 429, 584, 710, 1291, 801, 1309, 1196, 1202, 1206, 976, 1183, 782, 1286]. Formulas [474, 1423, 854]. Found [391]. Foundations [40]. Four [643, 632, 1129, 1230, 1345, 187, 779, 427, 793, 362]. Four-Square [1230]. Fourier [72, 285, 1433, 219]. Fourth [648]. fractals [1490, 590]. Fraction [1258]. Fractional [1008, 441, 8]. Fractions [594, 929, 1219, 901, 161, 155]. Fragmentation [1432]. Francis [192]. Free [55, 59, 548, 822]. Freely [841]. Fresh [1436]. Freshman [1141]. Friend [244]. Friendly [1108, 633]. Friends [915]. Friendship [1023]. Friezes [459]. Frobenius [486, 1101]. Frogs [1108]. Front [1, 13, 26, 38, 50, 64, 77, 89, 102, 115, 128, 139, 152, 166, 179, 191, 204, 216, 229, 243, 255, 266, 278, 289, 304, 316, 332, 346, 360, 372, 386, 398, 411, 423, 436, 453, 467, 482, 495, 509, 523, 538, 557, 576, 592, 610, 629, 647, 662, 663, 674, 675, 696, 715, 733, 751, 769, 791, 812, 829, 850, 870, 1035, 891, 910, 925, 939, 954, 972, 987, 1001, 1017, 1180, 1050, 1066, 1079, 1092, 1107, 1122, 1138, 1154, 1168]. Frontier [671]. FTC [588]. Fuchs [419]. Fueter [1040]. Function [740, 1459, 81, 582, 295, 488, 757, 731, 747, 517, 259, 1163, 340, 34, 1469, 1204, 1132, 892, 184, 1291, 902, 1309, 51, 463, 1283, 1353, 500, 701, 997, 1495]. Functional [21, 1217, 79]. Functionals [1392, 111]. Functions [1111, 816, 956, 1128, 377, 563, 1409, 1021, 958, 803, 341, 650, 247, 1321, 259, 995, 1042, 555, 1310, 728, 1210, 1420, 1086, 881, 718, 70, 699, 878, 782, 134, 1485, 396]. Functor [711]. Fundamental [1156, 361, 72, 532, 1325, 699, 362, 153]. Funny [836]. Furstenberg [602]. Further [1221, 1200, 1142]. Future [1004].

G [434]. Gaetana [1149]. Galileo [319]. Gallery [1006]. Gallian [167]. Galois [443, 640]. Gambling [892]. Game [1125, 367, 1041, 792, 512, 438, 294, 912, 364, 497]. Games [443, 485, 581, 574, 527, 302, 1483, 1486, 1493]. Gamma [1249, 488, 517, 110, 1291]. Gap [1162]. Gaps [1103, 1231]. Gardner [561]. Gareth [1192]. Gauss [1156, 1209, 339, 248, 544, 623, 1364, 1131]. Gaussian [505, 96, 324]. Gelfand [738]. Gem [786, 339]. Gems [632, 550, 898]. General [863]. Generalization [515, 1279, 391, 502, 162, 299, 603, 895, 161, 725, 935, 1320]. Generalizations [772, 540, 947, 1342, 917]. Generalize [528]. Generalized [106, 563, 1114, 1071, 1277, 1028, 1070, 267, 389, 881, 526, 579, 1252, 1458, 539]. Generalizing [339, 760]. Generate [775, 597, 157, 841]. Generated [335, 1253, 1115]. Generating [740, 259, 997]. Generating-Function [740]. Generator [1098]. Generic [282]. Genesis [1171]. Genetic [667]. Geodesic [879]. Geometria [1214]. Geometric [1304, 594, 1116, 731, 1284, 146, 913, 761, 261, 1376, 444, 727, 737, 797, 618, 1495]. Geometrical [786]. Geometries [40]. Geometry [721, 630, 1316, 290, 454, 1475, 777, 706, 52, 1291, 280, 78, 767, 358, 767]. George [1247, 926]. German [382, 1480, 1488]. German-Speaking [382, 1488, 1480]. Gerrymandering [1317, 1416]. Gersgorin [444]. get [1497]. Ghys [421]. Gian [330]. Gian-Carlo [330]. Gin [1080, 1222, 39, 167, 1359, 734, 926, 558, 279, 424]. Gini [14, 374]. Girgensohn [148]. Given [944, 1019, 1290, 1435, 134]. Glasser [391]. GM [1100]. Gödel [1119]. Goldbach [563, 60, 133]. Golden [206, 95, 1437]. Golomb [1103]. gon [159]. gons [29]. Good [2, 350]. Goose [1278]. Gordian [1052]. Grade [1095, 104]. Gradient [1465]. GrahaM [1229]. Graph [1052, 875, 1144, 57, 1175, 1378]. Graphs [816, 681, 1129, 893, 1437, 174, 45, 1158, 52, 1175]. Grasshopper [141]. Gray [731]. Greater [913]. Greatest [1113, 1099]. Greedy [443]. Green [894]. Gronwall [1255]. Gronwall-type [1255]. Group [1093, 1337, 199, 649, 1356, 1435, 1439, 1243, 57, 736, 1405]. Group-Theoretic [1405]. Groups [21, 681, 59, 250, 822, 486, 899, 744]. Grows [897]. Growth [217]. Guess [298]. Guessing [527]. Guest [664, 1169, 412]. Guide [1221, 465]. Guided [767]. Guinand [170]. Gulick [590]. Gung [1080, 1222, 39, 167, 1359, 734, 558, 279, 424, 926]. Gurvits [18].

H [555, 148, 330, 302]. Hackensack [1221]. Hadamard [569, 1366]. Hadwiger [1059]. Hahn [1367]. Hair [496]. Hairy [967, 1257]. Half [365]. Half-integers [365]. Hall [590, 1061, 1449]. Hall/CRC [590]. Halmos [1418]. Halmos-Lester [1418]. Halving [1461]. Hamilton [945, 299, 617]. Hamiltonian [548, 1404, 379]. Hanging [872]. Hangover [270]. Hanoi [57]. Happens [1282]. hardcover [731]. Hardest [1389, 1400]. Hardy [958, 1241, 1196]. Harmonic [1128, 1021, 69, 1228, 1370, 1242, 1312, 1225]. Harmony [731, 1332, 1495]. Hartogs [84]. Harvard [788]. Hasse [220]. Hatchet [414]. Hating [287, 1484]. Hausdorffization [657, 1062]. Hawkins [424]. Heads [633]. Heidelberg [382, 420]. Helen [810]. Helge [918]. Help [1340]. Henle [627]. Henry [1247]. Hermite [53, 567]. Heron [539, 267]. Heronian [402]. Hersh [287]. Heuristic [618]. Hex [533]. Hexagons [977]. Hidden [1495, 731]. High [438, 503]. High-Order [503]. Higher [487, 928, 835]. Higher-Dimensional [487]. Highly [1436]. Hilbert [1419, 1265]. Hilbert’s [404]. Histograms [257]. Historical [153, 701]. History [735, 788, 813, 484]. Hitting [237]. Hoboken [409]. Hölder [804]. Holditch [1109]. Homomorphism [283]. Homomorphisms [534, 1144, 1175]. Homothetic [783]. Homotopy [1062]. Hopf [1267]. Hopkins [1192]. Horseshoe [974]. Hot [1458]. Houston [465]. http [507]. Hu [1080, 1222, 39, 167, 1359, 734, 926, 558, 279, 424]. Hua [618]. Hull [783]. Human [1123]. Humans [749]. Hundred [1166]. Hyper [772]. Hyper- [772]. Hyperbola [1452]. Hyperbolas [743]. Hyperbolic [777, 976]. Hypercube [980]. Hyperplane [378]. Hypotenuses [1339]. Hypothesis [1261].

IAS [396]. IAS/Park [396]. Ice [1307]. Ideal [1250]. Idempotent [1268]. Identically [1469, 1283]. Identities [1245, 1361, 368, 6, 703, 53, 470, 723, 156]. Identity [447, 391, 310, 162, 392, 529, 251, 131, 921, 1025, 1322, 460, 935, 917, 618, 619, 885, 182, 211]. If [1190, 1292, 559]. Illuminated [209]. Illuminating [430]. Illustrate [468]. Image [1067]. Images [642]. Imaginary [200]. Imaging [470]. Implies [835, 428]. Imply [919]. Importance [1144]. Impossibility [1317]. Improved [1241]. Improving [175]. Income [374]. Indefinite [82]. Independence [108]. Independent [284, 1159, 340, 194, 1175]. Index [150, 384, 521, 14, 374]. Indian [349]. Induced [1404, 1086, 578]. Inducing [159]. Inductive [533]. Industry [217]. Inequalities [1304, 55, 30, 1461, 517, 1446, 272, 804, 500]. Inequality [840, 1271, 14, 1284, 1308, 374, 1241, 978, 46, 1437, 862, 569, 1059, 1233, 1015, 824, 797, 352, 1100, 1255, 784, 1218, 837, 404, 445]. Infinite [1334, 658, 1368, 1071, 994, 328, 1087, 762, 491, 1147, 736, 1173, 19]. Infinitely [747, 1186, 1262, 704, 1073]. Infinitésimal [465]. Infinitesimals [785, 472, 519]. Infinitude [602, 185, 1134, 776, 1297]. Infinity [130, 34]. Information [1477, 409]. Inscribed [123, 874]. Inserting [931]. Insights [1453]. Instead [913]. Institute [1237, 1178]. Integer [1387, 944, 1424, 940, 943, 1435, 586, 1259, 1468, 941, 157, 335, 1431, 1444, 1132, 1147]. Integer-Valued [940]. Integers [143, 225, 916, 884, 1386, 543, 157, 261, 1434, 96, 1354, 1309, 793, 365]. Integrability [757, 699, 134]. Integrable [34]. Integral [391, 1045, 153, 1211, 1412, 842, 180, 1433, 844, 83, 531, 465]. Integrals [958, 1380, 1392, 588, 473, 860]. Integration [894, 604, 1133, 705, 68]. Interesting [1406, 400, 1224, 32]. Intermediate [1453]. Interpolation [884, 1219, 960]. Interpretation [261, 1376]. Intersecting [416, 43]. Intersection [681]. Intersections [1349]. Interval [817, 82, 221, 1204, 902, 1270]. Intervals [878]. Introducing [1125]. Introduction [248, 493, 713, 280, 1496, 713]. Invariance [1108, 82]. Invariant [257]. Inverse [122, 917, 135]. Inversion [72, 1433]. inversive [1374]. Invitation [420, 1500, 871, 554]. Involutes [106, 415]. Involving [563, 658, 1218]. Irrational [964, 85]. Irrationality [601, 95, 949, 480, 693, 1056, 58]. Irreducibility [475, 1054, 1259, 1468, 1265, 457]. Irreducible [426, 120]. Irreducibles [1081]. Isaacs [358]. ISBN [287, 382, 1192, 645, 554, 465, 344, 810, 1221, 731, 788, 396, 519, 493, 574, 555, 660, 479, 1047, 507, 420, 409, 590, 148, 330, 536, 713, 358, 302, 434, 627, 607, 370, 314, 688, 450, 419]. ISBN-13 [420, 419]. Island [1047]. Isolated [864, 1254]. Isometrically [1143]. Isometries [900]. Isoptic [1009]. Isosynchronous [815]. Iterated [456, 997]. Iterations [903]. Itself [1351]. ix [344].

J [148, 713, 370]. J. [635]. Jacobi [131, 268, 427, 885]. Jacobsthal [182]. James [734]. Jan [11]. Jefferson [574]. Jensen [978]. Jeremy [731]. Jessica [574]. Jewish [382, 1480, 1488]. Joan [558]. Johannes [1149]. John [287, 409, 279]. John-Steiner [287]. Johns [1192]. Johnson [1450].

Jonathan [1018]. Jordan [934]. Jos [421]. Joseph [167, 330]. Joukowsky [1333]. Joy [1235, 409]. Jüdische [1480]. Julia [117]. Julius [117]. Jumble [1340]. Jumbles [1389, 1400]. Jump [1041]. Jumps [141]. Justin [571].

Kantorovich [928]. Keeler [542]. Keith [574]. Kenneth [39]. Kenschaft [344]. Kepler [1149, 1087, 393]. Kernel [995]. Khayyam [913]. Killer [373]. Kin [351]. Kinoshita [1378]. Kirkman [142]. Klarner [1003]. Klein [1129]. Knar [1291]. Knot [615]. Knotoids [1402]. Knots [1172, 1402, 1052]. Knotted [1378]. Know [559, 940, 1292, 196, 1303]. Knowledge [499]. Koch [918, 29, 997]. Kontsevich [416]. Korovkin [1392]. Kot [1047]. Kristopher [450]. Kronecker [855, 709]. Krull [1280]. Krylov [928]. Kultur [1480]. Kung [330]. Kuratowski [1405, 28].

L [391, 396]. L-functions [396]. Labeled [1070]. Lackmann [420]. Lagrange [1230, 161, 619]. Languages [1438, 1297]. Laplace [1390, 6, 613]. Laplacian [933]. Large [1410, 978, 605, 784]. Large-Deviation [605]. Larry [536]. Last [334, 1445]. Lattice [1277, 652, 402, 1022]. Lattices [1162]. Laurent [950]. Law [963, 545, 978, 655, 784, 169, 727, 737, 217, 1462]. Lawrence [507]. LCM [873]. LCM-Sequences [873]. Lead [310]. Leading [601]. Learn [749]. Learners [913]. Learning [1285]. Least [1466, 238]. Lebesgue [947, 588, 1270, 843]. Lectures [419, 1478]. Left [249]. Legendre [354, 1247, 169]. LEGO [955]. Lehmer [400]. Leibniz [515]. Leila [607]. Leitzel [558]. Lemma [1160, 929, 947, 729, 1027, 861, 719, 117, 1012, 428, 1369]. Lemniscate [903]. Length [560, 650, 683, 1270]. Lengths [1140, 1019, 778, 1318, 1038, 426]. Lens [539]. Leonid [18]. Less [877, 1293]. Lester [1418]. Letter [664, 256, 437, 524, 697, 1018, 1036, 1051, 1330, 412]. Letters [913]. Level [1041, 1318, 562]. Levi [797]. Lewis [734]. Lexell [1096]. Leys [421]. Library [396, 1047]. Lie [568]. Lies [230]. life [1484, 287]. Light [1166, 775]. Lights [1185]. Like [1173, 367, 338]. Liljedahl [810]. Lill [181]. Limit [295, 863, 291, 44]. Limiting [1127]. Limits [645, 1392, 501, 981, 193, 1489, 1498]. Limsup [1447]. Lindemann [1324]. Lindenstrauss [1450]. Line [118, 337, 561, 580, 1239, 776, 1216, 1448]. Line-of-Sight [580]. Linear [817, 258, 636, 612, 108, 943, 511, 262, 1419, 1127, 982, 111, 1479, 1136, 461, 765, 434]. Linear-algebraic [1419]. Linearly [1159, 340]. Lines [159, 1277, 16, 173, 1335]. Link [1387]. Links [52, 666]. Liouville [320, 958, 881]. List [399]. Little [1097, 407, 1131]. Littlewood [635]. Lives [1415]. Ljunggren [549]. Loaded [912]. Local [1111]. Locus [31]. Logarithmic [1422, 908]. logic [1498, 1221, 1221]. Logical [1166]. London [645]. Lonely [1008]. Long [1036]. Longest [335]. Look [957, 539, 1436, 172]. Looking [350]. Looks [417]. Looping [893]. Lorentz [1214, 1251]. Lost [305, 209]. Lottery [2]. Lover [1208]. Loving [1484, 287]. Lowell [107, 235, 347, 483, 648, 830, 1002, 1155, 1289, 1429]. Lower [1164, 1466]. Lozano [396]. Lozano-Robledo [396]. Lucas [1342, 616, 819, 623, 564]. Luenberger [434]. Luke [148].

M [465, 1018, 1221, 409, 148]. M. [391]. MA [645, 148, 302]. MAA [465, 697, 698]. maa.pinnaclecart.com [507]. Macaulay [587]. Machine [462]. Made [361]. Magic [1108, 363]. Mahler [755, 1305]. Mail [305]. Majorization [539]. Majorization-Minimization [539]. Make [334, 848]. Makes [1302]. Malte [420]. Mamikon [393]. Mancala [485]. Many [874, 1188, 281, 1186, 1262, 704, 1415, 1073, 498]. Map [487, 1333, 871]. Mapping [603]. Mappings [321, 800, 273]. Maps [817, 636, 742, 943, 221, 271, 948, 1325, 1086, 324]. Marden [625, 1116]. Margaret [713]. Maria [1149]. Marjorie [11]. Mark [1047]. Market [468, 1229]. Markov [990, 1461, 1144, 65]. Marriage [1108, 708, 439, 1061, 1449]. Martha [1080]. Martin [358, 561]. Martínez [660]. Mascheroni [175]. Mastering [117]. Masterpieces [1208]. Matchings [705]. Matchstick [174]. Math [660, 1497, 1068, 1492, 607]. math.org [421]. Mathemagical [187]. Mathematical [287, 1166, 1198, 244, 554, 735, 465, 344, 810, 1221, 247, 1112, 1478, 1055, 396, 670, 107, 235, 347, 483, 648, 830, 1002, 1155, 493, 1289, 1429, 555, 479, 827, 1047, 507, 1393, 665, 536, 713, 358, 627, 669, 314, 688, 450, 419, 1484, 421, 1494, 1178, 419]. Mathematically [749]. Mathematician [510, 1292, 196, 536]. Mathematicians [382, 1480, 555, 1488, 810]. Mathematics [287, 1192, 349, 1077, 974, 1222, 1208, 788, 1119, 670, 374, 1152, 420, 825, 11, 148, 927, 1359, 734, 926, 1285, 558, 770, 713, 1500, 1493, 1123, 1247, 484, 607, 154, 1489, 1496, 1478, 1484, 1476, 1498, 1080, 39, 167, 279, 424, 344, 419, 749, 574]. Mathematiker [1480]. Matrices [636, 1319, 702, 1024, 781, 208, 915, 745, 596, 1431, 1086, 841, 1299, 1226, 1411, 1324, 268]. Matrix [1350, 1334, 338, 1446, 838, 995, 5, 709, 1449, 1131, 1189]. Matrix-norm [1446]. Matter [12, 25, 37, 49, 63, 76, 88, 101, 127, 1, 13, 26, 38, 50, 64, 77, 89, 102, 115, 138, 151, 165, 178, 190, 203, 215, 228, 242, 254, 128, 139, 152, 166, 179, 191, 204, 216, 229, 243, 265, 277, 288, 303, 315, 331, 345, 359, 371, 385, 255, 266, 278, 289, 304, 316, 332, 346, 360, 372, 397, 410, 422, 435, 452, 466, 481, 494, 508, 522, 386, 398, 411, 423, 436, 453, 467, 482, 495, 509, 537, 556, 575, 591, 609, 628, 646, 661, 672]. Matter [673, 694, 695, 523, 538, 557, 576, 592, 610, 629, 647, 662, 663, 674, 675, 714, 732, 750, 768, 790, 811, 828, 849, 869, 890, 696, 715, 733, 751, 769, 791, 812, 829, 850, 870, 1049, 909, 924, 938, 953, 971, 986, 1000, 1016, 1034, 1035, 891, 910, 925, 939, 954, 972, 987, 1001, 1017, 1195, 1065, 1078, 1091, 1106, 1121, 1137, 1153, 1167, 1179, 1180, 1050, 1066, 1079, 1092, 1107, 1122, 1138, 1154, 1168, 770, 114]. Matters [1367, 291]. Max [736]. Maxim [416]. Maximal [1075, 1337, 1009, 784, 799, 194]. Maximality [949]. Maximally [1248]. Maximum [874, 650, 503, 431]. May [388]. Mazur [1261]. McDougall [572]. McFarland [574]. MD [1192]. Meal [334]. Mean [877, 612, 1128, 341, 1422, 570, 933, 1088, 908, 836]. Means [284, 350, 1284, 267]. Measurability [232]. Measurable [961, 44, 134]. Measure [80, 493, 717, 1173, 1270]. Measures [14, 477, 848]. Measuring [1416]. media [1493, 574]. Medial [234]. Median [282]. Medical [470]. Medieval [1152]. Medium [614]. Meets [1149, 43, 393]. members [358]. Memoir [536]. Memorable [1022]. Memory [497]. Menelau [1174]. Menelaus [1417]. Menger [1143]. Mensa [1237]. Mercator [281]. Meromorphic [212]. Mersenne [614]. Metagrobologists [1166]. Method [1145, 929, 1232, 455, 1363, 1305, 392]. Methods [928, 1088, 65]. Metric [499, 1143, 803, 864, 756]. Michael [465, 627]. Microorganisms [668]. Middle [814]. Midpoints [1128]. Miller [713]. Milnor [929]. Minded [1457]. Mineola [1221]. Minimal [818, 305, 1060]. Minimalist [1462]. Minimax [222]. Minimization [539]. Minimizing [741]. Minimum [561]. Minkowski [1164, 1183]. Minorities [344, 1476]. minutes [421]. Miranda [1366]. Mirrors [324]. Mishandling [1027]. Missed [392]. Missing [305]. Mitscherlich [282]. Mixed [816, 808]. Mixing [1173, 498]. Mixing-Like [1173]. Mocposite [956]. Model [348, 380, 1251]. Modeling [507]. Models [418]. Modern [519, 1090]. Modified [806]. Modular [396, 743, 960, 70, 1485]. Modules [387]. Moduli [1251]. Modulo [858, 586, 120, 238]. Modulus [146, 431]. Moessner [401]. Moivre [1206]. Moll [555, 148]. Mom [1105]. Moments [1245, 779]. Momentum [393]. Money [309]. Money-Changing [309]. Mongolian [1022]. Monic [60]. Monochromatic [300]. Monographs [314]. Monoids [143, 611]. Monotone [210, 83]. Monotonic [650, 463]. Monotonicity [135]. Monsters [1296]. Monte [65]. Month [1198, 1393]. Monthly [690, 889, 1048, 1194, 716, 1068, 550, 898, 692, 1329, 1198, 1474, 735, 1249, 1393, 691, 1320]. Moore [507, 1145, 571]. Moritz [382]. Morley [761]. Morrie [727, 737]. Moser [945, 187, 1083]. Most [251, 334, 1161]. Mostly [1487, 688]. Mothers [1105]. Motif [218]. Motion [97, 1379, 462]. Motivated [1272, 899]. Moving [15]. Multi [1041]. Multi-Jump [1041]. Multichoose [921]. Multiconnected [932]. Multidimensional [353]. Multinomial [904]. Multiple [1070, 1466]. Multiplication [1334, 1148]. Multiplicities [1019, 444]. Multiplier [1215]. Multiplying [1226]. Multiset [1386]. Multistep [1023]. Music [1192]. Must [1469, 1283]. My [1201]. Mycielski [1346]. Mysteries [1007]. Mysterious [784]. Myths [287, 1484, 1492, 660].

N [148, 159]. Nader [519]. Nadis [788]. Nakajima [861]. Napoleon [853, 700, 322, 1011, 1335]. Napoleon-style [1335]. Napoleonic [1382]. Narayana [244]. Nash [1125]. Natural [171]. Nature [291]. NC [574]. Near [307, 130, 22]. Nearest [335]. Nearly [1181, 617]. Nearly-Forgotten [617]. Needed [193]. Neighbors [1052]. Nemirovski [30]. Nested [528, 859]. Nets [1444]. Network [430, 990]. Networks [667]. Neumann [222, 633]. Neuroscience [669]. Never [718]. Newman [1082]. Newton [929, 368, 884, 6, 1348]. NFL [1037]. Nillsen [314]. Nilpotent [1268]. Ninth [1429]. Nitecki [465]. NJ [287, 1221, 409, 370]. No [388, 561, 358, 421]. No- [561]. Nodes [430]. Noether [1189]. Noise [1404]. Noise-Induced [1404]. Non [1030, 448, 543, 40, 872, 1007]. Non-Abelian [1007]. Non-Euclidean [1030, 40]. Non-recursive [543]. Non-Supercyclicity [448]. Non-Uniform [872]. Nonaxis [964]. Noncommuting [472]. Nondefinability [78]. Nonexistence [501]. Noninductive [1206]. Noninteger [156]. Nonlinear [1479, 434]. Nonnegative [1411]. Nonopenness [1111]. Nonresidues [1072]. Nonreversible [990]. Nonsymmetric [678]. Nontransitive [1248]. Nontrivial [1355, 1243]. Nonuniqueness [1281]. Nonvanishing [782]. norm [1446]. Normal [559, 581]. Normality [1213, 833]. Norms [81, 340, 111, 965]. North [1152]. Noson [645]. Note [685, 367, 1021, 1452, 1368, 1380, 1267, 1354, 1074, 1364, 908, 445]. Notes [970, 1379]. Nothing [504]. Notion [799]. Notions [1173]. Novel [1056]. Nowhere [1321]. Null [856]. Nullstellensatz [121]. Number [858, 1387, 325, 143, 310, 1190, 1375, 1337, 1377, 298, 16, 1090, 336, 417, 156, 364, 639, 795, 899, 980, 744, 314]. Numbers [759, 1426, 1261, 499, 262, 963, 1436, 611, 1043, 805, 597, 1407, 119, 941, 978, 921, 1444, 638, 677, 441, 158, 351, 1336, 1083, 784, 920, 627, 1223, 796, 1491, 1497, 555]. Numerical [983, 1203, 1133, 838, 1177, 677]. NY [1221].

Observation [416]. Observations [1453, 614, 824]. Obtain [93]. Occasion [794]. Occurrence [979]. Occurs [297]. Oceanic [1094]. Octahedron [17]. Odd [1268, 59, 639, 158]. ODEs [894]. Often [747]. Old [1160, 258, 651, 40, 915, 417, 526, 414]. Olympiad [1022]. Omar [913]. Omnibus [419, 1478]. Once [876, 585, 480]. Oncology [670]. One [1166, 1190, 580, 208, 1163, 1090, 693, 776, 455, 238, 782, 1252]. One-Formula [782]. One-Line [776]. One-Sentence [580]. One-to-One [1163]. Online [1125]. onto [742, 122]. Open [742, 355, 1130]. Operations [362]. Operator [1446, 995, 22, 135]. Operator-valued [1446]. Operators [22, 678]. Opinions [1432]. Optimal [1029]. Optimality [1215]. Optimally [911]. Optimizing [1041]. Orbit [1026]. Orbits [1403, 911]. Order [1268, 1337, 59, 486, 503, 1253]. Ordered [679, 1217]. Ordering [707]. Orderings [144, 1171]. Origami [52]. Origin [455]. Original [175]. Origins [736]. Orthocenters [131]. Other [1145, 1410, 15, 915, 574, 186, 1493]. Out-of-Sight [536]. Outer [645, 1270, 1498]. Ownership [868].

P [330, 340]. PA [660]. Packing [195]. Packings [180]. Page [993]. Pair [123, 912]. Pairs [348, 841]. Pairwise [66]. Paleo [349]. Paleo-Mathematics [349]. Paper [454]. paperback [1221]. Parabolas [966]. Parabolic [516]. Paradox [1023, 1373]. Paradoxical [1200]. Parallelograms [123, 326]. Parameterizations [280]. Parametrization [336]. Parbelos [516, 600]. Parent [564]. Parents [427]. Parity [45, 236, 795]. Park [396]. Parking [816, 1044]. Part [1205]. Partial [1219, 808, 901, 1397, 51]. Partition [582, 875, 1322, 1196]. Partitional [930].

Partitioning [746, 1448]. Partitions [858, 1387, 1339, 1132, 1254, 1176]. Partnership [1196]. Parts [740, 1339, 894, 441, 8]. Pascal [915, 261, 5, 211, 1210, 513]. Passengers [1457]. Passing [1140, 1385]. Pasta [1247]. Path [1196, 784]. Paths [815, 319]. Patricia [344]. Pattern [665]. Paul [1418]. Payments [716]. Peano [530, 1421]. Pedagogical [813]. Pedals [205]. Peer [1285]. Pennsylvania [927]. Pentagonal [920]. People [1125]. Perfect [499, 388, 979, 1407, 1434, 158, 351, 490, 326]. Perimeter [1435]. Period [835]. Periodic [320, 943, 835, 260, 1403, 528, 800]. Periodicities [469]. Periodicity [577]. Perish [465]. Permanents [18]. Permutation [300, 236]. Permutations [1131]. Perpetual [462]. Perron [678]. Perron-type [678]. Person [388]. Personal [1105]. Perspectiva [1214]. Perspectives [767]. Perturbations [982]. Perturbed [1404]. Peter [554, 810]. Peters [148]. Ph.D [927]. Phasor [976]. Phi [1293]. Philip [1359]. Philosophies [1077]. Physical [983, 1015]. Physics [291]. Pi [559, 845, 735, 469, 1423, 1293, 903, 559]. Pick [771, 95]. Pictorial [900]. PIDs [1030]. Piecewise [817]. Pigeonholes [620]. Pittsburgh [660]. Places [15, 915]. Plan [334]. Planar [1140, 893]. Plane [943, 1005, 1347, 40, 43, 16, 634, 66, 1388, 977]. Planes [1349]. Planimeters [414]. Plank [1124]. Play [1125, 912, 1229]. Player [1041]. Playing [364]. Please [1201]. Plus [931, 1377, 1090]. Poincaré [729, 1366, 882]. Poinsot [1087]. Point [864, 1277, 596, 555, 336, 1366, 706, 1272, 1311, 1205]. Points [1140, 1256, 1111, 1281, 835, 880, 964, 173, 1070, 800, 639, 44, 239]. Poisson [1010]. Poker [831]. Polar [1307]. Political [1416]. Politics [927]. Polya [883, 41, 1040, 653, 327]. Pólya’s [598]. Polygonal [1182]. Polygons [700, 1075, 778, 1347, 292, 754, 356, 852, 459, 652, 280, 504, 853]. Polyhedra [42]. Polyhedral [741]. Polyhedron [1087]. Polynomial [944, 546, 312, 884, 407, 97, 543, 1464, 979, 1060, 130, 173, 1215, 1115, 1280, 133, 186, 564]. Polynomials [1304, 1029, 1323, 456, 1406, 1054, 940, 1219, 1114, 755, 323, 1113, 208, 1060, 293, 1374, 586, 1259, 1468, 941, 567, 146, 60, 1086, 677, 1040, 186, 170, 1362, 457, 531]. Polytopes [209]. Pompeiu [1279]. Poncelet [794]. Ponomarenko [1198]. Pool [978]. Popcorn [1163]. Popular [574, 1493]. Population [1483, 1486, 302]. Porism [777]. Positive [636, 157, 441, 1309]. Positivity [1459]. Possible [344, 1187, 1476]. Possibly [545]. Postulate [476]. Power [1424, 306, 337, 1094, 1157, 211, 981, 566, 272, 1272, 1397, 718, 1266]. Power-Sum [1157]. Powers [1426, 884, 658, 543, 1071, 745, 261, 1253, 641, 1161, 1286]. pp [287, 382, 645, 554, 465, 344, 810, 1221, 731, 788, 396, 519, 493, 574, 555, 660, 479, 1047, 420, 409, 590, 148, 330, 536, 713, 358, 302, 434, 627, 607, 370, 314, 688, 419]. Predicting [1004]. Prediction [468]. Prefer [845, 735]. Preimages [1203]. près [51]. Press [287, 1192, 645, 465, 788, 519, 660, 590, 330, 302, 370]. Price [1192, 507]. Priest [438]. Primality [717]. Primary [1083]. Prime [1426, 1146, 296, 1103, 1338, 747, 723, 1312, 725, 238, 1098, 1266, 1261]. Prime-Representing [1338]. Primes [839, 602, 469, 1190, 185, 658, 1410, 505, 1186, 1262, 1467, 704, 1073, 1134, 776, 599, 1299, 1297, 1142]. Primitive [1339, 1364]. Primitivity [1026]. Princeton [287, 370]. Principal [1250, 678]. Principle [307, 146, 503, 490, 431, 220, 353]. Principles [425]. Printing [1247]. Prisoner [1372]. Prisoners [1166]. Probabilistic [325, 162, 904, 460, 917, 1305, 1451]. Probabilistically [223]. Probability [468, 469, 1439, 765, 327]. Problem [1057, 474, 763, 487, 1323, 1209, 320, 42, 309, 875, 561, 955, 1278, 305, 1224, 780, 426, 1008, 1124, 1003, 905, 1070, 752, 267, 224, 1022, 308, 600, 579, 209, 1320, 539, 142, 141, 1272]. Problems [112, 136, 147, 163, 176, 188, 201, 213, 226, 240, 252, 263, 275, 286, 301, 313, 329, 343, 357, 369, 381, 384, 395, 408, 433, 449, 464, 478, 492, 506, 518, 521, 535, 553, 573, 589, 606, 626, 644, 659, 686, 687, 712, 730, 748, 766, 787, 809, 826, 846, 865, 886, 1046, 906, 922, 936, 951, 968, 984, 998, 1013, 1032, 1166, 1191, 1063, 1076, 1089, 1104, 1118, 1135, 1151, 1165, 1220, 1094, 413, 10, 35, 47, 61, 73, 86, 99, 1207, 1234, 1246, 1260, 1274, 1287, 1300, 1313, 1326]. Problems [1343, 1357, 1371, 1383, 1398, 1413, 1427, 1441, 1454, 1470, 23, 125, 373, 498, 1068]. Product [823, 568, 821, 328, 565, 959, 1199, 822, 68, 1147, 859, 885, 1440]. Products [598, 368, 542, 1071, 979, 879, 110, 19]. Professor [1105]. Professors [1105]. Program [484]. Programming [1479, 434]. Programs [337]. Progression [683, 704]. Progressions [1410, 1186, 1083]. Projectile [1379]. Projection [614]. Projections [281]. Projective [767]. Proof [1160, 1045, 1269, 625, 9, 1082, 764, 802, 1156, 222, 1458, 432, 530, 1396, 1116, 631, 722, 171, 145, 572, 162, 392, 823, 1257, 864, 407, 808, 545, 361, 1367, 580, 448, 805, 1284, 1419, 1011, 794, 774, 568, 1028, 328, 533, 838, 1031, 904, 761, 581, 95, 616, 862, 18, 532, 603, 547, 1059, 921, 946, 84, 638, 56, 551, 967, 476, 1450, 1073, 1134, 121, 655, 727, 737, 1395, 1015, 429, 33, 1280, 584, 693, 161, 776, 460, 1088, 1430, 1309, 1196, 473, 797, 352, 477, 1100, 489, 431, 1202, 1150, 1206, 1270, 549, 198]. Proof [1101, 622, 699, 1189, 844, 799, 806, 878, 1117, 843, 837, 1451, 600, 404, 782, 1462, 1265, 124, 882, 885, 20, 719, 1286]. Proofs [602, 46, 211, 1366, 981, 1378, 1295, 917, 457, 58]. Properties [232, 728]. Property [1263, 540, 991, 966, 133, 899, 600]. Proposed [384, 521]. Prove [1141, 1346, 1056]. Proved [129]. Providence [554, 465, 344, 396, 493, 555, 1047, 713, 358, 688, 419]. Proving [552, 85, 491, 223]. Prüfer [320]. Pseudo [1336]. Pseudoperfect [1083]. Pseudospheres [832]. Publish [465]. Pure [358]. Pursuit [370, 1489]. Putnam [107, 235, 347, 483, 648, 830, 1002, 1155, 1289, 1429, 1057, 1055, 820]. Puzzle [770, 1306, 1315, 1372]. Puzzles [1166, 621]. Pythagoras [842, 1492]. Pythagorean [1350, 631, 722, 1339, 637, 726, 231, 774, 491, 946]. Pythagorus [660].

Quadratic [1072, 1304, 802, 683, 82, 544, 743, 200, 1443, 70, 1462, 169]. Quadrature [656]. Quadrilaterals [1348, 777]. Quadruples [300]. Quantization [92]. Quantum [80, 291]. Quartic [1419]. Quasi [54].

Quasi-Cauchy [54]. Quasiconvex [982]. Quaternions [617]. Question [123, 738]. Questions [208]. Quick [510]. Quotient [632, 233]. Quotients [505, 1043].

R [1418, 148]. Rabbits [187]. Race [927]. Radial [919]. Radical [914]. Radicals [260, 446, 528, 859]. Ramanujan [94, 1146, 244, 251, 476, 710, 1196, 701]. Ramsey [818]. Random [1245, 1347, 840, 284, 1447, 883, 41, 653, 578]. Randomly [335]. Randomness [1482, 314]. Range [1203, 988]. Rank [1201]. Rate [175, 893]. Rates [1352]. Ratio [1437, 95]. Rational [1290, 611, 1348, 964, 941, 155, 1085]. Rationals [144, 249]. Ratios [1424, 517]. Raton [590]. Rayleigh [432]. Raymond [1221]. Real [685, 1166, 534, 1184, 361, 805, 781, 1239, 441, 282, 1216, 440, 878, 767, 782, 1448, 1491, 1482, 627, 519]. Real-Algebraic [361]. Real-Valued [534, 878]. Realization [44]. Really [877, 322, 198]. Reals [144]. Rearrangement [69]. Rearrangements [1228]. reason [1498, 645]. Reciprocal [499, 1054]. Reciprocals [293, 725]. Reciprocity [802, 70, 169, 1462]. Reconstruction [234]. Record [621]. Rectangles [1097]. Rectangular [1022, 67]. Recurrence [729, 1102, 297, 1253, 314, 1482]. Recurrences [612, 765]. Recurrent [58]. Recursion [1057, 834]. Recursive [175, 543]. Recursively [1368]. Rediscovery [753]. Reduction [958]. Redux [1143, 544]. Reed [1238]. Referees [384, 521, 690, 889, 1048, 1194, 1329, 1474]. Refinement [635]. Refinements [1230]. Reflected [775]. Reflections [153, 614]. Regarding [123, 1267]. Region [1024]. Regions [246, 16]. Regular [1174, 1087, 720, 174, 22, 1175]. Regularity [22]. Rejection [613]. Related [1172, 1402, 1355, 1057, 1459, 1367, 1380, 834, 1322]. Relates [291]. Relation [1102, 616, 1253, 819, 617, 564]. Relations [832, 1342]. Relationships [779]. Relativity [82, 1408]. Remainder [485]. Remark [185, 1469, 701]. Rental [1332]. Repeatedly [597, 157]. Replacement [605, 799]. Representation [594, 1412, 762, 960, 1295]. Representations [1455, 295, 234, 206]. Represented [683]. Representing [1338, 901]. Repulsion [218]. Repunits [620]. Research [1172, 1178, 1500, 116, 420]. Resembling [903]. Residue [1197]. Residues [1072, 743]. Resisting [614]. Resolvability [864]. Resolvent [1213]. Restrictions [1409]. Result [1468, 486]. Resulting [775]. Results [651, 40, 528, 132]. Rethinking [595]. Reuben [287]. Reveals [1333]. Reverse [440, 135]. Review [287, 382, 969, 1192, 1261, 241, 645, 554, 465, 113, 344, 810, 1105, 189, 1221, 264, 276, 731, 788, 1033, 952, 907, 396, 87, 887, 214, 866, 519, 867, 749, 1119, 985, 937, 1152, 202, 1014, 493, 574, 1090, 555, 177, 660, 479, 827, 1047, 507, 420, 480, 409, 48, 11, 24, 36, 62, 74, 100, 126, 164, 590, 923, 1136, 148, 421, 330, 253, 999, 536, 713, 358, 302, 434, 627, 1247, 767, 847, 607, 149, 370, 314, 688, 137, 227, 450, 419]. Reviewed [713]. Reviews [1384, 1358, 1442, 1314, 1428, 1327, 1471, 1414, 1288, 1301, 1399, 1064, 1275, 1456, 1344, 969, 952, 907, 887, 866, 867, 985, 937, 923, 847]. Revisited [1361, 1187, 918, 30, 160, 992, 1422, 237, 883, 1006, 801, 820, 1223].

Revisiting [5]. Revolution [425, 305]. Rhode [1047]. RI [554, 465, 344, 396, 493, 555, 713, 358, 688, 419]. Richard [688]. Riddles [366]. Riemann [1045, 1261, 1459, 947, 295, 406, 1211, 69, 271, 1132, 844, 83]. Riesz [1219, 1295]. Right [1290, 249, 157]. Rigorous [512]. Ring [534, 1188, 1115, 1280, 1353]. Ring-theoretic [1353]. Rings [442, 1268, 200, 133, 796]. Rise [731, 1495]. Road [1036]. Roads [205]. Robbins [512]. Roberts [1192]. Robledo [396]. Rodney [314]. Roll [418]. Rolle [366]. Rolling [1160, 775, 318]. Roommate [1332]. Root [1184, 97, 1147, 1369, 283]. Roots [1029, 349, 1323, 755, 441, 709, 1074, 1298, 1364]. Ropes [140]. Ross [39]. Rota [330, 1481]. Rotating [815, 1070]. Rotations [617]. Rotor [988]. Roulette [1160]. Roulettes [205]. Round [831, 1363]. Routh [502, 1110]. Rubik [403]. Ruderman [224]. Ruehr [1361]. Rule [1302, 959, 641, 1440, 515]. Rules [611, 656, 981]. Run [335]. Rupert [1263, 1264, 1463]. Ruti [382].

S [465, 574, 330, 713]. S. [244, 710]. Sacks [1276]. Saddle [741]. salesman [1489, 370]. Sally [358]. Same [157]. Sample [284]. Sampling [1144, 605]. Sandholm [302]. Sandpile [893]. Sank [333]. Satisfying [442]. Scheme [175]. Schemes [598]. Schleicher [420]. Schneps [607]. Schönemann [129]. School [1095, 104]. Schoolgirl [142]. Schur [94, 300, 896]. Schwartz [688]. Schwarz [117, 824, 804, 837, 445]. science [1498, 11]. Scientific [1166, 1221]. Scores [1037]. Seat [624]. Secant [785, 455]. Secants [994]. Second [1269, 465, 808, 507, 347]. Secret [298, 510]. Secretive [1332]. Sections [916, 378]. Seeing [1236]. Seems [529]. seen [778]. Segerman [1247]. Segmentally [654]. Segments [1216]. Selection [1420]. Self [1182, 1374, 1306, 1315]. Self-Descriptive [1306, 1315]. Self-inversive [1374]. Self-Similar [1182]. Semi [885]. Semi-Finite [885]. Semigroups [1177, 677, 961]. Semiperimeter [231]. Senechal [11]. Sentence [580]. Separating [194]. Separation [1184, 934]. Sequence [740, 269, 1114, 157, 91]. Sequences [109, 1355, 1256, 950, 54, 296, 348, 207, 930, 1253, 856, 873]. Serge [419]. Series [950, 1094, 563, 1271, 400, 1102, 212, 1368, 69, 1228, 1370, 206, 863, 259, 762, 1242, 272, 1312, 1225, 725, 285, 654, 358, 618, 1244]. Service [1302, 1080, 1222, 39, 167, 1359, 734, 926, 558, 279, 424]. Set [1331, 632, 642, 595, 1254, 28, 134, 1235]. Sets [759, 1447, 1410, 233, 1159, 1460, 1038, 1129, 207, 1164, 634, 232, 1070, 562, 1084, 856, 1130, 526, 857, 44, 194, 1175, 1448]. Seventeenth [1232]. Seventh [1155]. Seventieth [107]. Seventy [1055, 1289, 1429, 235, 347, 483, 648, 830, 1002, 1155]. Seventy-Eighth [1289]. Seventy-Fifth [830]. Seventy-First [235]. Seventy-Five [1055]. Seventy-Fourth [648]. Seventy-Ninth [1429]. Seventy-Second [347]. Seventy-Seventh [1155]. Seventy-Sixth [1002]. Seventy-Third [483]. Several [457]. Shadows [1425]. Shahriar [465]. Shape [1140, 752, 11].

Shared [334]. Sharhriari [465]. Sharing [1126]. Sharkovsky [171]. Sharp [719]. Sheldon [1430]. Shi [816]. Shift [1212]. Shing [788]. Shing-Tung [788]. Short [764, 432, 145, 572, 1257, 838, 547, 56, 967, 1134, 121, 110, 161, 1202, 1270, 1286]. Shorten [1381]. Shortening [825]. Should [940, 196]. Shouldn’t [2]. Shouting [1372]. Shuffles [974]. Side [778, 536]. Sided [1061]. Siebeck [1116, 71]. Siegel [1080]. Sierpiński [1396]. Sight [580, 1277, 536]. Signed [1390, 141, 477]. Signs [931, 1377]. Similar [1182, 341]. Simple [159, 109, 474, 625, 9, 394, 1082, 1396, 773, 738, 1228, 794, 1028, 1366, 1450, 486, 960, 110, 1346, 477, 1295, 549, 198, 806, 878, 843, 404]. Simpler [222, 1395]. Simplex [1174, 720]. Simplices [1110]. Simplicity [1376]. Simply [892]. Simpson [641]. Simultaneous [1124, 1176, 699]. Sinc [81, 333, 651]. Sine [823, 541, 328, 566]. Sines [1298]. Single [1037]. Singletons [1254]. Singular [892, 184, 1205]. Sinusoidal [318]. Six [992, 1438]. Sixpartite [1281]. Sixth [1002]. Sizes [1044]. Sklar [574]. Skolem [1189]. Slabs [1349]. Sleeping [1139]. Slender [668]. Slender-Body [668]. Slice [980]. Slightness [200, 962]. Slinky [1053]. Slow [1256, 930]. Small [1075, 963, 130]. Smaller [1181]. Smallest [1426, 1355, 1377, 1299]. Smith [507]. Smooth [1325, 934]. Smoothing [1461]. Smullyan [1221]. Sneaky [431]. Snowflake [918, 997]. Social [1294]. Societies [3, 4, 1199]. Society [554, 465, 344, 396, 493, 555, 1047, 713, 358, 688, 419]. Soifer [479]. Solids [1263, 548]. Solitaire [485]. Solomon [1238]. Solution [1355, 1400, 612, 901, 913, 335, 1272, 1315, 1299, 600, 579]. Solutions [112, 136, 147, 163, 176, 188, 201, 213, 226, 240, 252, 263, 275, 286, 301, 313, 329, 343, 357, 369, 381, 384, 395, 408, 433, 449, 464, 478, 492, 506, 518, 521, 535, 553, 573, 589, 606, 626, 644, 659, 686, 687, 712, 730, 748, 766, 787, 809, 826, 846, 865, 886, 1046, 906, 922, 936, 951, 968, 984, 998, 1013, 1032, 1191, 1063, 1076, 1089, 1104, 1118, 1135, 1151, 1165, 1220, 10, 35, 47, 61, 73, 86, 99, 1207, 1234, 1246, 1260, 1274, 1287, 1300, 1313, 1326, 1343, 1357, 1371]. Solutions [1383, 1398, 1413, 1427, 1441, 1454, 1470, 23, 125, 1348, 855, 1127, 79]. Solvable [586]. Solve [739]. Solved [384, 521]. Solving [181, 267]. Some [947, 780, 66, 566, 272, 800, 860]. Somewhat [702]. Sondow [600]. Sons [409]. Sore [1197]. Sort [1240]. Sound [851]. Sourcebook [1152]. Space [311, 1143, 471, 864, 111, 1162, 718, 1251]. Spaces [1143, 803, 448, 80, 879, 1212, 1325]. Spanning [1159]. Spatial [1024, 415]. Speaking [382, 1480, 1488]. Special [82]. Species [666]. Specified [899]. Spectral [685]. Specular [1236]. Speed [291]. Sperner [428]. Sphere [140, 752, 321]. Sphere-Filling [140]. Spheres [425, 832]. Spherical [756, 342]. Spiral [118, 365]. Spirals [318]. Spirographic [560]. Spivak [465]. Splines [656]. Splitting [374]. Sporadic [1093]. Sports [1302]. Spot [1458]. Spreads [636]. Springer [382, 731, 479, 420, 434, 450]. Springer-Verlag [420]. Square [349, 1067, 1230, 634, 1147, 709, 1298, 1369, 1299, 283]. Squares [1186, 979, 363, 1434, 1161, 1085, 427]. Stability [439]. Stabilization [1404]. Stable [1108, 1465]. Stammler [1452]. Stamping [183]. Standard [284]. Standards [196]. Standards-Based [196]. Stanley [1322]. Stars [317]. Statement [868]. States [484]. Stein [1261]. Steiner [287, 1110]. Steinhaus [250]. Stereographic [281, 703]. Stern [1114, 91]. Steve [788]. Steven [713]. Still [559, 1209]. Stirling [488, 583, 547, 584]. Stochastic [702, 1385]. Stock [1229]. Stöer [1214]. Stokes [1307]. Stone [552]. Stone-Cech [552]. store.doverpublications.com [1221]. store.doverpublications.com/0486492370.html [1221]. Stories [344, 1476]. Story [1460]. Stråhle [1020]. Straight [337, 16]. Straightforward [1440]. Strange [1271]. Strategies [581]. Strategy [1041]. Streamlined [124]. String [206]. Stroll [1142]. Strong [919, 456, 963, 873]. Student [1178, 396, 1047]. Student-Driven [1178]. students [1475, 358]. Sturm [320, 361]. Sturmian [792]. style [1335]. Sub [772]. Sub- [772]. Subalgebras [1115]. Subdivision [471]. Subgroups [899, 744]. Submajorization [290]. Subordinate [1212]. Subrings [1394]. Subseries [396, 1352]. Subset [122]. Subsets [341, 1239]. Subspaces [643]. Substitution [894]. Substitutions [21]. Subsum [856]. Subsums [1370, 1242]. Success [1027]. Such [657]. Sudoku [363]. Sufficient [282]. Suggested [1250]. Sum [1426, 788, 901, 543, 544, 206, 259, 1157, 211, 1161, 1397, 820, 268, 1286]. Summing [1090]. Sums [1072, 1424, 499, 306, 365, 884, 7, 994, 293, 634, 261, 566, 1132, 641, 996, 920, 1386]. Sunny [536]. Supercyclicity [448]. Superfactorials [772]. Superpowered [1098]. Surd [1443]. Surface [832, 815, 305, 454, 1183]. Surfaces [957, 1236, 90, 1021, 741, 1487, 688]. Surjective [1204, 902]. Surjectivity [1325, 1086]. Surprise [758]. Surprised [168]. Surprises [983]. Surprising [1387]. Sury [616, 1025]. Süss [861]. SVD [375]. Sylow [197, 1336]. Symbols [354]. Symmetric [1406, 840, 377, 781, 183, 1411]. Symmetrically [90]. Symmetries [57]. Symmetry [391, 851, 1494, 450]. Synthetic [671, 905, 600]. System [27, 718]. Systems [546, 1262, 1404, 1069, 911, 525, 997, 1465, 1482, 314]. Szego [680].

Tabachnikov [419]. Tails [633]. Take [2, 941]. Tale [103, 989]. Tangency [706, 600]. Tangent [785]. Tangents [994, 308]. Tanks [498]. Tanvolutes [106]. Tao [493]. Tapp [450]. Tarski [1124]. Tau [701]. Taylor [1363]. Teaching [153, 1453, 116]. Technique [604]. Telescoping [1100, 996]. Television [574, 1493]. tell [1498]. Tempered [1020]. Terence [493]. Terms [1102, 1368, 156]. Ternary [642, 1419]. Territorial [665]. Tessellation [27]. Test [863]. Tests [196]. Tetrahedra [684, 402]. Tetrahedron [142, 1463, 1233]. Textbooks [465]. Texts [358]. th [1253, 1397]. Their [246, 1238, 1345, 396, 1348, 52, 351, 1485]. Them [2, 1190, 1070]. Theme [740, 442]. Theodorus [365, 1365]. Theorem [685, 1317, 1045, 625, 1082, 1156, 222, 546, 1116, 153, 239, 299, 894, 529, 729, 1356, 726, 708, 361, 1367, 69, 552, 794, 774, 1392, 1028, 992, 237, 1031, 904, 401, 95, 596, 883, 72, 491, 41, 532, 603, 250, 1325, 84, 1162, 1110, 551, 967, 934, 45, 405, 1450, 71, 1006, 1388, 895, 33, 525, 1040, 678, 623, 161, 682, 653, 274, 881, 1216, 1088, 158, 725, 1346, 477, 1405, 1061, 28, 1295, 1150, 198, 1101, 699, 1189, 844, 799, 806, 1117, 83, 843, 513, 680, 1265, 427, 793, 882, 635, 362, 20, 772, 1279, 1269, 764]. Theorem [1307, 502, 1382, 1458, 432, 530, 1396, 171, 145, 572, 1174, 185, 1257, 407, 542, 580, 1453, 1230, 1419, 322, 1011, 485, 761, 946, 1096, 319, 1366, 1322, 836, 1395, 1364, 197, 489, 366, 1231, 1449, 1131, 1324]. Theorems [1345, 777, 1366, 1417, 351, 170, 989]. Theoretic [1405, 1353]. Theoretical [1178]. Theorist [336, 778]. Theory [387, 143, 499, 1477, 649, 1435, 438, 595, 417, 1176, 668, 736, 376, 1183, 1495, 731, 493, 409, 1090]. There [388, 1073]. Thin [296]. Thing [300]. Think [749]. Third [483]. Thirty [419, 1478]. Thomas [409]. Thomson [465]. Thoughts [1200]. Thrackles [210]. Three [763, 928, 615, 749, 46, 1162, 752, 1417, 920, 989, 1231]. Three-Body [752]. Throwing [1187]. Thumb [1197]. Tides [1094]. Tiling [1182, 977, 379]. Tilings [756]. Time [237, 319]. Times [639]. Tire [414]. Title [384, 521]. Today [510]. Too [1460]. Top [187]. Topological [250]. Topologically [911]. Topology [103, 355, 736]. Torricelli [579]. Torus [403]. Tossing [298, 624]. Totals [1276]. Totient [1353]. Tour [767]. Tournaments [945]. Tower [1200, 57]. Toy [380]. Trace [838]. Tracing [666]. Tracks [414]. Tractrices [414]. Tradition [382, 1488]. Trajectory [1351]. Transcendence [389]. Transcendental [532, 638, 871]. Transcending [1488, 382]. Transform [6, 1433, 219]. Transit [577]. Transition [596]. Transitive [911, 52]. Translates [639]. Translation [465]. Transpositions [542]. Transversals [852, 67]. Trapezoids [1382]. Traveling [370, 1489]. Treatment [822]. Tree [771, 1443, 1202]. Treisman [1359]. Tri [1356]. Tri-cycle [1356]. Trial [607, 1497]. Triangle [502, 880, 1007, 1130, 1058]. Triangles [1095, 471, 1339, 104, 1005, 1347, 874, 1290, 637, 1435, 876, 684, 977, 352, 1058]. Triangular [795, 1251]. Triangulations [459, 652]. Trick [230]. Trigonometric [312, 1219, 703, 53, 1255, 58]. Triple [885]. Triples [1350, 231]. Trisecting [637]. Tritangency [170]. Triumph [1020]. Tropical [630]. Trouble [29]. True [1460]. Truncated [1463]. Tung [788]. Turán [325]. Turning [1172]. Twice [1293]. Twin [405]. Twisted [1415]. Twisting [548]. Two [1426, 1361, 540, 1113, 208, 374, 728, 319, 1025, 1022, 1378, 132, 1161, 1061, 1305]. Two-Chord [319]. Two-Colored [1022]. Two-Sided [1061]. TX [465]. Type [611, 1087, 19, 1252, 598, 239, 678, 1255].

Ulam [428]. Ultimate [825]. Ultrasmall [119]. Umberto [731]. Umlaufsatz [1267]. Unbiased [708, 1061, 1449]. Unboundedness [919]. Unconstrained [282]. Uncountability [95, 638]. Uncountable [805, 1239, 1448]. Uncountably [1448]. Undergraduate [1172, 468, 358, 484]. Undetermined [739]. Undominated [581].

Unexpected [915]. Ungar [382]. Unified [894, 1136]. Uniform [803, 757, 872, 1295, 198, 878]. Unique [427]. Uniquely [1386, 879]. Uniqueness [894, 533, 79]. Unit [901, 964, 634, 1216]. United [484]. Units [1188]. Unity [755, 1074]. Universe [932]. University [287, 1192, 788, 519, 660, 927, 330, 370]. Unknots [593]. Unknotting [593]. Unknowable [291]. Unnoticed [530, 680]. Unordered [290]. Unreasonable [200, 962]. Unsettleable [413]. Unsolved [1209]. untitled [241, 113, 189, 264, 276, 87, 214, 202, 177, 48, 24, 36, 62, 74, 100, 126, 164, 253, 137, 227]. Unveiled [1109]. Update [1201]. Upon [559]. Upper [780, 342]. Uri [1359]. Urn [598, 41, 327]. USAMO [905]. Use [391]. Used [913, 607, 1497]. Useful [1468]. Using [468, 1141, 775, 916, 471, 1339, 739, 1377, 27, 1262, 785, 1461, 670, 995, 528, 353, 1366, 981, 960, 285, 473, 618, 1433, 1297, 996].

V [148]. Vadim [1198]. Vakil [519]. Valid [747]. Value [1094, 1128, 580, 1113, 1453, 836, 1088, 806, 1117]. Valued [534, 940, 878, 1446]. Values [306, 979, 941, 838, 110, 933, 531, 58, 365]. Vandermonde [834, 1252]. Variable [1045, 739, 844]. Variable-Coefficient [739]. Variables [1245, 840, 284]. Variants [1185]. Variation [309, 416]. Variational [490]. Variations [740, 442, 1012, 1372, 28, 1499, 1047]. Vec [709]. Vector [111, 352, 1136]. Venus [577]. Vera [287]. Verhulst [282]. Verlag [382, 420]. Version [532, 84, 1088, 1061, 1449, 806]. Versions [1131]. Versus [624]. Vertex [1351, 634, 1437]. Vertical [614]. Vertices [1425, 980]. Very [1134]. vi [1221]. via [447, 1125, 511, 1174, 1114, 1435, 361, 212, 552, 72, 491, 41, 1144, 532, 949, 211, 68, 765, 1305]. Vicsek [997]. Victor [555]. Video [1041, 421]. Viète [1291]. View [336, 1311, 376, 555]. vii [358]. viii [465, 574]. Viscosity [837]. Visibility [1277]. Visual [393, 1251]. Visualizing [1247]. Vitali [232, 477]. Vol [1047, 396]. Volterra [79]. Volume [150, 832, 1390, 783, 1024, 353, 1233]. Volume/Surface [832]. Voronoï [27]. Voting [3, 1199]. vs [1467].

W. [734]. Waerden [839]. Wahba [1070]. Walk [988, 883, 653, 578]. Walks [41, 1144]. Wall [1458]. Wallis [94, 598, 821, 760, 565, 19]. Wallis- [19]. Wallis-type [598]. Walls [274]. Walnut [465]. Wanted [1303]. War [367, 294]. War-like [367]. Warning [1269]. Was [168]. Washington [465, 810, 507, 536, 627, 314]. Wave [219]. Way [105, 259, 93, 1056, 1481, 330]. Wedges [425]. Weierstrass [368, 51, 489]. Weighted [567]. Well [707]. Well-Classification [707]. Well-Ordering [707]. Wellesley [148]. Were [1303]. Weyl [376]. Wheels [205]. Where [193, 1299, 1085]. Which [874, 747, 1491, 16, 913, 627]. Who [1105]. Whose [1128, 122, 463]. Wikler [11]. Wild [1278]. Wiley [409]. Wilf [975, 392, 1443]. William [1261, 107, 235, 347, 483, 648, 830, 1002, 1155, 1289, 1429, 302, 370, 424, 927]. Wilson [772]. Win [978]. Winding [1036]. within [426]. Without [864, 981, 655, 605, 282]. Women [344, 1476]. Words [335, 655]. Work [1171, 181, 617, 169, 713]. Works [1282]. World [1221]. Worlds [749]. worldscibooks [1221]. worldscientific.com [1221]. worldscientific.com/worldscibooks/10.1142/9865 [1221]. Worpitzky [1223]. Would [406]. Writing [11]. Wronskians [108]. www.dimensions [421]. www.dimensions-math.org [421].

x [465, 1221, 1047, 627]. xi [382, 607]. Xiang [824]. xii [287, 148, 330]. xiii [554, 370, 688]. xiv [645, 465, 420, 434, 450]. xix [519]. xv [419]. xvi [465, 493, 590]. xvii [465, 731]. xviii [810, 788, 713, 314]. xxii [660]. xxiv [409]. xxv [302]. xxx [479].

Yan [330]. Yang [118]. Yanofsky [645]. Yau [788]. Ye [434]. Year [465, 697, 414]. Years [1198, 788, 1055, 1393, 691, 692, 698]. Yin [118]. Yinyu [434]. York [731, 479, 330, 434, 607, 450]. You’re [1292]. Yueh [1080, 1222, 39, 167, 1359, 734, 926, 558, 279, 424]. Yueh-Gin [1080, 1222, 39, 167, 1359, 734, 926, 558, 279, 424].

Zagier [123]. Zariski [258]. Zbigniew [465]. Zeckendorf [206]. Zeilberger [392, 392]. Zero [634]. Zeros [173, 1362]. Zeta [1459, 365, 563, 295, 585, 1132, 1309]. Zeta-values [365]. Zhai [810]. Zwicky [11].

References

Anonymous:2010:FMa

[1] Anonymous. Front matter. American Mathematical Monthly, 117(1): ??, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

      0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
      4169/amermathmont.117.1.FM.pdf.

Abrams:2010:FGB

[2] Aaron Abrams and Skip Garibaldi. Finding good bets in the lottery, and why you shouldn’t take them. American Mathematical Monthly, 117(1): 3–26, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

      0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
      4169/000298910X474952.pdf.

Berg:2010:VAS

[3] Deborah E. Berg, Serguei Norine, Francis Edward Su, Robin Thomas, and Paul Wollan. Voting in agreeable societies. American Mathematical Monthly, 117(1):27–39, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

      org/stable/pdfplus/10.4169/000298910X474961.pdf.

Hardin:2010:ACS

[4] Christopher S. Hardin. Agreement in circular societies. American Mathematical Monthly, 117(1):40–49, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X474970.pdf.

Lewis:2010:RPM

[5] Barry Lewis. Revisiting the Pascal matrix. American Mathematical Monthly, 117(1):50–66, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/000298910X474989.pdf.

Cirnu:2010:NIL

[6] Mircea I. Cîrnu. Newton’s identities and the Laplace transform. American Mathematical Monthly, 117(1):67–71, January 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X474998.pdf.

Girstmair:2010:FSD

[7] Kurt Girstmair. Farey sums and Dedekind sums. American Mathematical Monthly, 117(1):72–78, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X475005.pdf.

Pillichshammer:2010:ECA

[8] Friedrich Pillichshammer. Euler’s constant and averages of fractional parts. American Mathematical Monthly, 117(1):78–83, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X475014.pdf.

Bagby:2010:SP

[9] Richard Bagby. A simple proof that Γ′(1) = −γ. American Mathematical Monthly, 117(1):83–85, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X475023.pdf.

Edgar:2010:PSa

[10] Gerald A. Edgar, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(1):86–93, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X475032.pdf.

Nunemacher:2010:BRBa

[11] Jeffrey Nunemacher and Amir Alexander. Book review: The Shape of Content: Creative Writing in Mathematics and Science, by Chandler Davis; Marjorie Wikler Senechal; Jan Zwicky. American Mathematical Monthly, 117(1):94–96, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X475041.pdf.

Anonymous:2010:BMa

[12] Anonymous. Back matter. American Mathematical Monthly, 117(1): ??, January 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.1.BM.pdf.

Anonymous:2010:FMb

[13] Anonymous. Front matter. American Mathematical Monthly, 117(10): ??, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.10.FM.pdf.

Farris:2010:GIM

[14] Frank A. Farris. The Gini index and measures of inequality. American Mathematical Monthly, 117(10):851–864, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X523344.pdf.

Gordon:2010:MFO

[15] Gary Gordon and Elizabeth McMahon. Moving faces to other places: Facet derangements. American Mathematical Monthly, 117(10):865–880, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X523353.pdf.

Ivanov:2010:NRW

[16] Oleg A. Ivanov. On the number of regions into which n straight lines divide the plane. American Mathematical Monthly, 117(10):881–888, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X523362.pdf.

Langer:2010:WCO

[17] Joel C. Langer and David A. Singer. When is a curve an octahedron? American Mathematical Monthly, 117(10):889–902, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X523371.pdf.

Laurent:2010:LGP

[18] Monique Laurent and Alexander Schrijver. On Leonid Gurvits’s proof for permanents. American Mathematical Monthly, 117(10):903–911, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X523380.pdf.

Sondow:2010:NWC

[19] Jonathan Sondow and Huang Yi. New Wallis- and Catalan-type infinite products for π, e and \sqrt{2 + \sqrt{2}}. American Mathematical Monthly, 117 (10):912–917, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X523399.pdf.

deSilva:2010:CEP

[20] Nadish de Silva. A concise, elementary proof of Arzelà’s bounded convergence theorem. American Mathematical Monthly, 117(10):918–920, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X523407.pdf.

Bessenyei:2010:FEF

[21] Mihály Bessenyei. Functional equations and finite groups of substitutions. American Mathematical Monthly, 117(10):921–927, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X523416.pdf.

Liu:2010:RON

[22] Shibo Liu. On the regularity of operators near a regular operator. American Mathematical Monthly, 117(10):927–928, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X523425.pdf.

Edger:2010:PSa

[23] Gerald A. Edger, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(10):929–936, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X523434.pdf.

Nunemacher:2010:BRBb

[24] Jeffrey Nunemacher and Ethan Akin. Book review: [untitled]. American Mathematical Monthly, 117(10):937–940, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X523443.pdf.

Anonymous:2010:BMb

[25] Anonymous. Back matter. American Mathematical Monthly, 117(10): ??, December 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.10.BM.pdf.

Anonymous:2010:FMc

[26] Anonymous. Front matter. American Mathematical Monthly, 117(2): ??, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.2.FM.pdf.

Frank:2010:DSU

[27] Natalie Priebe Frank and Sean M. Hart. A dynamical system using the Voronoï tessellation. American Mathematical Monthly, 117(2):99– 112, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X476022.pdf.

Sherman:2010:VKS

[28] David Sherman. Variations on Kuratowski’s 14-set theorem. American Mathematical Monthly, 117(2):113–123, February 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X476031.pdf.

Keleti:2010:TKC

[29] Tamás Keleti and Elliot Paquette. The trouble with von Koch curves built from n gons. American Mathematical Monthly, 117(2):124–137, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X476040.pdf.

Dumbgen:2010:NIR

[30] Lutz Dümbgen, Sara A. van de Geer, Mark C. Veraar, and Jon A. Wellner. Nemirovski’s inequalities revisited. American Mathematical Monthly, 117(2):138–160, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X476059.pdf.

Wetzel:2010:AEL

[31] John E. Wetzel and Ed Scheinerman. An ancient elliptic locus. American Mathematical Monthly, 117(2):161–167, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X476068.pdf.

Pudwell:2010:CIE

[32] Lara K. Pudwell and Eric S. Rowland. Counting interesting elections. American Mathematical Monthly, 117(2):167–174, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdfplus/10.4169/
000298910X476077.pdf.

Nadler:2010:PDT

[33] Sam B. Nadler, Jr. A proof of Darboux’s theorem. American Mathematical Monthly, 117(2):174–175, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X476086.pdf.

Lesigne:2010:BII

[34] Emmanuel Lesigne. On the behavior at infinity of an integrable function. American Mathematical Monthly, 117(2):175–181, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X476095.pdf.

Edgar:2010:PSb

[35] Gerald A. Edgar, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(2):182–189, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X476103.pdf.

Nunemacher:2010:BRBc

[36] Jeffrey Nunemacher and Daniel S. Silver. Book review: [untitled]. American Mathematical Monthly, 117(2):190–192, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X476112.pdf.

Anonymous:2010:BMc

[37] Anonymous. Back matter. American Mathematical Monthly, 117(2): ??, February 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.2.BM.pdf.

Anonymous:2010:FMd

[38] Anonymous. Front matter. American Mathematical Monthly, 117(3): ??, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.3.FM.pdf.

Faires:2010:YGG

[39] Barbara Faires. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2010 to Kenneth A. Ross for Distinguished Service to Mathematics. American Mathematical Monthly, 117(3):195–197, March 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X480054.pdf.

Greenberg:2010:ONR

[40] Marvin Jay Greenberg. Old and new results in the foundations of elementary plane Euclidean and non-Euclidean geometries. American Mathematical Monthly, 117(3):198–219, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/pdfplus/10.4169/000298910X480063.pdf.

Levin:2010:PTR

[41] David A. Levin and Yuval Peres. Pólya’s theorem on random walks via Pólya’s urn. American Mathematical Monthly, 117(3):220–231, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480072.pdf.

Borisov:2010:CPP

[42] Alexander Borisov, Mark Dickinson, and Stuart Hastings. A congruence problem for polyhedra. American Mathematical Monthly, 117(3):232– 249, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480081.pdf.

Hilmar:2010:EMB

[43] Jan Hilmar and Chris Smyth. Euclid meets Bézout: Intersecting algebraic plane curves with the Euclidean algorithm. American Mathematical Monthly, 117(3):250–260, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X480090.pdf.

Tanaka:2010:RMS

[44] Jun Tanaka and Peter F. McLoughlin. A realization of measurable sets as limit points. American Mathematical Monthly, 117(3):261– 266, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480108.pdf.

McQuillan:2010:PTD

[45] Dan McQuillan and R. Bruce Richter. A parity theorem for drawings of complete and complete bipartite graphs. American Mathematical Monthly, 117(3):267–273, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X480117.pdf.

Khattri:2010:TPI

[46] Sanjay K. Khattri. Three proofs of the inequality e < (1 + \frac{1}{n})^{n+0.5}. American Mathematical Monthly, 117(3):273–277, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X480126.pdf.

Edgar:2010:PSc

[47] Gerald A. Edgar, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(3):278–285, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480135.pdf.

Nerode:2010:BRB

[48] Anil Nerode. Book review: [untitled]. American Mathematical Monthly, 117(3):286–288, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X480144.pdf.

Anonymous:2010:BMd

[49] Anonymous. Back matter. American Mathematical Monthly, 117(3): ??, March 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.3.BM.pdf.

Anonymous:2010:FMe

[50] Anonymous. Front matter. American Mathematical Monthly, 117(4): ??, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.4.FM.pdf.

Rice:2010:FPC

[51] Adrian Rice. “to a factor près”: Cayley’s partial anticipation of the Weierstrass ℘-function. American Mathematical Monthly, 117(4):291– 302, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480766.pdf.

Pearce:2010:TDG

[52] Geoffrey Pearce. Transitive decompositions of graphs and their links with geometry and origami. American Mathematical Monthly, 117(4): 303–310, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480775.pdf.

Johnson:2010:TIH

[53] Warren P. Johnson. Trigonometric identities à la Hermite. American Mathematical Monthly, 117(4):311–327, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/pdfplus/10.4169/000298910X480784.pdf.

Burton:2010:QCS

[54] David Burton and John Coleman. Quasi-Cauchy sequences. American Mathematical Monthly, 117(4):328–333, April 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X480793.pdf.

Bennett:2010:FI

[55] Grahame Bennett. p-free ℓ^{p} inequalities. American Mathematical Monthly, 117(4):334–351, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/000298910X480801.pdf.

Marshall:2010:SP

[56] Timothy Marshall. A short proof of ζ(2) = π^{2}/6. American Mathematical Monthly, 117(4):352–353, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X480810.pdf.

Park:2010:GST

[57] So Eun Park. The group of symmetries of the Tower of Hanoi graph. American Mathematical Monthly, 117(4):353–360, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X480829.pdf.

Zhou:2010:RPI

[58] Li Zhou and Lubomir Markov. Recurrent proofs of the irrationality of certain trigonometric values. American Mathematical Monthly, 117(4): 360–362, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480838.pdf.

Curran:2010:GCF

[59] M. John Curran. Groups of cube-free odd order. American Mathematical Monthly, 117(4):363–365, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X480847.pdf.

Kozek:2010:AFG

[60] Mark Kozek. An asymptotic formula for Goldbach’s conjecture with monic polynomials in Z[x]. American Mathematical Monthly, 117(4): 365–369, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480856.pdf.

Edgar:2010:PSd

[61] Gerald A. Edgar, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(4):370–377, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X480865.pdf.

Nunemacher:2010:BRBd

[62] Jeffrey Nunemacher and Sanford L. Segal. Book review: [untitled]. American Mathematical Monthly, 117(4):378–380, April 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X480874.pdf.

Anonymous:2010:BMe

[63] Anonymous. Back matter. American Mathematical Monthly, 117(4): ??, April 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.4.BM.pdf.

Anonymous:2010:FMf

[64] Anonymous. Front matter. American Mathematical Monthly, 117(5): ??, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.5.FM.pdf.

Richey:2010:EMC

[65] Matthew Richey. The evolution of Markov chain Monte Carlo methods. American Mathematical Monthly, 117(5):383–413, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X485923.pdf.

Kharazishvili:2010:SCE

[66] A. B. Kharazishvili and T. Sh. Tetunashvili. On some coverings of the Euclidean plane with pairwise congruent circles. American Mathematical Monthly, 117(5):414–423, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X485932.pdf.

Stein:2010:TRA

[67] Sherman Stein. Transversals in rectangular arrays. American Mathematical Monthly, 117(5):424–433, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X485941.pdf.

Mortici:2010:PAA

[68] Cristinel Mortici. Product approximations via asymptotic integration. American Mathematical Monthly, 117(5):434–441, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X485950.pdf.

Freniche:2010:RRT

[69] Francisco J. Freniche. On Riemann’s rearrangement theorem for the alternating harmonic series. American Mathematical Monthly, 117(5):

442–448, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X485969.pdf.

Steiner:2010:MDF

[70] Richard Steiner. Modular divisor functions and quadratic reciprocity. American Mathematical Monthly, 117(5):448–451, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X485978.pdf.

Mercer:2010:AAS

[71] Peter R. Mercer. Another application of Siebeck’s theorem. American Mathematical Monthly, 117(5):452–454, May 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X485987.pdf.

Lazer:2010:FTA

[72] Alan C. Lazer and Mark Leckband. The fundamental theorem of algebra via the Fourier inversion formula. American Mathematical Monthly, 117(5):455–457, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X485996.pdf.

Edgar:2010:PSe

[73] Gerald A. Edgar, Dough Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(5):458– 465, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X486003.pdf.

Nunemacher:2010:BRBe

[74] Jeffrey Nunemacher and Carol S. Schumacher. Book review: [untitled]. American Mathematical Monthly, 117(5):466–468, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X486012.pdf.

Velleman:2010:EE

[75] Daniel J. Velleman. Editor’s endnotes. American Mathematical Monthly, 117(5):469–470, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X486021.pdf.

Anonymous:2010:BMf

[76] Anonymous. Back matter. American Mathematical Monthly, 117(5): ??, May 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.5.BM.pdf.

Anonymous:2010:FMg

[77] Anonymous. Front matter. American Mathematical Monthly, 117(6): ??, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.6.FM.pdf.

Smith:2010:DNG

[78] James T. Smith. Definitions and nondefinability in geometry. American Mathematical Monthly, 117(6):475–489, June 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X492781.pdf.

Logemann:2010:VFD

[79] Hartmut Logemann and Eugene P. Ryan. Volterra functional differential equations: Existence, uniqueness, and continuation of solutions. American Mathematical Monthly, 117(6):490–511, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X492790.pdf.

Gudder:2010:FQM

[80] Stan Gudder. Finite quantum measure spaces. American Mathematical Monthly, 117(6):512–527, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X492808.pdf.

Borwein:2010:NSF

[81] David Borwein, Jonathan M. Borwein, and Isaac E. Leonard. L_{p} norms and the sinc function. American Mathematical Monthly, 117(6):528– 539, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X492817.pdf.

Elton:2010:IQF

[82] John H. Elton. Indefinite quadratic forms and the invariance of the interval in special relativity. American Mathematical Monthly, 117(6):

540–547, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X492826.pdf.

Thomson:2010:MCT

[83] Brian S. Thomson. Monotone convergence theorem for the Riemann integral. American Mathematical Monthly, 117(6):547–550, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X492835.pdf.

Manetti:2010:PVT

[84] Marco Manetti. A proof of a version of a theorem of Hartogs. American Mathematical Monthly, 117(6):550–552, June 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X492844.pdf.

Jones:2010:DPI

[85] Timothy W. Jones. Discovering and proving that π is irrational. American Mathematical Monthly, 117(6):553–557, June 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X492853.pdf.

Edgar:2010:PSf

[86] Gerald A. Edgar, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(6):558– 565, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X492862.pdf.

Hampton:2010:BRB

[87] Charles R. Hampton and Jeffrey Nunemacher. Book review: [untitled]. American Mathematical Monthly, 117(6):566–568, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X492871.pdf.

Anonymous:2010:BMg

[88] Anonymous. Back matter. American Mathematical Monthly, 117(6): ??, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.6.BM.pdf.

Anonymous:2010:FMh

[89] Anonymous. Front matter. American Mathematical Monthly, 117(7): ??, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.7.FM.pdf.

Cavendish:2010:SBS

[90] William Cavendish and John H. Conway. Symmetrically bordered surfaces. American Mathematical Monthly, 117(7):571–580, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdfplus/10.4169/
000298910X496705.pdf.

Northshield:2010:SDS

[91] Sam Northshield. Stern’s diatomic sequence 0, 1, 1, 2, 1, 3, 2, 3, 1, 4, . . .. American Mathematical Monthly, 117(7):581–598, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X496714.pdf.

Olver:2010:DQ

[92] Peter J. Olver. Dispersive quantization. American Mathematical Monthly, 117(7):599–610, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X496723.pdf.

Stoll:2010:FWO

[93] Thomas Stoll. A fancy way to obtain the binary digits of 759250125\sqrt{2}. American Mathematical Monthly, 117(7):611–617, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X496732.pdf.

Amdeberhan:2010:WRS

[94] T. Amdeberhan, O. Espinosa, V. H. Moll, and A. Straub. Wallis– Ramanujan–Schur–Feynman. American Mathematical Monthly, 117 (7):618–632, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X496741.pdf.

Krebs:2010:CFU

[95] Mike Krebs and Thomas Wright. On Cantor’s first uncountability proof, Pick’s theorem, and the irrationality of the golden ratio. American Mathematical Monthly, 117(7):633–637, August 2010. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/pdfplus/10.4169/000298910X496750.pdf.

Najman:2010:DEG

[96] Filip Najman. The Diophantine equation x^{4} ± y^{4} = iz^{2} in Gaussian integers. American Mathematical Monthly, 117(7):637–641, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X496769.pdf.

Frayer:2010:PRM

[97] Christopher Frayer and James A. Swenson. Polynomial root motion. American Mathematical Monthly, 117(7):641–646, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X496778.pdf.

Fraenkel:2010:EA

[98] Aviezri S. Fraenkel. From enmity to amity. American Mathematical Monthly, 117(7):646–648, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/000298910X496787.pdf.

Edgar:2010:PSg

[99] Gerald A. Edgar, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(7):649–656, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X496796.pdf.

Nunemacher:2010:BRBf

[100] Jeffrey Nunemacher and Peter Sin. Book review: [untitled]. American Mathematical Monthly, 117(7):657–660, August 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X496804.pdf.

Anonymous:2010:BMh

[101] Anonymous. Back matter. American Mathematical Monthly, 117(7): ??, August 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.7.BM.pdf.

Anonymous:2010:FMi

[102] Anonymous. Front matter. American Mathematical Monthly, 117(8): ??, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.8.FM.pdf.

Folland:2010:TT

[103] Gerald B. Folland. A tale of topology. American Mathematical Monthly, 117(8):663–672, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X515730.pdf.

Calcut:2010:GST

[104] Jack S. Calcut. Grade school triangles. American Mathematical Monthly, 117(8):673–685, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X515749.pdf.

Conger:2010:BWD

[105] Mark A. Conger and Jason Howald. A better way to deal the cards. American Mathematical Monthly, 117(8):686–700, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdfplus/10.4169/
000298910X515758.pdf.

Apostol:2010:TGI

[106] Tom M. Apostol and Mamikon A. Mnatsakanian. Tanvolutes: Generalized involutes. American Mathematical Monthly, 117(8):701–713, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X515767.pdf.

Klosinski:2010:SWL

[107] Leonard F. Klosinski, Gerald L. Alexanderson, and Loren C. Larson. The Seventieth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 117(8):714–721, October 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X515776.pdf.

Bostan:2010:WLI

[108] Alin Bostan and Philippe Dumas. Wronskians and linear independence. American Mathematical Monthly, 117(8):722–727, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X515785.pdf.

Alhakim:2010:SCA

[109] Abbas M. Alhakim. A simple combinatorial algorithm for de Bruijn sequences. American Mathematical Monthly, 117(8):728–732, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X515794.pdf.

Nijenhuis:2010:SGP

[110] Albert Nijenhuis. Short gamma products with simple values. American Mathematical Monthly, 117(8):733–737, October 2010. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/000298910X515802.pdf.

Kwon:2010:DLF

[111] Miyeon Kwon. Dimension, linear functionals, and norms in a vector space. American Mathematical Monthly, 117(8):738–740, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X515811.pdf.

Anonymous:2010:PS

[112] Anonymous. Problems and solutions. American Mathematical Monthly, 117(8):741–748, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X515820.pdf.

Cohen:2010:BRB

[113] Marion Deutsche Cohen. Book review: [untitled]. American Mathematical Monthly, 117(8):749–753, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X515839.pdf.

Anonymous:2010:BMi

[114] Anonymous. Back matter. American Mathematical Monthly, 117(8): ??, October 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.8.BM.pdf.

Anonymous:2010:FMj

[115] Anonymous. Front matter. American Mathematical Monthly, 117(9): ??, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.9.FM.pdf.

Su:2010:TRE

[116] Francis Edward Su. Teaching research: Encouraging discoveries. American Mathematical Monthly, 117(9):759–769, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X521634.pdf.

Boas:2010:JJM

[117] Harold P. Boas. Julius and Julia: Mastering the art of the Schwarz Lemma. American Mathematical Monthly, 117(9):770–785, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X521643.pdf.

Banakh:2010:FSL

[118] Taras Banakh, Oleg Verbitsky, and Yaroslav Vorobets. Fermat’s spiral and the line between Yin and Yang. American Mathematical Monthly, 117(9):786–800, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X521652.pdf.

Hrbacek:2010:AUN

[119] Karel Hrbacek, Olivier Lessmann, and Richard O’Donovan. Analysis with ultrasmall numbers. American Mathematical Monthly, 117(9):801– 816, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X521661.pdf.

Lenstra:2010:ICM

[120] Hendrik Lenstra. Irreducible cubics modulo five. American Mathematical Monthly, 117(9):817–821, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X521670.pdf.

Michalek:2010:SPC

[121] Mateusz Michałek. A short proof of combinatorial Nullstellensatz. American Mathematical Monthly, 117(9):821–823, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/pdfplus/10.4169/000298910X521689.pdf.

Creswell:2010:CBS

[122] Sam H. Creswell. A continuous bijection from ℓ^{2} onto a subset of ℓ^{2} whose inverse is everywhere discontinuous. American Mathematical Monthly, 117(9):823–828, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/000298910X521698.pdf.

Anghel:2010:QCZ

[123] Nicolae Anghel. On a question of Connes and Zagier regarding parallelograms inscribed in a pair of confocal ellipses. American Mathematical Monthly, 117(9):828–831, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/000298910X521706.pdf.

Walk:2010:SPE

[124] Stephen M. Walk. A streamlined proof of an essential calculus fact. American Mathematical Monthly, 117(9):832–833, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdfplus/10.4169/

000298910X521715.pdf. We present a short, classroom-friendly proof of a standard theorem of calculus: continuity on an interval implies Riemann integrability on that interval.

Edger:2010:PSb

[125] Gerald A. Edger, Doug Hensley, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 117(9):834–843, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/000298910X521724.pdf.

Nunemacher:2010:BRBg

[126] Jeffrey Nunemacher and Rebecca F. Goldin. Book review: [untitled]. American Mathematical Monthly, 117(9):844–847, November 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdfplus/10.4169/
000298910X521733.pdf.

Anonymous:2010:BMj

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amermathmont.117.9.BM.pdf.

Anonymous:2011:FMa

[128] Anonymous. Front matter. American Mathematical Monthly, 118(1): 1–2, January 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.01.fm.pdf.

Cox:2011:WEP

[129] David A. Cox. Why Eisenstein proved the Eisenstein criterion and why Schönemann discovered it first. American Mathematical Monthly, 118(1): 3–21, January 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.01.003.pdf.

Johnson:2011:HSC

[130] Jennifer M. Johnson and János Kollár. How small can a polynomial be near infinity? American Mathematical Monthly, 118(1):22– 40, January 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.01.022.pdf.

Ivanov:2011:AJI

[131] Nikolai V. Ivanov. Arnol’d, the Jacobi identity, and orthocenters. American Mathematical Monthly, 118(1):41–65, January 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.01.041.pdf.

Pippenger:2011:TER

[132] Nicholas Pippenger. Two extensions of results of Archimedes. American Mathematical Monthly, 118(1):66–71, January 2011. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.118.
01.066.pdf.

Pollack:2011:PRG

[133] Paul Pollack. On polynomial rings with a Goldbach property. American Mathematical Monthly, 118(1):71–77, January 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.01.071.pdf.

Villani:2011:MFG

[134] Alfonso Villani. Measurable functions with a given set of integrability exponents. American Mathematical Monthly, 118(1):77–82, January 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.01.077.pdf.

Toda:2011:ORM

[135] Alexis Akira Toda. Operator reverse monotonicity of the inverse. American Mathematical Monthly, 118(1):82–83, January 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.01.082.pdf.

Anonymous:2011:PSa

[136] Anonymous. Problems and solutions. American Mathematical Monthly, 118(1):84–91, January 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.01.084.pdf.

Warfield:2011:BRB

[137] Virginia McShane Warfield. Book review: [untitled]. American Mathematical Monthly, 118(1):92–95, January 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.01.092.pdf.

Anonymous:2011:BMa

[138] Anonymous. Back matter. American Mathematical Monthly, 118(1): 96, January 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.01.bm.pdf.

Anonymous:2011:FMb

[139] Anonymous. Front matter. American Mathematical Monthly, 118 (10):861–862, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.10.fm.pdf.

Gerlach:2011:SFR

[140] Henryk Gerlach and Heiko von der Mosel. On sphere-filling ropes. American Mathematical Monthly, 118(10):863–876, December 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.10.863.pdf.

Kos:2011:GPS

[141] Géza Kós. On the Grasshopper Problem with signed jumps. American Mathematical Monthly, 118(10):877–886, December 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.10.877.pdf.

Falcone:2011:KTF

[142] Giovanni Falcone and Marco Pavone. Kirkman’s tetrahedron and the Fifteen Schoolgirl Problem. American Mathematical Monthly, 118(10):887– 900, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.10.887.pdf.

Baginski:2011:FAI

[143] Paul Baginski and Scott T. Chapman. Factorizations of algebraic integers, block monoids, and additive number theory. American Mathematical Monthly, 118(10):901–920, December 2011. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.118.
10.901.pdf.

Ardal:2011:COR

[144] Hayri Ardal, Tom Brown, and Veselin Jungić. Chaotic orderings of the rationals and reals. American Mathematical Monthly, 118(10):921–925, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.10.921.pdf.

Cerf:2011:SPC

[145] Raphaël Cerf and Pierre Petit. A short proof of Cramér’s Theorem in R. American Mathematical Monthly, 118(10):925–931, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.10.925.pdf.

Kalantari:2011:GMP

[146] Bahman Kalantari. A geometric modulus principle for polynomials. American Mathematical Monthly, 118(10):931–935, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.10.931.pdf.

Anonymous:2011:PSb

[147] Anonymous. Problems and solutions. American Mathematical Monthly, 118(10):936–945, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.10.936.pdf.

Odlyzko:2011:BRE

[148] Andrew Odlyzko. Book review: Experimental Mathematics in Action, by D. H. Bailey, J. M. Borwein, N. J. Calkin, R. Girgensohn, D. R. Luke, and V. H. Moll. A K Peters, Wellesley, MA, 2007. xii + 322 pp., ISBN 978-1-56881-271-7. $65. American Mathematical Monthly, 118(10):946– 951, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL https://www.jstor.org/stable/10.
4169/amer.math.monthly.118.10.946;   https://www.jstor.org/
stable/pdfplus/10.4169/amer.math.monthly.118.10.946.pdf.

Velleman:2011:BRB

[149] Daniel J. Velleman. Book review: Endnotes. American Mathematical Monthly, 118(10):952–953, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.10.952.pdf.

Anonymous:2011:IV

[150] Anonymous. Index to volume 118. American Mathematical Monthly, 118(10):954–964, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.10.954.pdf.

Anonymous:2011:BMb

[151] Anonymous. Back matter. American Mathematical Monthly, 118(10): ??, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.10.bm.pdf.

Anonymous:2011:FMc

[152] Anonymous. Front matter. American Mathematical Monthly, 118(2):97– 98, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.02.fm.pdf.

Bressoud:2011:HRT

[153] David M. Bressoud. Historical reflections on teaching the Fundamental Theorem of Integral Calculus. American Mathematical Monthly, 118(2):99–115, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.02.099.pdf.

Vakil:2011:MD

[154] Ravi Vakil. The mathematics of doodling. American Mathematical Monthly, 118(2):116–129, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.02.116.pdf.

Short:2011:FCC

[155] Ian Short. Ford circles, continued fractions, and rational approximation. American Mathematical Monthly, 118(2):130–135, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.02.130.pdf.

Muller:2011:HAN

[156] Markus Müller and Dierk Schleicher. How to add a noninteger number of terms: From axioms to new identities. American Mathematical Monthly, 118(2):136–152, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.02.136.pdf.

Jones:2011:WDA

[157] Lenny Jones. When does appending the same digit repeatedly on the right of a positive integer generate a sequence of composite integers? American Mathematical Monthly, 118(2):153–160, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.02.153.pdf.

Pollack:2011:DTC

[158] Paul Pollack. On Dickson’s theorem concerning odd perfect numbers. American Mathematical Monthly, 118(2):161–164, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.02.161.pdf.

Ackerman:2011:ESA

[159] Eyal Ackerman, Rom Pinchasi, Ludmila Scharf, and Marc Scherfenberg. Every simple arrangement of n lines contains an inducing simple ngon. American Mathematical Monthly, 118(2):164–167, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.02.164.pdf.

Gamez-Merino:2011:CCR

[160] José L. Gámez-Merino, Gustavo A. Muñoz-Fernández, and Juan B. Seoane-Sepúlveda. A characterization of continuity revisited. American Mathematical Monthly, 118(2):167–170, February 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.02.167.pdf.

Northshield:2011:SPG

[161] Sam Northshield. A short proof and generalization of Lagrange’s theorem on continued fractions. American Mathematical Monthly, 118(2):171– 175, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.02.171.pdf.

Chang:2011:GPP

[162] Guisong Chang and Chen Xu. Generalization and probabilistic proof of a combinatorial identity. American Mathematical Monthly, 118(2):175– 177, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.02.175.pdf.

Anonymous:2011:PSc

[163] Anonymous. Problems and solutions. American Mathematical Monthly, 118(2):178–185, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.02.178.pdf.

Nunemacher:2011:BRB

[164] Jeffrey Nunemacher. Book review: [untitled]. American Mathematical Monthly, 118(2):186–189, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.02.186.pdf.

Anonymous:2011:BMc

[165] Anonymous. Back matter. American Mathematical Monthly, 118(2):190– 192, February 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.02.bm.pdf.

Anonymous:2011:FMd

[166] Anonymous. Front matter. American Mathematical Monthly, 118 (3):193–194, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.03.fm.pdf.

Faires:2011:YGG

[167] Barbara Faires. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2011 to Joseph A. Gallian for Distinguished Service to Mathematics. American Mathematical Monthly, 118(3):195–197, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.03.195.pdf.

Gouvea:2011:WCS

[168] Fernando Q. Gouvêa. Was Cantor surprised? American Mathematical Monthly, 118(3):198–209, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.03.198.pdf.

Weintraub:2011:LWL

[169] Steven H. Weintraub. On Legendre’s work on the law of quadratic reciprocity. American Mathematical Monthly, 118(3):210–216, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.03.210.pdf.

Ryba:2011:EPT

[170] Alexander Ryba and Joseph Stern. Equimodular polynomials and the tritangency theorems of Euler, Feuerbach, and Guinand. American Mathematical Monthly, 118(3):217–228, March 2011. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.118.
03.217.pdf.

Burns:2011:STN

[171] Keith Burns and Boris Hasselblatt. The Sharkovsky Theorem: a natural direct proof. American Mathematical Monthly, 118(3):229–244, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.03.229.pdf.

Hubbard:2011:FLD

[172] John H. Hubbard and Benjamin E. Lundell. A first look at differential algebra. American Mathematical Monthly, 118(3):245–261, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.03.245.pdf.

Keady:2011:LBF

[173] Grant Keady. Lines of best fit for the zeros and for the critical points of a polynomial. American Mathematical Monthly, 118(3):262– 264, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.03.262.pdf.

Kurz:2011:RMG

[174] Sascha Kurz and Rom Pinchasi. Regular matchstick graphs. American Mathematical Monthly, 118(3):264–267, March 2011. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.118.
03.264.pdf.

Chlebus:2011:RSI

[175] Edward Chlebus. A recursive scheme for improving the original rate of convergence to the Euler–Mascheroni constant. American Mathematical Monthly, 118(3):268–274, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.03.268.pdf.

Anonymous:2011:PSd

[176] Anonymous. Problems and solutions. American Mathematical Monthly, 118(3):275–282, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.03.275.pdf.

Leary:2011:BRB

[177] Christopher C. Leary. Book review: [untitled]. American Mathematical Monthly, 118(3):283–286, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.03.283.pdf.

Anonymous:2011:BMd

[178] Anonymous. Back matter. American Mathematical Monthly, 118(3):287– 288, March 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.03.bm.pdf.

Anonymous:2011:FMe

[179] Anonymous. Front matter. American Mathematical Monthly, 118(4): 289–290, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.04.fm.pdf.

Sarnak:2011:IAP

[180] Peter Sarnak. Integral Apollonian packings. American Mathematical Monthly, 118(4):291–306, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.118.04.291.pdf.

Hull:2011:SCC

[181] Thomas C. Hull. Solving cubics with creases: The work of Beloch and Lill. American Mathematical Monthly, 118(4):307–315, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.04.307.pdf.

Chan:2011:CAJ

[182] Heng Huat Chan, Ling Long, and YiFan Yang. A cubic analogue of the Jacobsthal Identity. American Mathematical Monthly, 118(4):316– 326, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.04.316.pdf.

Hilden:2011:ASS

[183] H. M. Hilden, J. M. Montesinos, D. M. Tejada, and M. M. Toro. Artifacts for stamping symmetric designs. American Mathematical Monthly, 118(4):327–343, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.04.327.pdf.

Paradis:2011:NSF

[184] Jaume Paradís, Pelegrí Viader, and Lluís Bibiloni. A new singular function. American Mathematical Monthly, 118(4):344–354, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.04.344.pdf.

Cooke:2011:RET

[185] Roger Cooke. A remark on Euclid’s Theorem on the infinitude of the primes. American Mathematical Monthly, 118(4):355–358, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.04.355.pdf.

Rickards:2011:WPC

[186] James Rickards. When is a polynomial a composition of other polynomials? American Mathematical Monthly, 118(4):358–363, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.04.358.pdf.

Moree:2011:TMF

[187] Pieter Moree. A top hat for Moser’s four mathemagical rabbits. American Mathematical Monthly, 118(4):364–370, April 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.04.364.pdf.

Anonymous:2011:PSe

[188] Anonymous. Problems and solutions. American Mathematical Monthly, 118(4):371–378, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.04.371.pdf.

Colley:2011:BRB

[189] Susan Jane Colley. Book review: [untitled]. American Mathematical Monthly, 118(4):379–382, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.118.04.379.pdf.

Anonymous:2011:BMe

[190] Anonymous. Back matter. American Mathematical Monthly, 118(4): 383–384, April 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.04.bm.pdf.

Anonymous:2011:FMf

[191] Anonymous. Front matter. American Mathematical Monthly, 118(5): 385–386, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.05.fm.pdf.

Watkins:2011:FA

[192] David S. Watkins. Francis’s algorithm. American Mathematical Monthly, 118(5):387–403, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL https://www.jstor.org/stable/info/
10.4169/amer.math.monthly.118.05.387.

Range:2011:WLN

[193] R. Michael Range. Where are limits needed in calculus? American Mathematical Monthly, 118(5):404–417, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.05.404.pdf.

Vatter:2011:MIS

[194] Vincent Vatter. Maximal independent sets and separating covers. American Mathematical Monthly, 118(5):418–423, May 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.05.418.pdf.

Skopenkov:2011:PCB

[195] M. Skopenkov. Packing a cake into a box. American Mathematical Monthly, 118(5):424–433, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.118.05.424.pdf.

Tucker:2011:WEM

[196] Alan Tucker. What every mathematician should know about standardsbased tests. American Mathematical Monthly, 118(5):434–447, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.05.434.pdf.

Robinson:2011:CTS

[197] Geoffrey R. Robinson. Cauchy’s Theorem and Sylow’s Theorem from the cyclic case. American Mathematical Monthly, 118(5):448–449, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.05.448.pdf.

Sokal:2011:RSE

[198] Alan D. Sokal. A really simple elementary proof of the uniform boundedness theorem. American Mathematical Monthly, 118(5):450– 452, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.05.450.pdf.

Deaconescu:2011:GAG

[199] Marian Deaconescu and Gary L. Walls. On the group of automorphisms of a group. American Mathematical Monthly, 118(5):452–455, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.05.452.pdf.

Nica:2011:USI

[200] Bogdan Nica. The unreasonable slightness of E_{2} over imaginary quadratic rings. American Mathematical Monthly, 118(5):455–462, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.05.455.pdf. See corrigendum [962].

Anonymous:2011:PSf

[201] Anonymous. Problems and solutions. American Mathematical Monthly, 118(5):463–470, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.05.463.pdf.

Kidwell:2011:BRB

[202] Peggy Aldrich Kidwell. Book review: [untitled]. American Mathematical Monthly, 118(5):471–474, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.05.471.pdf.

Anonymous:2011:BMf

[203] Anonymous. Back matter. American Mathematical Monthly, 118(5): 475–476, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.05.bm.pdf.

Anonymous:2011:FMg

[204] Anonymous. Front matter. American Mathematical Monthly, 118(6): 477–478, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.06.fm.pdf.

Kuczmarski:2011:RWR

[205] Fred Kuczmarski. Roads and wheels, roulettes and pedals. American Mathematical Monthly, 118(6):479–496, June 2011. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.118.
06.479.pdf.

Griffiths:2011:GSZ

[206] Martin Griffiths. The golden string, Zeckendorf representations, and the sum of a series. American Mathematical Monthly, 118(6):497– 507, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.06.497.pdf.

Jones:2011:ASS

[207] Rafe Jones. Achievement sets of sequences. American Mathematical Monthly, 118(6):508–521, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.118.06.508.pdf.

Gorkin:2011:PEM

[208] Pamela Gorkin and Elizabeth Skubak. Polynomials, ellipses, and matrices: Two questions, one answer. American Mathematical Monthly, 118(6):522–533, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.06.522.pdf.

Wotzlaw:2011:LCP

[209] Ronald F. Wotzlaw and Günter M. Ziegler. A lost counterexample and a problem on illuminated polytopes. American Mathematical Monthly, 118(6):534–543, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.06.534.pdf.

Pach:2011:CCM

[210] János Pach and Ethan Sterling. Conway’s Conjecture for monotone thrackles. American Mathematical Monthly, 118(6):544–548, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.06.544.pdf.

MacMillan:2011:PPS

[211] Kieren MacMillan and Jonathan Sondow. Proofs of power sum and binomial coefficient congruences via Pascal’s Identity. American Mathematical Monthly, 118(6):549–551, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.06.549.pdf.

Emmons:2011:MCD

[212] Caleb Emmons. Meromorphic continuation of Dirichlet series via derivations. American Mathematical Monthly, 118(6):551–556, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.06.551.pdf.

Anonymous:2011:PSg

[213] Anonymous. Problems and solutions. American Mathematical Monthly, 118(6):557–564, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.118.06.557.

Hartshorne:2011:BRB

[214] Robin Hartshorne. Book review: [untitled]. American Mathematical Monthly, 118(6):565–568, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.118.06.565.pdf.

Anonymous:2011:BMg

[215] Anonymous. Back matter. American Mathematical Monthly, 118(6): ??, June 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.06.bm.pdf.

Anonymous:2011:FMh

[216] Anonymous. Front matter. American Mathematical Monthly, 118 (7):569–570, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.07.fm.pdf.

Ross:2011:BLG

[217] Kenneth A. Ross. Benford’s Law, a growth industry. American Mathematical Monthly, 118(7):571–583, August/September 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169;   https://
www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.118.
07.571.pdf; http://www.maa.org/history/presidents/ross.html.

Everest:2011:RMD

[218] Graham Everest and Thomas Ward. A repulsion motif in Diophantine equations. American Mathematical Monthly, 118(7):584–598, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.07.584.pdf.

Torchinsky:2011:FTW

[219] Alberto Torchinsky. The Fourier transform and the wave equation. American Mathematical Monthly, 118(7):599–609, August 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.07.599.pdf.

Aitken:2011:CHP

[220] W. Aitken and F. Lemmermeyer. Counterexamples to the Hasse Principle. American Mathematical Monthly, 118(7):610–628, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.07.610.pdf.

Eslami:2011:EEM

[221] Peyman Eslami and Pawel Góra. On eventually expanding maps of the interval. American Mathematical Monthly, 118(7):629–635, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.07.629.pdf.

Ben-El-Mechaiekh:2011:SPN

[222] Hichem Ben-El-Mechaiekh and Robert W. Dimand. A simpler proof of the von Neumann minimax theorem. American Mathematical Monthly, 118(7):636–641, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.07.636.pdf.

Pace:2011:PPZ

[223] Luigi Pace. Probabilistically proving that ζ(2) = π^{2}/6. American Mathematical Monthly, 118(7):641–643, August 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.07.641.pdf.

Murty:2011:PR

[224] M. Ram Murty and V. Kumar Murty. On a problem of Ruderman. American Mathematical Monthly, 118(7):644–650, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.07.644.pdf.

Beardon:2011:CAD

[225] Alan F. Beardon. Cyclic absolute differences of integers. American Mathematical Monthly, 118(7):650–652, August 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.07.650.pdf.

Anonymous:2011:PSh

[226] Anonymous. Problems and solutions. American Mathematical Monthly, 118(7):653–660, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.07.653.pdf.

Winkler:2011:BRB

[227] Peter Winkler. Book review: [untitled]. American Mathematical Monthly, 118(7):661–664, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.07.661.pdf.

Anonymous:2011:BMh

[228] Anonymous. Back matter. American Mathematical Monthly, 118(7): ??, August 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.07.bm.pdf.

Anonymous:2011:FMi

[229] Anonymous. Front matter. American Mathematical Monthly, 118 (8):665–666, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.08.fm.pdf.

Ozaydin:2011:BTL

[230] Murad Őzaydın and Charlotte Simmons. The birthday trick with lies. American Mathematical Monthly, 118(8):667–679, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.08.667.pdf.

Gordon:2011:SPT

[231] Russell A. Gordon. Semiperimeter and Pythagorean triples. American Mathematical Monthly, 118(8):680–692, October 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.08.680.pdf.

Kharazishvili:2011:MPV

[232] A. B. Kharazishvili. Measurability properties of Vitali sets. American Mathematical Monthly, 118(8):693–703, October 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.08.693.pdf.

Garcia:2011:QSD

[233] Stephan Ramon Garcia, Vincent Selhorst-Jones, Daniel E. Poore, and Noah Simon. Quotient sets and Diophantine equations. American Mathematical Monthly, 118(8):704–711, October 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.08.704.pdf.

Giblin:2011:RMR

[234] P. J. Giblin and J. P. Warder. Reconstruction from medial representations. American Mathematical Monthly, 118(8):712–725, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.08.712.pdf.

Klosinski:2011:SFW

[235] Leonard F. Klosinski, Gerald L. Alexanderson, Loren C. Larson, and Mark Krusemeyer. The Seventy-First William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 118(8):726– 733, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.08.726.pdf.

Oliver:2011:PP

[236] R. K. Oliver. On the parity of a permutation. American Mathematical Monthly, 118(8):734–735, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.08.734.pdf.

Kager:2011:HTT

[237] Wouter Kager. The hitting time theorem revisited. American Mathematical Monthly, 118(8):735–737, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.08.735.pdf.

Thangadurai:2011:LPC

[238] R. Thangadurai and A. Vatwani. The least prime congruent to one modulo n. American Mathematical Monthly, 118(8):737–742, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.08.737.pdf.

Cao:2011:BTT

[239] Penghao Cao and Liping Yuan. A Blichfeldt-type theorem for Hpoints. American Mathematical Monthly, 118(8):743–746, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.08.743.pdf.

Anonymous:2011:PSi

[240] Anonymous. Problems and solutions. American Mathematical Monthly, 118(8):747–754, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.08.747.pdf.

Askey:2011:BRB

[241] Richard Askey. Book review: [untitled]. American Mathematical Monthly, 118(8):755–762, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.08.755.pdf.

Anonymous:2011:BMi

[242] Anonymous. Back matter. American Mathematical Monthly, 118(8): ??, October 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.08.bm.pdf.

Anonymous:2011:FMj

[243] Anonymous. Front matter. American Mathematical Monthly, 118 (9):765–766, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.09.fm.pdf.

Berndt:2011:CAM

[244] Bruce C. Berndt. The chief accountant and mathematical friend of Ramanujan — S. Narayana Aiyar. American Mathematical Monthly, 118(9):767–776, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.09.767.pdf.

Haxell:2011:FC

[245] P. Haxell. On forming committees. American Mathematical Monthly, 118(9):777–788, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.09.777.pdf.

Apostol:2011:CDC

[246] Tom M. Apostol and Mamikon A. Mnatsakanian. Complete dissections: Converting regions and their boundaries. American Mathematical Monthly, 118(9):789–798, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.09.789.pdf.

Drago:2011:BDF

[247] Giacomo Drago, Pier Domenico Lamberti, and Paolo Toni. A “bouquet” of discontinuous functions for beginners in mathematical analysis. American Mathematical Monthly, 118(9):799–811, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.09.799.pdf.

Cosgrave:2011:IGF

[248] John B. Cosgrave and Karl Dilcher. An introduction to Gauss factorials. American Mathematical Monthly, 118(9):812–829, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.09.812.pdf.

Glasby:2011:ERL

[249] S. P. Glasby. Enumerating the rationals from left to right. American Mathematical Monthly, 118(9):830–835, November 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.09.830.pdf.

Litvinov:2011:BST

[250] Semyon N. Litvinov. On the Banach–Steinhaus theorem for topological groups. American Mathematical Monthly, 118(9):836–838, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.09.836.pdf.

Hirschhorn:2011:RMB

[251] Michael D. Hirschhorn. Ramanujan’s “Most Beautiful Identity”. American Mathematical Monthly, 118(9):839–845, November 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.118.09.839.pdf.

Anonymous:2011:PSj

[252] Anonymous. Problems and solutions. American Mathematical Monthly, 118(9):846–853, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.118.09.846.pdf.

Richey:2011:BRB

[253] Matthew Richey. Book review: [untitled]. American Mathematical Monthly, 118(9):854–858, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.118.09.854.pdf.

Anonymous:2011:BMj

[254] Anonymous. Back matter. American Mathematical Monthly, 118(9):??, November 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.09.bm.pdf.

Anonymous:2012:FMa

[255] Anonymous. Front matter. American Mathematical Monthly, 119(1): 1–2, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.01.fm.pdf.

Chapman:2012:LE

[256] Scott Chapman. A letter from the Editor. American Mathematical Monthly, 119(1):3, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.01.003.pdf.

Brinkman:2012:IH

[257] Daniel Brinkman and Peter J. Olver. Invariant histograms. American Mathematical Monthly, 119(1):4–24, January 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
01.004.pdf.

Bauer:2012:ZDN

[258] Thomas Bauer, Mirel Caibăr, and Gary Kennedy. Zariski decomposition: a new (old) chapter of linear algebra. American Mathematical Monthly, 119(1):25–41, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.01.025.pdf.

Kalman:2012:AWS

[259] Dan Kalman and Mark McKinzie. Another way to sum a series: Generating functions, Euler, and the dilog function. American Mathematical Monthly, 119(1):42–51, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.01.042.pdf.

Efthimiou:2012:CPC

[260] Costas J. Efthimiou. A class of periodic continued radicals. American Mathematical Monthly, 119(1):52–58, January 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
01.052.pdf.

Laosinchai:2012:GIP

[261] Parames Laosinchai and Bhinyo Panijpan. A geometric interpretation of Pascal’s formula for sums of powers of integers. American Mathematical Monthly, 119(1):58–64, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.01.058.pdf.

Clark:2012:CNL

[262] Pete L. Clark. Covering numbers in linear algebra. American Mathematical Monthly, 119(1):65–67, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.01.065.pdf.

Anonymous:2012:PSa

[263] Anonymous. Problems and solutions. American Mathematical Monthly, 119(1):68–75, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.01.068.pdf.

Durfee:2012:BRB

[264] Alan Durfee. Book review: [untitled]. American Mathematical Monthly, 119(1):76–82, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.01.076.pdf.

Anonymous:2012:BMa

[265] Anonymous. Back matter. American Mathematical Monthly, 119(1):83– 84, January 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.01.bm.pdf.

Anonymous:2012:FMb

[266] Anonymous. Front matter. American Mathematical Monthly, 119(2):85– 86, February 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.02.fm.pdf.

Mordukhovich:2012:SGH

[267] Boris S. Mordukhovich, Nguyen Mau Nam, and Juan Salinas, Jr. Solving a generalized heron problem by means of convex analysis. American Mathematical Monthly, 119(2):87–99, February 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.02.087.pdf.

Vandervelde:2012:JSM

[268] Sam Vandervelde. Jacobi sum matrices. American Mathematical Monthly, 119(2):100–114, February 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.02.100.pdf.

Bindner:2012:AS

[269] Donald J. Bindner and Martin Erickson. Alcuin’s sequence. American Mathematical Monthly, 119(2):115–121, February 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.02.115.pdf.

Polster:2012:CCH

[270] Burkard Polster, Marty Ross, and David Treeby. A case of continuous hangover. American Mathematical Monthly, 119(2):122–139, February 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.02.122.pdf.

Herron:2012:RMD

[271] David A. Herron. Riemann maps and diameter distance. American Mathematical Monthly, 119(2):140–147, February 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.02.140.pdf.

Mortici:2012:PSA

[272] Cristinel Mortici. A power series approach to some inequalities. American Mathematical Monthly, 119(2):147–151, February 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.02.147.pdf.

Rosen:2012:CMF

[273] Julian Rosen, Zachary Scherr, Benjamin Weiss, and Michael E. Zieve. Chebyshev mappings of finite fields. American Mathematical Monthly, 119(2):151–155, February 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.02.151.pdf.

Pak:2012:CWT

[274] Igor Pak and Rom Pinchasi. Collapsing walls theorem. American Mathematical Monthly, 119(2):156–160, February 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.02.156.pdf.

Anonymous:2012:PSb

[275] Anonymous. Problems and solutions. American Mathematical Monthly, 119(2):161–168, February 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.02.161.pdf.

Ferreiros:2012:BRB

[276] José Ferreirós. Book review: [untitled]. American Mathematical Monthly, 119(2):169–176, February 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.02.169.pdf.

Anonymous:2012:BMb

[277] Anonymous. Back matter. American Mathematical Monthly, 119(2): ??, February 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.02.bm.pdf.

Anonymous:2012:FMc

[278] Anonymous. Front matter. American Mathematical Monthly, 119 (3):177–178, March 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.03.fm.pdf.

Watkins:2012:YGG

[279] Ann E. Watkins. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2012 to John Ewing for Distinguished Service to Mathematics. American Mathematical Monthly, 119(3):179–182, March 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.03.179.pdf.

Sidman:2012:IAG

[280] Jessica Sidman. An introduction to algebraic geometry: Polygons, parameterizations, and equations. American Mathematical Monthly, 119(3):183–198, March 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.03.183.pdf.

Daners:2012:MSP

[281] Daniel Daners. The Mercator and stereographic projections, and many in between. American Mathematical Monthly, 119(3):199–210, March 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.03.199.pdf.

Nievergelt:2012:RGD

[282] Yves Nievergelt. Real and generic data without unconstrained bestfitting Verhulst curves and sufficient conditions for median Mitscherlich and Verhulst curves to exist. American Mathematical Monthly, 119(3):211–234, March 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.03.211.pdf.

Waterhouse:2012:SRH

[283] William C. Waterhouse. Square root as a homomorphism. American Mathematical Monthly, 119(3):235–239, March 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
03.235.pdf.

Cohen:2012:SMI

[284] Michael P. Cohen. Sample means of independent standard Cauchy random variables are standard Cauchy: a new approach. American Mathematical Monthly, 119(3):240–244, March 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
03.240.pdf.

Reem:2012:CAU

[285] Daniel Reem. Character analysis using Fourier series. American Mathematical Monthly, 119(3):245–246, March 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
03.245.pdf.

Anonymous:2012:PSc

[286] Anonymous. Problems and solutions. American Mathematical Monthly, 119(3):247–254, March 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.03.247.pdf.

Alexanderson:2012:BRBa

[287] Gerald L. Alexanderson. Book review: Loving and Hating Mathematics: Challenging the Myths of Mathematical Life, by Reuben Hersh and Vera John-Steiner. Princeton University Press, Princeton, NJ, 2011, xii + 416 pp. ISBN 978-0-691-14247-0, $29.95. American Mathematical Monthly, 119(3):255–259, March 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.03.255.pdf.

Anonymous:2012:BMc

[288] Anonymous. Back matter. American Mathematical Monthly, 119(3): 260, March 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.03.bm.pdf.

Anonymous:2012:FMd

[289] Anonymous. Front matter. American Mathematical Monthly, 119(4): 261–262, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.04.fm.pdf.

Hall:2012:SGU

[290] Rachel Wells Hall and Dmitri Tymoczko. Submajorization and the geometry of unordered collections. American Mathematical Monthly, 119(4):

263–283, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.04.263.pdf.

McAdam:2012:UMH

[291] Stephen McAdam. Unknowable matters: How Nature’s speed limit on communication relates to quantum physics. American Mathematical Monthly, 119(4):284–299, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.04.284.pdf.

Cox:2012:DP

[292] Barry Cox and Stan Wagon. Drilling for polygons. American Mathematical Monthly, 119(4):300–312, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.04.300.pdf.

Hicks:2012:SRP

[293] Kenneth Hicks, Xiang dong Hou, and Gary L. Mullen. Sums of reciprocals of polynomials over finite fields. American Mathematical Monthly, 119(4):313–317, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.04.313.pdf.

Lakshtanov:2012:FCG

[294] Evgeny Lakshtanov and Vera Roshchina. On finiteness in the card game of War. American Mathematical Monthly, 119(4):318–323, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.04.318.pdf.

Cvijovic:2012:LRR

[295] Djurdje Cvijović and Hari M. Srivastava. Limit representations of Riemann’s zeta function. American Mathematical Monthly, 119(4):324– 330, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.04.324.pdf.

Elsholtz:2012:PDT

[296] Christian Elsholtz. Prime divisors of thin sequences. American Mathematical Monthly, 119(4):331–333, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.04.331.pdf.

Leonard:2012:FRO

[297] I. E. Leonard and A. C. F. Liu. A familiar recurrence occurs again. American Mathematical Monthly, 119(4):333–336, April 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.04.333.pdf.

Fokkink:2012:TCG

[298] Robbert Fokkink. Tossing coins to guess a secret number. American Mathematical Monthly, 119(4):337–339, April 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
04.337.pdf.

Chen:2012:GCH

[299] Lizhou Chen. A generalization of the Cayley–Hamilton theorem. American Mathematical Monthly, 119(4):340–342, April 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.04.340.pdf.

McCutcheon:2012:MPQ

[300] R. McCutcheon. Monochromatic permutation quadruples — a Schur thing in S_{n}. American Mathematical Monthly, 119(4):342–343, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdfplus/10.4169/
amer.math.monthly.119.04.342.pdf.

Anonymous:2012:PSd

[301] Anonymous. Problems and solutions. American Mathematical Monthly, 119(4):344–351, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.04.344.pdf.

Shulman:2012:BRB

[302] Bonnie Shulman. Book review: Population Games and Evolutionary Dynamics, by William H. Sandholm. MIT Press, Cambridge, MA, 2011, xxv + 589 pp., ISBN 978-0-262-19587-4, $65. American Mathematical Monthly, 119(4):352–356, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.04.352.pdf.

Anonymous:2012:BMd

[303] Anonymous. Back matter. American Mathematical Monthly, 119(4): ??, April 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.04.bm.pdf.

Anonymous:2012:FMe

[304] Anonymous. Front matter. American Mathematical Monthly, 119(5): 357–358, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.05.fm.pdf.

Gilbert:2012:LMM

[305] Tony Gilbert. Lost mail: The missing envelope in the problem of the minimal surface of revolution. American Mathematical Monthly, 119(5): 359–372, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.05.359.pdf.

Beukers:2012:PVD

[306] Frits Beukers, Florian Luca, and Frans Oort. Power values of divisor sums. American Mathematical Monthly, 119(5):373–380, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.05.373.pdf.

Boualem:2012:WAN

[307] Hassan Boualem and Robert Brouzet. On what is the almost-near principle. American Mathematical Monthly, 119(5):381–397, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.05.381.pdf.

Tabachnikov:2012:ETP

[308] Serge Tabachnikov. The (un)equal tangents problem. American Mathematical Monthly, 119(5):398–405, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.05.398.pdf.

Bryant:2012:VMC

[309] Lance Bryant, James Hamblin, and Lenny Jones. A variation on the money-changing problem. American Mathematical Monthly, 119(5):406– 414, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.05.406.pdf.

Callan:2012:FLC

[310] David Callan. Flexagons lead to a Catalan number identity. American Mathematical Monthly, 119(5):415–419, May 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
05.415.pdf.

Arthan:2012:ACE

[311] R. D. Arthan. Aronszajn’s criterion for Euclidean space. American Mathematical Monthly, 119(5):419–422, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.05.419.pdf.

Borzellino:2012:WTP

[312] Joseph E. Borzellino and Morgan Sherman. When is a trigonometric polynomial not a trigonometric polynomial? American Mathematical Monthly, 119(5):422–425, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.119.05.422.pdf.

Anonymous:2012:PSe

[313] Anonymous. Problems and solutions. American Mathematical Monthly, 119(5):426–433, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.05.426.pdf.

Walsh:2012:BRB

[314] James A. Walsh. Book review: Randomness and Recurrence in Dynamical Systems, by Rodney Nillsen. The Carus Mathematical Monographs Number 31, Mathematical Association of America, Washington, DC, 2010, xviii + 357 pp., ISBN 978-0-88385-043-5, $52.95. American Mathematical Monthly, 119(5):434–438, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.05.434.pdf.

Anonymous:2012:BMe

[315] Anonymous. Back matter. American Mathematical Monthly, 119(5): 439–440, May 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.05.bm.pdf.

Anonymous:2012:FMf

[316] Anonymous. Front matter. American Mathematical Monthly, 119(6): 441–442, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.06.fm.pdf.

Koukoulopoulos:2012:ASA

[317] Dimitris Koukoulopoulos and Johann Thiel. Arrangements of stars on the American flag. American Mathematical Monthly, 119(6):443– 450, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.06.443.pdf.

Kuczmarski:2012:RSS

[318] Fred Kuczmarski. Rolling sinusoidal spirals. American Mathematical Monthly, 119(6):451–467, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.119.06.451.pdf.

Mandl:2012:AAG

[319] Robert Mandl, Thomas Pühringer, and Maximilian Thaler. Analytic approach to Galileo’s Theorem on the descent time along two-chord paths in a circle. American Mathematical Monthly, 119(6):468–476, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.06.468.pdf.

Binding:2012:PAA

[320] Paul Binding and Hans Volkmer. A Prüfer angle approach to the periodic Sturm–Liouville problem. American Mathematical Monthly, 119(6):477– 484, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.06.477.pdf.

Mortini:2012:MES

[321] Raymond Mortini and Rudolf Rupp. Mappings into the Euclidean sphere. American Mathematical Monthly, 119(6):485–494, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.06.485.pdf.

Grunbaum:2012:NTR

[322] Branko Grünbaum. Is Napoleon’s Theorem really Napoleon’s Theorem? American Mathematical Monthly, 119(6):495–501, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.06.495.pdf.

Dorfling:2012:PD

[323] Michael J. Dorfling and Johan H. Meyer. On polynomials and divisibility. American Mathematical Monthly, 119(6):502–506, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.06.502.pdf.

Roitman:2012:GCM

[324] Pedro Roitman. Gaussian curvature, mirrors, and maps. American Mathematical Monthly, 119(6):506–509, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.06.506.pdf.

Aw:2012:TNP

[325] Alan J. Aw. The Turán number and probabilistic combinatorics. American Mathematical Monthly, 119(6):510–513, June 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.06.510.pdf.

Wyss:2012:PP

[326] Walter Wyss. Perfect parallelograms. American Mathematical Monthly, 119(6):513–515, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.06.513.pdf.

Wallstrom:2012:EPP

[327] Timothy C. Wallstrom. The equalization probability of the Pólya urn. American Mathematical Monthly, 119(6):516–518, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.06.516.pdf.

Holst:2012:PEI

[328] Lars Holst. A proof of Euler’s infinite product for the sine. American Mathematical Monthly, 119(6):518–521, June 2012. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.119.
06.518.pdf.

Anonymous:2012:PSf

[329] Anonymous. Problems and solutions. American Mathematical Monthly, 119(6):522–529, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.06.522.pdf.

Quinn:2012:BRB

[330] Jennifer J. Quinn. Book review: Combinatorics the Rota Way, by Joseph P. S. Kung, Gian-Carlo Rota, and Catherine H. Yan. Cambridge University Press, New York, 2009, xii + 396 pp., ISBN 978-0-521-73794-4, $40.99. American Mathematical Monthly, 119(6):530–532, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.06.530.pdf.

Anonymous:2012:BMf

[331] Anonymous. Back matter. American Mathematical Monthly, 119(6): ??, June 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.06.bm.pdf.

Anonymous:2012:FMg

[332] Anonymous. Front matter. American Mathematical Monthly, 119 (7):533–534, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.07.fm.pdf.

Borwein:2012:SS

[333] David Borwein, Jonathan M. Borwein, and Armin Straub. A sinc that sank. American Mathematical Monthly, 119(7):535–549, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/1391/
;   https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.07.535.pdf.

Levine:2012:HMM

[334] Lionel Levine and Katherine E. Stange. How to make the most of a shared meal: Plan the last bite first. American Mathematical Monthly, 119(7):550–565, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.07.550.pdf.

Kristensen:2012:NIS

[335] Kai F. Kristensen. A nearest integer solution to the longest run of randomly generated words. American Mathematical Monthly, 119(7):566– 572, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.07.566.pdf.

Lemmermeyer:2012:PAC

[336] F. Lemmermeyer. Parametrization of algebraic curves from a number theorist’s point of view. American Mathematical Monthly, 119(7):573– 583, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.07.573.pdf.

Borwein:2012:EPD

[337] Peter Borwein and Joe Hobart. The extraordinary power of division in straight line programs. American Mathematical Monthly, 119(7):584– 592, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.07.584.pdf.

Bostan:2012:BLM

[338] Alin Bostan and Thierry Combot. A binomial-like matrix equation. American Mathematical Monthly, 119(7):593–597, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.07.593.pdf.

Brown:2012:GGG

[339] Ezra Brown and Marc Chamberland. Generalizing Gauss’s Gem. American Mathematical Monthly, 119(7):597–601, August 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.07.597.pdf.

Kuperberg:2012:NFP

[340] Greg Kuperberg. Norms as a function of p are linearly independent in finite dimensions. American Mathematical Monthly, 119(7):601–603, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.07.601.pdf.

DeGroote:2012:FCM

[341] Cédric De Groote and Mitia Duerinckx. Functions with constant mean on similar countable subsets of R^{2}. American Mathematical Monthly, 119(7):603–605, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.07.603.pdf.

Tkocz:2012:UBS

[342] Tomasz Tkocz. An upper bound for spherical caps. American Mathematical Monthly, 119(7):606–607, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.07.606.pdf.

Anonymous:2012:PSg

[343] Anonymous. Problems and solutions. American Mathematical Monthly, 119(7):608–615, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.07.608.pdf.

Cohen:2012:BRB

[344] Marion Cohen. Book review: Change Is Possible: Stories of Women and Minorities in Mathematics, by Patricia Clark Kenschaft. American Mathematical Society, Providence, RI, 2005, ix + 211 pp., ISBN 0-8218-3748-6, $32.00. American Mathematical Monthly, 119(7):616– 620, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.07.616.pdf.

Anonymous:2012:BMg

[345] Anonymous. Back matter. American Mathematical Monthly, 119(7): ??, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.07.bm.pdf.

Anonymous:2012:FMh

[346] Anonymous. Front matter. American Mathematical Monthly, 119 (8):621–622, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.08.fm.pdf.

Klosinski:2012:SSW

[347] Leonard F. Klosinski, Gerald L. Alexanderson, and Mark Krusemeyer. The Seventy-Second William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 119(8):623–635, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.08.623.pdf.

Ginosar:2012:MPB

[348] Yuval Ginosar and Ilan Yona. A model for pairs of Beatty sequences. American Mathematical Monthly, 119(8):636–645, October 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.08.636.pdf.

Bailey:2012:AIS

[349] David H. Bailey and Jonathan M. Borwein. Ancient Indian square roots: An exercise in forensic paleo-mathematics. American Mathematical Monthly, 119(8):646–657, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://docserver.carma.

newcastle.edu.au/1385/;   https://www.jstor.org/stable/10.4169;
 https://www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.
119.08.646.pdf.

Ebanks:2012:LFG

[350] Bruce Ebanks. Looking for a few good means. American Mathematical Monthly, 119(8):658–669, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.08.658.pdf.

Pollack:2012:FTP

[351] Paul Pollack. Finiteness theorems for perfect numbers and their kin. American Mathematical Monthly, 119(8):670–681, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.08.670.pdf.

Sakurai:2012:VAP

[352] Akira Sakurai. Vector analysis proof of Erdős’ inequality for triangles. American Mathematical Monthly, 119(8):682–684, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.08.682.pdf.

Makarov:2012:CVM

[353] Boris M. Makarov. Calculation of the volume of multidimensional cones and balls using Cavalieri’s Principle. American Mathematical Monthly, 119(8):685–689, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.08.685.pdf.

Angell:2012:ECL

[354] David Angell. Evaluation of certain Legendre symbols. American Mathematical Monthly, 119(8):690–693, October 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.08.690.pdf.

Groves:2012:ECC

[355] Jonathan Groves. An elementary counterexample in the compactopen topology. American Mathematical Monthly, 119(8):693–694, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.08.693.pdf.

Dawson:2012:AP

[356] Robert Dawson. Arithmetic polygons. American Mathematical Monthly, 119(8):695–698, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.08.695.pdf.

Anonymous:2012:PSh

[357] Anonymous. Problems and solutions. American Mathematical Monthly, 119(8):699–706, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.08.699.pdf.

Schwartz:2012:BRB

[358] Mark Schwartz. Book review: Geometry for College Students, by I. Martin Isaacs. American Mathematical Society, The Sally Series, Pure and Applied Undergraduate Texts No. 8, Providence, RI, 2001. vii + 222 pp., ISBN 978-0-8218-4794-7. $62 ($49.60 to AMS members). American Mathematical Monthly, 119(8):707–711, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.08.707.pdf.

Anonymous:2012:BMh

[359] Anonymous. Back matter. American Mathematical Monthly, 119(8): 712, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.08.bm.pdf.

Anonymous:2012:FMi

[360] Anonymous. Front matter. American Mathematical Monthly, 119 (9):713–714, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.09.fm.pdf.

Eisermann:2012:FTA

[361] Michael Eisermann. The fundamental theorem of algebra made effective: An elementary real-algebraic proof via Sturm chains. American Mathematical Monthly, 119(9):715–752, November 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.09.715.pdf.

deOliveira:2012:FTA

[362] Oswaldo Rio Branco de Oliveira. The fundamental theorem of algebra: From the four basic operations. American Mathematical Monthly, 119(9):753–758, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.09.753.pdf.

Lorch:2012:MSS

[363] John Lorch. Magic squares and Sudoku. American Mathematical Monthly, 119(9):759–770, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.09.759.pdf.

Panagopoulou:2012:PGB

[364] Panagiota N. Panagopoulou and Paul G. Spirakis. Playing a game to bound the chromatic number. American Mathematical Monthly, 119 (9):771–778, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.09.771.pdf.

Brink:2012:STS

[365] David Brink. The spiral of Theodorus and sums of zeta-values at the half-integers. American Mathematical Monthly, 119(9):779–786, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.09.779.pdf.

Shapiro:2012:FRB

[366] B. Shapiro and M. Shapiro. A few riddles behind Rolle’s Theorem. American Mathematical Monthly, 119(9):787–793, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.09.787.pdf.

Alexeev:2012:NWL

[367] Boris Alexeev and Jacob Tsimerman. Note on a War-like card game. American Mathematical Monthly, 119(9):793–795, November 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.09.793.pdf.

Breuer:2012:NIW

[368] Florian Breuer. Newton identities for Weierstrass products. American Mathematical Monthly, 119(9):796–799, November 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.09.796.pdf.

Anonymous:2012:PSi

[369] Anonymous. Problems and solutions. American Mathematical Monthly, 119(9):800–807, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.09.800.pdf.

Wagon:2012:BRB

[370] Stan Wagon. Book review: In Pursuit of the Traveling Salesman, by William J. Cook. Princeton University Press, Princeton, NJ, 2012 xiii + 228 pp., ISBN 978-0-691-15270-7, $27.95. American Mathematical Monthly, 119(9):808–811, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.09.808.pdf.

Anonymous:2012:BMi

[371] Anonymous. Back matter. American Mathematical Monthly, 119(9):812, November 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.09.bm.pdf.

Anonymous:2012:FMj

[372] Anonymous. Front matter. American Mathematical Monthly, 119 (10):813–814, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.10.fm.pdf.

Khovanova:2012:KP

[373] Tanya Khovanova and Alexey Radul. Killer problems. American Mathematical Monthly, 119(10):815–823, December 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.10.815.pdf.

Jantzen:2012:MII

[374] Robert T. Jantzen and Klaus Volpert. On the mathematics of income inequality: Splitting the Gini index in two. American Mathematical Monthly, 119(10):824–837, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.10.824.pdf.

Martin:2012:ES

[375] Carla D. Martin and Mason A. Porter. The extraordinary SVD. American Mathematical Monthly, 119(10):838–851, December 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.10.838.pdf.

Trench:2012:EVW

[376] William F. Trench. An elementary view of Weyl’s Theory of Equal Distribution. American Mathematical Monthly, 119(10):852–861, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.10.852.pdf.

Chen:2012:ESF

[377] Yong-Gao Chen and Min Tang. On the elementary symmetric functions of 1, 1/2, …, 1/n. American Mathematical Monthly, 119(10):862–867, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.10.862.pdf.

Frank:2012:HSD

[378] Rolfdieter Frank and Harald Riede. Hyperplane sections of the ndimensional cube. American Mathematical Monthly, 119(10):868–872, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.10.868.pdf.

Siehler:2012:THC

[379] Jacob A. Siehler. Tiling Hamiltonian cycles on the 24-cell. American Mathematical Monthly, 119(10):872–875, December 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.119.10.872.pdf.

Sadowski:2012:TFM

[380] Witold Sadowski. A toy forecast model. American Mathematical Monthly, 119(10):876–879, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.119.10.876.pdf.

Anonymous:2012:PSj

[381] Anonymous. Problems and solutions. American Mathematical Monthly, 119(10):880–887, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.10.880.pdf.

Alexanderson:2012:BRBb

[382] Gerald L. Alexanderson. Book review: Transcending Tradition: Jewish Mathematicians in German-Speaking Academic Culture, by Birgit Bergmann, Moritz Epple, and Ruti Ungar, editors. Springer Verlag, Berlin and Heidelberg, 2012, xi + 289 pp., ISBN 978-3-642-22463-8, $49.95. American Mathematical Monthly, 119(10):888–891, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.10.888.pdf.

Anonymous:2012:EE

[383] Anonymous. Editor’s endnotes. American Mathematical Monthly, 119 (10):892–894, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.10.892.pdf.

Anonymous:2012:TIA

[384] Anonymous. Title index, author index, problems proposed, problems solved, solutions, referees. American Mathematical Monthly, 119 (10):895–905, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.10.895.pdf.

Anonymous:2012:BMj

[385] Anonymous. Back matter. American Mathematical Monthly, 119 (10):906–908, December 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.119.10.bm.pdf.

Anonymous:2013:FMa

[386] Anonymous. Front matter. American Mathematical Monthly, 120(1): 1–2, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.01.fm.pdf.

Baeth:2013:FTD

[387] Nicholas R. Baeth and Roger Wiegand. Factorization theory and decompositions of modules. American Mathematical Monthly, 120(1):3– 34, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.01.003.pdf.

Brams:2013:PCC

[388] Steven J. Brams, Michael A. Jones, and Christian Klamler. N-person cake-cutting: There may be no perfect division. American Mathematical Monthly, 120(1):35–47, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.01.035.pdf.

Murty:2013:TGE

[389] M. Ram Murty and Anastasia Zaytseva. Transcendence of generalized Euler constants. American Mathematical Monthly, 120(1):48–54, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.01.048.pdf.

Hiriart-Urruty:2013:WC

[390] Jean-Baptiste Hiriart-Urruty and Patrice Lassère. When is L^{r}(R) contained in L^{p}(R) + L^{q}(R)? American Mathematical Monthly, 120(1):

55–61, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.01.055.pdf.

Anghel:2013:USG

[391] Vinicius Nicolae Petre Anghel. A use of symmetry: Generalization of an integral identity found by M. L. Glasser. American Mathematical Monthly, 120(1):62–69, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.01.062.pdf.

Chen:2013:PZM

[392] Yi Jun Chen. A proof that Zeilberger missed: a new proof of an identity by Chaundy and Bullard based on the Wilf–Zeilberger Method. American Mathematical Monthly, 120(1):69–70, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.01.069.pdf.

Withers:2013:VAM

[393] L. P. Withers, Jr. Visual angular momentum: Mamikon meets Kepler. American Mathematical Monthly, 120(1):71–73, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.01.071.pdf.

Bak:2013:SCD

[394] Wlodzimierz Bak. A simple characterization of differentiation. American Mathematical Monthly, 120(1):74–75, January 2013. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.120.
01.074.pdf.

Anonymous:2013:PSa

[395] Anonymous. Problems and solutions. American Mathematical Monthly, 120(1):76–83, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.01.076.pdf.

Gouvea:2013:BRB

[396] Fernando Q. Gouvêa. Book review: Elliptic Curves, Modular Forms, and Their L-functions (Student Mathematical Library IAS/Park City Subseries, vol, by Álvaro Lozano-Robledo. 58), American Mathematical Society, Providence, RI, 2011, 195 pp., ISBN 978-0-8218-5242-2, $29.60. American Mathematical Monthly, 120(1):84–94, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.01.084.pdf.

Anonymous:2013:BMa

[397] Anonymous. Back matter. American Mathematical Monthly, 120(1):95– 96, January 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.01.bm.pdf.

Anonymous:2013:FMb

[398] Anonymous. Front matter. American Mathematical Monthly, 120(2):97– 98, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.02.fm.pdf.

Teismann:2013:TMC

[399] Holger Teismann. Toward a more complete list of completeness axioms. American Mathematical Monthly, 120(2):99–114, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.02.099.pdf.

Dyson:2013:LIS

[400] Freeman J. Dyson, Norman E. Frankel, and M. Lawrence Glasser. Lehmer’s interesting series. American Mathematical Monthly, 120 (2):116–130, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.02.116.pdf.

Kozen:2013:MT

[401] Dexter Kozen and Alexandra Silva. On Moessner’s theorem. American Mathematical Monthly, 120(2):131–139, February 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.02.131.pdf.

Marshall:2013:HTL

[402] Susan H. Marshall and Alexander R. Perlis. Heronian tetrahedra are lattice tetrahedra. American Mathematical Monthly, 120(2):140–149, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.02.140.pdf.

Alm:2013:RT

[403] Jeremy Alm, Michael Gramelspacher, and Theodore Rice. Rubik’s on the torus. American Mathematical Monthly, 120(2):150–160, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.02.150.pdf.

Ullrich:2013:SEP

[404] David C. Ullrich. A simple elementary proof of Hilbert’s inequality. American Mathematical Monthly, 120(2):161–164, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.02.161.pdf.

Melman:2013:TTC

[405] Aaron Melman. The twin of a theorem by Cauchy. American Mathematical Monthly, 120(2):164–168, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.02.164.pdf.

Dalai:2013:HWR

[406] Marco Dalai. How would Riemann evaluate ζ(2n)? American Mathematical Monthly, 120(2):169–171, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.02.169.pdf.

Dragovic:2013:PDP

[407] Vladimir Dragović. Polynomial dynamics and a proof of the Fermat Little Theorem. American Mathematical Monthly, 120(2):171–173, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.02.171.pdf.

Anonymous:2013:PSb

[408] Anonymous. Problems and solutions. American Mathematical Monthly, 120(2):174–181, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.02.174.pdf.

Nathanson:2013:BRB

[409] Michael Nathanson. Book review: Elements of Information Theory, by Thomas M. Cover and Joy A. Thomas. John Wiley and Sons, Inc., Hoboken, NJ, 2006, xxiv + 748 pp., ISBN 0-471-24195-4, $111.00. American Mathematical Monthly, 120(2):182–187, February 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.02.182.pdf.

Anonymous:2013:BMb

[410] Anonymous. Back matter. American Mathematical Monthly, 120(2): 188, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.02.bm.pdf.

Anonymous:2013:FMc

[411] Anonymous. Front matter. American Mathematical Monthly, 120 (3):189–190, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.03.fm.pdf.

Schleicher:2013:LGE

[412] Dierk Schleicher and Serge Tabachnikov. Letter from the Guest Editors. American Mathematical Monthly, 120(3):191, March 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.03.191.pdf.

Conway:2013:UAP

[413] John H. Conway. On unsettleable arithmetical problems. American Mathematical Monthly, 120(3):192–198, March 2013. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.120.
03.192.pdf.

Foote:2013:TBT

[414] Robert Foote, Mark Levi, and Serge Tabachnikov. Tractrices, bicycle tire tracks, hatchet planimeters, and a 100-year-old conjecture. American Mathematical Monthly, 120(3):199–216, March 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.03.199.pdf.

Fuchs:2013:EIS

[415] Dmitry Fuchs. Evolutes and involutes of spatial curves. American Mathematical Monthly, 120(3):217–231, March 2013. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.120.
03.217.pdf.

Ghys:2013:ICV

[416] Étienne Ghys. Intersecting curves (variation on an observation of Maxim Kontsevich). American Mathematical Monthly, 120(3):232–242, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.03.232.pdf.

Malter:2013:NLO

[417] Aimeric Malter, Dierk Schleicher, and Don Zagier. New looks at old number theory. American Mathematical Monthly, 120(3):243–264, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.03.243.pdf.

Tokieda:2013:RM

[418] Tadashi Tokieda. Roll models. American Mathematical Monthly, 120(3):265–282, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.03.265.pdf.

Zivaljevic:2013:BRB

[419] Rade T. Zivaljević. Book review: Mathematical Omnibus: Thirty Lectures on Classic Mathematics, by Dmitry Fuchs and Serge Tabachnikov. American Mathematical Society, Providence, RI, 2007, xv + 463 pp., ISBN-13: 978-0-8218-4316-1, $61. American Mathematical Monthly, 120(3):283–286, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.03.283.pdf.

Moree:2013:BRB

[420] Pieter Moree. Book review: An Invitation to Mathematics: From Competitions to Research, [edited] by Dierk Schleicher and Malte Lackmann, editors. Springer-Verlag, Heidelberg, 2011, xiv + 220 pp. ISBN-13: 978-3-642-19532-7, $34.95. American Mathematical Monthly, 120(3):286– 288, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.03.286.pdf.

Poppe:2013:BRB

[421] Christoph Pöppe. Book review: Dimensions, by Jos Leys, Étienne Ghys, and Aurélien Alvarez. A mathematical video. www.dimensions-math.org, 117 minutes, no cost. American Mathematical Monthly, 120(3):288– 290, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.03.288.pdf.

Anonymous:2013:BMc

[422] Anonymous. Back matter. American Mathematical Monthly, 120(3):291– 292, March 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.03.bm.pdf.

Anonymous:2013:FMd

[423] Anonymous. Front matter. American Mathematical Monthly, 120(4): 293–294, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.04.fm.pdf.

Watkins:2013:YGG

[424] Ann E. Watkins. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2013 to William A. Hawkins for Distinguished Service to Mathematics. American Mathematical Monthly, 120(4):295–297, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.04.295.pdf.

Apostol:2013:NBP

[425] Tom M. Apostol and Mamikon A. Mnatsakanian. New balancing principles applied to circumsolids of revolution, and to n-dimensional spheres, cylindroids, and cylindrical wedges. American Mathematical Monthly, 120(4):298–321, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.04.298.pdf.

Jenssen:2013:IFL

[426] Matthew Jenssen, Daniel Montealegre, and Vadim Ponomarenko. Irreducible factorization lengths and the elasticity problem within N.

American Mathematical Monthly, 120(4):322–328, April 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.04.322.pdf.

Williams:2013:PJF

[427] Kenneth S. Williams. The parents of Jacobi’s four squares theorem are unique. American Mathematical Monthly, 120(4):329–345, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.04.329.pdf.

Nyman:2013:BUE

[428] Kathryn L. Nyman and Francis Edward Su. A Borsuk–Ulam equivalent that directly implies Sperner’s Lemma. American Mathematical Monthly, 120(4):346–354, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.120.04.346.pdf.

Muzaffar:2013:NPC

[429] Habib Bin Muzaffar. A new proof of a classical formula. American Mathematical Monthly, 120(4):355–358, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.04.355.pdf.

Alpern:2013:INN

[430] Steve Alpern and Robbert Fokkink. Illuminating a network from its nodes. American Mathematical Monthly, 120(4):358–359, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.04.358.pdf.

Shalit:2013:SPM

[431] Orr Moshe Shalit. A sneaky proof of the maximum modulus principle. American Mathematical Monthly, 120(4):359–362, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.04.359.pdf.

Bernardi:2013:SPR

[432] Olivier Bernardi. A short proof of Rayleigh’s Theorem with extensions. American Mathematical Monthly, 120(4):362–364, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.04.362.pdf.

Anonymous:2013:PSc

[433] Anonymous. Problems and solutions. American Mathematical Monthly, 120(4):365–372, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.04.365.pdf.

Snipes:2013:BRB

[434] Marie Snipes. Book review: Linear and Nonlinear Programming, 3rd ed, by David G. Luenberger and Yinyu Ye., Springer, New York, 2008, xiv + 546 pp., ISBN 0-387-74502-5, $109. American Mathematical Monthly, 120(4):373–379, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.120.04.373.pdf.

Anonymous:2013:BMd

[435] Anonymous. Back matter. American Mathematical Monthly, 120(4): 380, April 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.04.bm.pdf.

Anonymous:2013:FMe

[436] Anonymous. Front matter. American Mathematical Monthly, 120(5): 381–382, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.05.fm.pdf.

Chapman:2013:LE

[437] Scott Chapman. A letter from the Editor. American Mathematical Monthly, 120(5):383, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.05.383.pdf.

Kalai:2013:FHP

[438] Ehud Kalai. Foreword: The high priest of game theory. American Mathematical Monthly, 120(5):384–385, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.05.384.pdf.

Gale:2013:CAS

[439] D. Gale and L. S. Shapley. College admissions and the stability of marriage. American Mathematical Monthly, 120(5):386–391, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.05.386.pdf.

Propp:2013:RAR

[440] James Propp. Real analysis in reverse. American Mathematical Monthly, 120(5):392–408, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.05.392.pdf.

Nathanson:2013:FPR

[441] Melvyn B. Nathanson. On the fractional parts of roots of positive real numbers. American Mathematical Monthly, 120(5):409–429, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.05.409.pdf.

Buckley:2013:VTR

[442] Stephen M. Buckley and Desmond MacHale. Variations on a theme: Rings satisfying x^{3} = x are commutative. American Mathematical Monthly, 120(5):430–440, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.120.05.430.pdf.

Cooper:2013:GGG

[443] Joshua Cooper and Aaron Dutle. Greedy Galois games. American Mathematical Monthly, 120(5):441–451, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.05.441.pdf.

Marsli:2013:GMG

[444] Rachid Marsli and Frank J. Hall. Geometric multiplicities and Gersgorin discs. American Mathematical Monthly, 120(5):452–455, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.05.452.pdf.

Xiang:2013:NCS

[445] Jim X. Xiang. A note on the Cauchy–Schwarz inequality. American Mathematical Monthly, 120(5):456–459, May 2013. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.120.
05.456.pdf.

Efthimiou:2013:CCR

[446] Costas J. Efthimiou. A class of continued radicals. American Mathematical Monthly, 120(5):459–461, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.05.459.pdf.

Aharonov:2013:BID

[447] D. Aharonov and U. Elias. A binomial identity via differential equations. American Mathematical Monthly, 120(5):462–466, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.05.462.pdf.

Galaz-Fontes:2013:APN

[448] F. Galaz-Fontes. Another proof for non-supercyclicity in finite dimensional complex Banach spaces. American Mathematical Monthly, 120(5): 466–468, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.05.466.pdf.

Anonymous:2013:PSd

[449] Anonymous. Problems and solutions. American Mathematical Monthly, 120(5):469–476, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.05.469.pdf.

Woods:2013:BRB

[450] Kevin Woods. Book review: Symmetry: A Mathematical Exploration, by Kristopher Tapp. Springer, New York, 2012, xiv + 215pp., ISBN 978-1-4614-0298-5, $49.95. American Mathematical Monthly, 120(5):477– 481, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.05.477.pdf.

Anonymous:2013:EEa

[451] Anonymous. Editor’s endnotes. American Mathematical Monthly, 120(5):482–483, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.05.482.pdf.

Anonymous:2013:BMe

[452] Anonymous. Back matter. American Mathematical Monthly, 120(5): 484, May 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.05.bm.pdf.

Anonymous:2013:FMf

[453] Anonymous. Front matter. American Mathematical Monthly, 120(6): 485–486, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.06.fm.pdf.

Hwang:2013:PSG

[454] Andrew D. Hwang. Paper surface geometry. American Mathematical Monthly, 120(6):487–499, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.06.487.pdf.

Papakonstantinou:2013:OES

[455] Joanna M. Papakonstantinou and Richard A. Tapia. Origin and evolution of the secant method in one dimension. American Mathematical Monthly, 120(6):500–518, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.06.500.pdf.

Bliss:2013:SDC

[456] Nathan Bliss, Ben Fulan, Stephen Lovett, and Jeff Sommars. Strong divisibility, cyclotomic polynomials, and iterated polynomials. American Mathematical Monthly, 120(6):519–536, June 2013. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.120.
06.519.pdf.

Weintraub:2013:SPI

[457] Steven H. Weintraub. Several proofs of the irreducibility of the cyclotomic polynomials. American Mathematical Monthly, 120(6):537– 545, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.06.537.pdf.

Girstmair:2013:EIS

[458] Kurt Girstmair. The equation x^{2} + m̄y^{2} = k̄ in Z/pZ. American Mathematical Monthly, 120(6):546–552, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.06.546.pdf.

Henry:2013:CFT

[459] Claire-Soizic Henry. Coxeter friezes and triangulations of polygons. American Mathematical Monthly, 120(6):553–558, June 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.06.553.pdf.

Peterson:2013:PPB

[460] Jonathon Peterson. A probabilistic proof of a binomial identity. American Mathematical Monthly, 120(6):558–562, June 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.06.558.pdf.

Rankin:2013:EAL

[461] S. A. Rankin. The Euclidean algorithm and the linear Diophantine equation ax + by = gcd(a, b). American Mathematical Monthly, 120(6):562– 564, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.06.562.pdf.

Tokieda:2013:BDP

[462] Tadashi Tokieda. A buoyancy-driven perpetual motion machine. American Mathematical Monthly, 120(6):564–566, June 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.06.564.pdf.

Sahoo:2013:EME

[463] Manas R. Sahoo. Example of a monotonic, everywhere differentiable function on R whose derivative is not continuous. American Mathematical Monthly, 120(6):566–568, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.06.566.pdf.

Anonymous:2013:PSe

[464] Anonymous. Problems and solutions. American Mathematical Monthly, 120(6):569–576, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.06.569.pdf.

Bressoud:2013:BRB

[465] David Bressoud. Book review: Calculus, by Michael Spivak; Calculus Deconstructed: A Second Course in First-Year Calculus. MAA Textbooks, Mathematical Association of America, Washington, DC, 2009, xvi + 491 pp., ISBN 978-0-88385-756-4, $ 59.95. by Zbigniew Nitecki; Approximately Calculus. American Mathematical Society, Providence, RI, 2006, xvii + 292 pp., ISBN 978-0-8218-3750-4, $50.00. by Shahriar Sharhriari; Guide to Cauchy’s Calculus: A Translation and Analysis of Calcul Infinitésimal. Fairview Academic Press, Walnut Creek, CA, 2012, viii + 616 pp., ISBN 978-0-9838837-0-8, $139.00. by Dennis M. Cates; The Calculus Integral. ClassicalRealAnalysis.com, Charleston, SC, 2010, x + 291, ISBN 978-1442180956, $14.95. by Brian S. Thomson. 4th edition. Publish or Perish, Houston, TX, 2008, xiv + 680 pp., ISBN 978-0-914098-91-1, $85.00. American Mathematical Monthly, 120(6):577–580, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.06.577.pdf.

Anonymous:2013:BMf

[466] Anonymous. Back matter. American Mathematical Monthly, 120(6): ??, June 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.06.bm.pdf.

Anonymous:2013:FMg

[467] Anonymous. Front matter. American Mathematical Monthly, 120 (7):581–582, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.07.fm.pdf.

Aldous:2013:UPM

[468] David J. Aldous. Using prediction market data to illustrate undergraduate probability. American Mathematical Monthly, 120(7):583–593, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.583.pdf.

Casey:2013:PPP

[469] Stephen D. Casey and Brian M. Sadler. Pi, the primes, periodicities, and probability. American Mathematical Monthly, 120(7):594– 608, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.594.pdf.

Kuchment:2013:ISX

[470] Peter Kuchment and Sergey Lvin. Identities for sin x that came from medical imaging. American Mathematical Monthly, 120(7):609–621, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.609.pdf.

Butler:2013:SUA

[471] Steve Butler and Ron Graham. Subdivision using angle bisectors is dense in the space of triangles. American Mathematical Monthly, 120(7):622– 630, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.622.pdf.

Katz:2013:CNI

[472] Mikhail G. Katz and Eric Leichtnam. Commuting and noncommuting infinitesimals. American Mathematical Monthly, 120(7):631–641, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.631.pdf.

Ritelli:2013:API

[473] Daniele Ritelli. Another proof of ζ(2) = \frac{π^{2}}{6} using double integrals. American Mathematical Monthly, 120(7):642–645, August 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.07.642.pdf.

Arnold:2013:SAF

[474] Matthias Arnold and Werner Glaß. Simple approximation formulas for the birthday problem. American Mathematical Monthly, 120(7):645– 648, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.645.pdf.

Bevelacqua:2013:AIC

[475] Anthony J. Bevelacqua. Another irreducibility criterion. American Mathematical Monthly, 120(7):648–650, August 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.07.648.pdf.

Meher:2013:RPB

[476] Jaban Meher and M. Ram Murty. Ramanujan’s proof of Bertrand’s postulate. American Mathematical Monthly, 120(7):650–653, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.07.650.pdf.

Samuel:2013:SPV

[477] T. Samuel. A simple proof of Vitali’s theorem for signed measures. American Mathematical Monthly, 120(7):654–659, August 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.07.654.pdf.

Anonymous:2013:PSf

[478] Anonymous. Problems and solutions. American Mathematical Monthly, 120(7):660–669, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.07.660.pdf.

Malkevitch:2013:BRB

[479] Joseph Malkevitch. Book review: The Mathematical Coloring Book, by Alexander Soifer. Springer, New York, xxx + 607 pp., ISBN 978-0-387-74640-1, $69.95. American Mathematical Monthly, 120(7):670– 674, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.670.pdf.

Moreno:2013:BRB

[480] Samuel G. Moreno and Esther M. García-Caballero. Book review: On the Irrationality of √2 Once Again. American Mathematical Monthly, 120 (7):674, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.674.pdf.

Anonymous:2013:BMg

[481] Anonymous. Back matter. American Mathematical Monthly, 120(7):675– 676, August 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.07.bm.pdf.

Anonymous:2013:FMh

[482] Anonymous. Front matter. American Mathematical Monthly, 120 (8):677–678, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.08.fm.pdf.

Klosinski:2013:STW

[483] Leonard F. Klosinski, Gerald L. Alexanderson, and Mark Krusemeyer. The Seventy-Third William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 120(8):679–688, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.08.679.pdf.

Tucker:2013:HUP

[484] Alan Tucker. The history of the undergraduate program in mathematics in the United States. American Mathematical Monthly, 120(8):689– 705, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.08.689.pdf.

Jones:2013:SMG

[485] Brant Jones, Laura Taalman, and Anthony Tongen. Solitaire Mancala games and the Chinese Remainder Theorem. American Mathematical Monthly, 120(8):706–724, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.08.706.pdf.

Monsky:2013:FRS

[486] Paul Monsky. Frobenius’ result on simple groups of order \frac{p^{3}−p^{2}}{2} . American Mathematical Monthly, 120(8):725–732, October 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.08.725.pdf.

Bagchi:2013:HDA

[487] Bhaskar Bagchi and Basudeb Datta. Higher-dimensional analogues of the map coloring problem. American Mathematical Monthly, 120(8):733– 737, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.08.733.pdf.

Dutkay:2013:SFE

[488] Dorin Ervin Dutkay, Constantin P. Niculescu, and Florin Popovici. Stirling’s formula and its extension for the gamma function. American Mathematical Monthly, 120(8):737–740, October 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.08.737.pdf.

Schaftingen:2013:DPE

[489] Jean Van Schaftingen. A direct proof of the existence of eigenvalues and eigenvectors by Weierstrass’s Theorem. American Mathematical Monthly, 120(8):741–746, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.08.741.pdf.

Ramm:2013:VPA

[490] A. G. Ramm. A variational principle and its application to estimating the electrical capacitance of a perfect conductor. American Mathematical Monthly, 120(8):747–751, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.08.747.pdf.

Lengvarszky:2013:PPT

[491] Zsolt Lengvárszky. Proving the Pythagorean theorem via infinite dissections. American Mathematical Monthly, 120(8):751–753, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.08.751.pdf.

Anonymous:2013:PSg

[492] Anonymous. Problems and solutions. American Mathematical Monthly, 120(8):754–761, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.08.754.pdf.

Konstantopoulos:2013:BRB

[493] Takis Konstantopoulos. Book review: An Introduction to Measure Theory, by Terence Tao. American Mathematical Society, Providence RI, 2011, xvi + 206 pp., ISBN 978-0-8218-6919-2, $53.00. American Mathematical Monthly, 120(8):762–768, October 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.08.762.pdf.

Anonymous:2013:BMh

[494] Anonymous. Back matter. American Mathematical Monthly, 120(8): ??, October 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.08.bm.pdf.

Anonymous:2013:FMi

[495] Anonymous. Front matter. American Mathematical Monthly, 120 (9):769–770, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.09.fm.pdf.

Vehel:2013:CH

[496] Jacques Lévy Véhel and Franklin Mendivil. Christiane’s hair. American Mathematical Monthly, 120(9):771–786, November 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.09.771.pdf.

Velleman:2013:WEG

[497] Daniel J. Velleman and Gregory S. Warrington. What to expect in a game of memory. American Mathematical Monthly, 120(9):787–805, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.09.787.pdf.

Slavik:2013:MPM

[498] Antonín Slavík. Mixing problems with many tanks. American Mathematical Monthly, 120(9):806–821, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.09.806.pdf.

Bayless:2013:RSK

[499] Jonathan Bayless and Dominic Klyve. Reciprocal sums as a knowledge metric: Theory, computation, and perfect numbers. American Mathematical Monthly, 120(9):822–831, November 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.09.822.pdf.

Villacorta:2013:CIA

[500] Jorge Luis Cimadevilla Villacorta. Certain inequalities associated with the divisor function. American Mathematical Monthly, 120(9):832–837, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.09.832.pdf.

Kahane:2013:NCL

[501] Charles S. Kahane. On the nonexistence of certain limits for the complex exponential. American Mathematical Monthly, 120(9):837–840, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.09.837.pdf.

Benyi:2013:GRT

[502] Árpád Bényi and Branko Ćurgus. A generalization of Routh’s Triangle Theorem. American Mathematical Monthly, 120(9):841–846, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.09.841.pdf.

Pan:2013:MPH

[503] David Pan. A maximum principle for high-order derivatives. American Mathematical Monthly, 120(9):846–848, November 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.09.846.pdf.

Woeginger:2013:NNA

[504] Gerhard J. Woeginger. Nothing new about equiangular polygons. American Mathematical Monthly, 120(9):849–850, November 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.09.849.pdf.

Garcia:2013:QGP

[505] Stephan Ramon Garcia. Quotients of Gaussian primes. American Mathematical Monthly, 120(9):851–853, November 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.09.851.pdf.

Anonymous:2013:PSh

[506] Anonymous. Problems and solutions. American Mathematical Monthly, 120(9):854–861, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.09.854.pdf.

Mayfield:2013:BRB

[507] Betty Mayfield. Book review: Calculus: Modeling and Application, second edition, by David A. Smith and Lawrence C. Moore. Mathematical Association of America, Washington, D.C., 2010. http://maa.pinnaclecart.com. Price: $35.00. ISBN 978-1-6144-610-1. American Mathematical Monthly, 120(9):862–864, November 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.09.862.pdf.

Anonymous:2013:BMi

[508] Anonymous. Back matter. American Mathematical Monthly, 120(9):??, November 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.09.bm.pdf.

Anonymous:2013:FMj

[509] Anonymous. Front matter. American Mathematical Monthly, 120 (10):865–866, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.10.fm.pdf.

Pengelley:2013:QDE

[510] David Pengelley. Quick, does 23/67 equal 33/97 ? A mathematician’s secret from Euclid to today. American Mathematical Monthly, 120(10):867– 876, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.10.867.pdf.

Campbell:2013:LAC

[511] Alexander P. Campbell and Daniel Daners. Linear algebra via complex analysis. American Mathematical Monthly, 120(10):877–892, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.10.877.pdf.

Haggstrom:2013:RCA

[512] Olle Häggström and Johan Wästlund. Rigorous computer analysis of the Chow–Robbins game. American Mathematical Monthly, 120(10):893– 900, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.10.893.pdf.

Traves:2013:PTD

[513] Will Traves. From Pascal’s theorem to d-constructible curves. American Mathematical Monthly, 120(10):901–915, December 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.10.901.pdf.

Howe:2013:CC

[514] Roger E. Howe. The Cuoco configuration. American Mathematical Monthly, 120(10):916–923, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.120.10.916.pdf.

Abel:2013:GLR

[515] Ulrich Abel. A generalization of the Leibniz Rule. American Mathematical Monthly, 120(10):924–928, December 2013. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.120.
10.924.pdf.

Sondow:2013:PPA

[516] Jonathan Sondow. The Parbelos, a parabolic analog of the Arbelos. American Mathematical Monthly, 120(10):929–935, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.10.929.pdf.

Jameson:2013:IGF

[517] G. J. O. Jameson. Inequalities for gamma function ratios. American Mathematical Monthly, 120(10):936–940, December 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.10.936.pdf.

Anonymous:2013:PSi

[518] Anonymous. Problems and solutions. American Mathematical Monthly, 120(10):941–948, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.10.941.pdf.

Henle:2013:BRB

[519] James M. Henle and Michael G. Henle. Book review: Real Analysis Through Modern Infinitesimals, Cambridge University Press, 2011, xix + 565 pp, by Nader Vakil., ISBN 978-1-107-00202-9, $127.00. American Mathematical Monthly, 120(10):949–953, December 2013. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.120.10.949.pdf.

Anonymous:2013:EEb

[520] Anonymous. Editor’s endnotes. American Mathematical Monthly, 120 (10):954–956, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.10.954.pdf.

Anonymous:2013:TIA

[521] Anonymous. Title index, author index, problems proposed, problems solved, solutions, referees. American Mathematical Monthly, 120 (10):957–967, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.120.10.957.pdf.

Anonymous:2013:BMj

[522] Anonymous. Back matter. American Mathematical Monthly, 120(10): 968, December 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.120.10.bm.pdf.

Anonymous:2014:FMa

[523] Anonymous. Front matter. American Mathematical Monthly, 121(1): 1–2, January 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.01.fm.pdf.

Chapman:2014:LEW

[524] Scott T. Chapman. A letter from the editor: What’s in a color? American Mathematical Monthly, 121(1):3–4, January 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.01.003.pdf.

Nathanson:2014:AST

[525] Melvyn B. Nathanson. Additive systems and a theorem of de Bruijn. American Mathematical Monthly, 121(1):5–17, January 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.01.005.pdf.

Ponce:2014:ESG

[526] Mario Ponce and Patricio Santibáñez. On equidistant sets and generalized conics: The old and the new. American Mathematical Monthly, 121(1):18–32, January 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.121.01.018.pdf.

Mendes:2014:GG

[527] Anthony Mendes and Kent E. Morrison. Guessing games. American Mathematical Monthly, 121(1):33–44, January 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
01.033.pdf.

Lynd:2014:UDE

[528] Chris D. Lynd. Using difference equations to generalize results for periodic nested radicals. American Mathematical Monthly, 121(1):45–59, January 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.01.045.pdf.

Conejero:2014:WIT

[529] J. A. Conejero, P. Jiménez-Rodríguez, G. A. Muñoz-Fernández, and J. B. Seoane-Sepúlveda. When the identity theorem “seems” to fail. American Mathematical Monthly, 121(1):60–68, January 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.01.060.pdf.

Besenyei:2014:PUP

[530] Ádám Besenyei. Peano’s unnoticed proof of Borel’s Theorem. American Mathematical Monthly, 121(1):69–72, January 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
01.069.pdf.

Weintraub:2014:VPI

[531] Steven H. Weintraub. Values of polynomials over integral domains. American Mathematical Monthly, 121(1):73–74, January 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.01.073.pdf.

Li:2014:DPT

[532] Bao Qin Li. A direct proof and a transcendental version of the fundamental theorem of algebra via Cauchy’s theorem. American Mathematical Monthly, 121(1):75–77, January 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.01.075.pdf.

Huneke:2014:IPH

[533] Samuel Clowes Huneke. An inductive proof of hex uniqueness. American Mathematical Monthly, 121(1):78–80, January 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.01.078.pdf.

Boulabiar:2014:RVR

[534] Karim Boulabiar. Real-valued ring homomorphisms on C(Ω). American Mathematical Monthly, 121(1):81–82, January 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
01.081.pdf.

Anonymous:2014:PSa

[535] Anonymous. Problems and solutions. American Mathematical Monthly, 121(1):83–90, January 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.121.01.083.pdf.

Ross:2014:BRB

[536] Kenneth A. Ross. Book review: In the Dark on the Sunny Side: a Memoir of an Out-of-Sight Mathematician, Mathematical Association of America, Washington, DC, 2012, 206 pp, by Larry Baggett., ISBN 978-0-88385-581-2, $50. American Mathematical Monthly, 121(1):91– 92, January 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.01.091.pdf.

Anonymous:2014:BMa

[537] Anonymous. Back matter. American Mathematical Monthly, 121(1): ??, January 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.01.bm.pdf.

Anonymous:2014:FMb

[538] Anonymous. Front matter. American Mathematical Monthly, 121(2):93– 94, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.02.fm.pdf.

Chi:2014:LGH

[539] Eric C. Chi and Kenneth Lange. A look at the Generalized Heron Problem through the lens of majorization-minimization. American Mathematical Monthly, 121(2):95–108, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.02.095.pdf.

Coll:2014:TGP

[540] Vincent Coll and Mike Harrison. Two generalizations of a property of the catenary. American Mathematical Monthly, 121(2):109–119, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.02.109.pdf.

Damianou:2014:BSF

[541] Pantelis A. Damianou. A beautiful sine formula. American Mathematical Monthly, 121(2):120–135, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.02.120.pdf.

Evans:2014:KTP

[542] Ron Evans, Lihua Huang, and Tuan Nguyen. Keeler’s Theorem and products of distinct transpositions. American Mathematical Monthly, 121(2):136–144, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.121.02.136.pdf.

Gaspar:2014:NRP

[543] Jaime Gaspar. Non-recursive polynomial formula for the sum of the powers of the integers. American Mathematical Monthly, 121(2):144, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.02.144.pdf.

Grant:2014:QGS

[544] David Grant. The quadratic Gauss sum redux. American Mathematical Monthly, 121(2):145–149, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.02.145.pdf.

Edwards:2014:PNP

[545] Miles Dillon Edwards. A possibly new proof of the law of cos. American Mathematical Monthly, 121(2):149, February 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.02.149.pdf.

Billig:2014:ETS

[546] Yuly Billig and John D. Dixon. An eigenvalue theorem for systems of polynomial equations. American Mathematical Monthly, 121(2):150–153, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.02.150.pdf.

Lou:2014:SPS

[547] Hongwei Lou. A short proof of Stirling’s formula. American Mathematical Monthly, 121(2):154–157, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.02.154.pdf.

Ehrenborg:2014:HCA

[548] Richard Ehrenborg. Hamiltonian cycles on Archimedean solids are twisting free. American Mathematical Monthly, 121(2):158–161, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.02.158.pdf.

Siong:2014:SPL

[549] Chua Cheong Siong. A simple proof of Ljunggren’s binomial congruence. American Mathematical Monthly, 121(2):162–164, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.02.162.pdf.

Velleman:2014:MG

[550] Daniel J. Velleman. Monthly gems. American Mathematical Monthly, 121(2):164, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.121.02.164.pdf.

McGrath:2014:APC

[551] Peter J. McGrath. Another proof of Clairaut’s theorem. American Mathematical Monthly, 121(2):165–166, February 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.02.165.pdf.

Giv:2014:PBA

[552] Hossein Hosseini Giv. Proving the Banach–Alaoglu theorem via the existence of the Stone-Čech compactification. American Mathematical Monthly, 121(2):167–169, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.02.167.pdf.

Anonymous:2014:PSb

[553] Anonymous. Problems and solutions. American Mathematical Monthly, 121(2):170–177, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.121.02.170.pdf.

Boas:2014:BRB

[554] Harold P. Boas. Book review: Invitation to Classical Analysis, American Mathematical Society, Providence, RI, 2012, xiii + 392 pp, by Peter Duren., ISBN 978-0-8218-6932-1, $74.00. American Mathematical Monthly, 121(2):178–182, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.02.178.pdf.

Lax:2014:BRB

[555] Peter D. Lax. Book review: Numbers and Functions: From a Classical Experimental Mathematicians Point of View, American Mathematical Society, Providence, RI, 2012, 504 pp, by Victor H. Moll., ISBN 978-0-8218-8795-0, $58. American Mathematical Monthly, 121(2):183, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.02.183.pdf.

Anonymous:2014:BMb

[556] Anonymous. Back matter. American Mathematical Monthly, 121(2): 184, February 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.02.bm.pdf.

Anonymous:2014:FMc

[557] Anonymous. Front matter. American Mathematical Monthly, 121 (3):185–186, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.121.03.fm.pdf.

Ross:2014:YGG

[558] Kenneth A. Ross and Ann E. Watkins. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2014 to Joan Leitzel for distinguished service to mathematics. American Mathematical Monthly, 121(3):187–190, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.03.187.pdf.

Bailey:2014:PDUa

[559] David H. Bailey and Jonathan Borwein. Pi Day is upon us again and we still do not know if pi is normal. American Mathematical Monthly, 121(3):191–206, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.121.03.191.pdf. Chi-

nese translation in Mathematical Advances in Translation (Chinese Academy of Science) [metadata unknown].

Berendonk:2014:LSC

[560] Stephan Berendonk. The length of spirographic curves. American Mathematical Monthly, 121(3):207–212, March 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
03.207.pdf.

Cooper:2014:MGM

[561] Alec S. Cooper, Oleg Pikhurko, John R. Schmitt, and Gregory S. Warrington. Martin Gardner’s minimum no-3-in-a-line problem. American Mathematical Monthly, 121(3):213–221, March 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
03.213.pdf.

Maliszewski:2014:LSD

[562] Aleksander Maliszewski and Marcin Szyszkowski. Level sets on disks. American Mathematical Monthly, 121(3):222–228, March 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.03.222.pdf.

Choi:2014:SIZ

[563] Junesang Choi and Hari M. Srivastava. Series involving the zeta functions and a family of generalized Goldbach–Euler series. American Mathematical Monthly, 121(3):229–236, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.03.229.pdf.

Sury:2014:PPF

[564] B. Sury. A polynomial parent to a Fibonacci–Lucas relation. American Mathematical Monthly, 121(3):236, March 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
03.236.pdf.

Hyde:2014:WPC

[565] Trevor Hyde. A Wallis product on clovers. American Mathematical Monthly, 121(3):237–243, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.121.03.237.pdf.

Merca:2014:SPS

[566] Mircea Merca. On some power sums of sine or cosine. American Mathematical Monthly, 121(3):244–248, March 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
03.244.pdf.

Johnston:2014:WHP

[567] William Johnston. The weighted Hermite polynomials form a basis for L_{2}(R). American Mathematical Monthly, 121(3):249–253, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.03.249.pdf.

Herzog:2014:PLP

[568] Gerd Herzog. A proof of Lie’s product formula. American Mathematical Monthly, 121(3):254–257, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/pdfplus/10.4169/amer.math.monthly.121.03.254.pdf.

Lange:2014:HDI

[569] Kenneth Lange. Hadamard’s determinant inequality. American Mathematical Monthly, 121(3):258–259, March 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.121.
03.258.pdf.

Kocak:2014:MMM

[570] S. Koçak and M. Limoncu. Mean of the mean of the mean . . .. American Mathematical Monthly, 121(3):259, March 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/pdfplus/10.4169/amer.math.
monthly.121.03.259.pdf.

Moore:2014:JTM

[571] Justin Tatch Moore. Justin Tatch Moore. American Mathematical Monthly, 121(3):260–262, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/amer.math.monthly.121.03.260.pdf.

Chamberland:2014:SPM

[572] Marc Chamberland and Doron Zeilberger. A short proof of Mc-Dougall’s Circle Theorem. American Mathematical Monthly, 121(3):263– 265, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.03.263.pdf.

Anonymous:2014:PSc

[573] Anonymous. Problems and solutions. American Mathematical Monthly, 121(3):266–273, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

pdfplus/10.4169/amer.math.monthly.121.03.266.pdf.

Kozek:2014:BRB

[574] Mark Kozek. Book review: Mathematics in Popular Culture: Essays on Appearances in Film, Fiction, Games, Television and Other Media. Edited by Jessica K. Sklar and Elizabeth S. Sklar; Foreword by Keith Devlin. McFarland, Jefferson, NC, 2012, viii + 345 pp., ISBN 978-0-7864-4978-1, $45. American Mathematical Monthly, 121(3):274– 278, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.03.274.pdf.

Anonymous:2014:BMc

[575] Anonymous. Back matter. American Mathematical Monthly, 121(3):279– 280, March 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.03.bm.pdf.

Anonymous:2014:FMd

[576] Anonymous. Front matter. American Mathematical Monthly, 121(4): 281–282, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.04.fm.

Simoson:2014:PDT

[577] Andrew J. Simoson. Periodicity domains and the transit of Venus. American Mathematical Monthly, 121(4):283–298, April 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.04.283.

Velleman:2014:DIR

[578] Daniel J. Velleman. A drug-induced random walk. American Mathematical Monthly, 121(4):299–317, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.04.299.

Uteshev:2014:ASG

[579] Alexei Yu. Uteshev. Analytical solution for the generalized Fermat– Torricelli problem. American Mathematical Monthly, 121(4):318–331, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL https://www.jstor.org/stable/10.4169/amer.
math.monthly.121.04.318.

Ferguson:2014:OSL

[580] Samuel J. Ferguson. A one-sentence line-of-sight proof of the Extreme Value Theorem. American Mathematical Monthly, 121(4):331, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.04.331.

Kovar:2014:PEU

[581] Martin Kovár and Alena Chernikava. On the proof of the existence of undominated strategies in normal form games. American Mathematical Monthly, 121(4):332–337, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.04.332.

Chen:2014:AFX

[582] Chao-Ping Chen and Junesang Choi. An asymptotic formula for (1 + 1/x)^{x} based on the partition function. American Mathematical Monthly, 121(4):338–343, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.121.04.338.

Eger:2014:SAC

[583] Steffen Eger. Stirling’s approximation for central extended binomial coefficients. American Mathematical Monthly, 121(4):344–349, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.04.344.

Neuschel:2014:NPS

[584] Thorsten Neuschel. A new proof of Stirling’s formula. American Mathematical Monthly, 121(4):350–352, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.04.350.

Krause:2014:ZOA

[585] Ralph M. Krause. Zeta(2) once again. American Mathematical Monthly, 121(4):353–354, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.04.353.

Hyde:2014:PSM

[586] Andrea M. Hyde, Paul D. Lee, and Blair K. Spearman. Polynomials (x^{3} − n)(x^{2} + 3) solvable modulo any integer. American Mathematical Monthly, 121(4):355–358, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.04.355.

Sury:2014:ME

[587] B. Sury. Macaulay expansion. American Mathematical Monthly, 121(4): 359–360, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.04.359.

Koliha:2014:ELI

[588] J. J. Koliha. Evaluating Lebesgue integrals efficiently with the FTC. American Mathematical Monthly, 121(4):361–364, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.04.361.

Anonymous:2014:PSd

[589] Anonymous. Problems and solutions. American Mathematical Monthly, 121(4):365–372, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.04.365.

Nunemacher:2014:BRB

[590] Jeffrey Nunemacher. Book review: Encounters with Chaos and Fractals, 2nd edition. By Denny Gulick. Chapman and Hall/CRC Press, Boca Raton, 2012, xvi + 371 pp., ISBN 978-1-58488-517-7, $79.95. American Mathematical Monthly, 121(4):373–376, April 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.04.373.

Anonymous:2014:BMd

[591] Anonymous. Back matter. American Mathematical Monthly, 121(4): ??, April 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.04.bm.

Anonymous:2014:FMe

[592] Anonymous. Front matter. American Mathematical Monthly, 121(5): 377–378, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.05.fm.

Henrich:2014:UU

[593] Allison Henrich and Louis H. Kauffman. Unknotting unknots. American Mathematical Monthly, 121(5):379–390, May 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.05.379.

Beardon:2014:GRC

[594] Alan F. Beardon and Ian Short. A geometric representation of continued fractions. American Mathematical Monthly, 121(5):391–402, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.05.391.

Leinster:2014:RST

[595] Tom Leinster. Rethinking set theory. American Mathematical Monthly, 121(5):403–415, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.05.403.

Kroopnick:2014:EAB

[596] Allan J. Kroopnick. An elementary application of Brouwer’s fixed point theorem to transition matrices. American Mathematical Monthly, 121 (5):415, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.05.415.

Grantham:2014:RAD

[597] Jon Grantham, Witold Jarnicki, John Rickert, and Stan Wagon. Repeatedly appending any digit to generate composite numbers. American Mathematical Monthly, 121(5):416–421, May 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.05.416.

Ben-Ari:2014:WTP

[598] Iddo Ben-Ari, Diana Hay, and Alexander Roitershtein. On Wallis-type products and Pólya’s urn schemes. American Mathematical Monthly, 121(5):422–432, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.05.422.

Pollack:2014:PEF

[599] Paul Pollack and Enrique Treviño. The primes that Euclid forgot. American Mathematical Monthly, 121(5):433–437, May 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.05.433.

Tsukerman:2014:SSP

[600] Emmanuel Tsukerman. Solution of Sondow’s problem: A synthetic proof of the tangency property of the Parbelos. American Mathematical Monthly, 121(5):438–443, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.05.438.

Araujo:2014:DEL

[601] José Ángel Cid Araújo. A difference equation leading to the irrationality of √2. American Mathematical Monthly, 121(5):443, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.05.443.

Carlson:2014:CBF

[602] Nathan A. Carlson. A connection between Furstenberg’s and Euclid’s proofs of the infinitude of primes. American Mathematical Monthly, 121 (5):444, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.05.444.

Li:2014:EPG

[603] Ming-Chia Li. An elementary proof of a generalization of Banach’s mapping theorem. American Mathematical Monthly, 121(5):445–446, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.05.445.

Glasser:2014:TCI

[604] M. L. Glasser. A technique in contour integration. American Mathematical Monthly, 121(5):447–448, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.05.447.

Luh:2014:LDB

[605] Kyle Luh and Nicholas Pippenger. Large-deviation bounds for sampling without replacement. American Mathematical Monthly, 121(5):449–454, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.05.449.

Anonymous:2014:PSe

[606] Anonymous. Problems and solutions. American Mathematical Monthly, 121(5):455–462, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.05.455.

Ullman:2014:BRB

[607] Daniel Ullman. Book review: Math on Trial: How Mathematics Is Used and Abused in the Courtroom, By Leila Schneps and Coralie Colmez, Basic Books, New York, 2013, xi + 256 pp., ISBN 978-0-465-03292-1, $26.99. American Mathematical Monthly, 121(5):463–466, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.05.463.

Anonymous:2014:EEa

[608] Anonymous. Editor’s endnotes. American Mathematical Monthly, 121(5):467–468, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.05.467.

Anonymous:2014:BMe

[609] Anonymous. Back matter. American Mathematical Monthly, 121(5):??, May 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.05.bm.

Anonymous:2014:FMf

[610] Anonymous. Front matter. American Mathematical Monthly, 121(6): 469–470, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.06.fm.

Facchini:2014:RNF

[611] Alberto Facchini and Giulia Simonetta. Rational numbers, finite cyclic monoids, divisibility rules, and numbers of type 99 ⋯ 900 ⋯ 0. American Mathematical Monthly, 121(6):471–485, June 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.06.471.

Borwein:2014:SLM

[612] David Borwein, Jonathan M. Borwein, and Brailey Sims. On the solution of linear mean recurrences. American Mathematical Monthly, 121(6):486– 498, June/July 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://docserver.carma.newcastle.edu.au/
1505/;   https://www.jstor.org/stable/10.4169.

Rukavicka:2014:RLD

[613] Josef Rukavicka. Rejection of Laplace’s demon. American Mathematical Monthly, 121(6):498, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.06.498.

Groetsch:2014:VPR

[614] C. W. Groetsch and S. A. Yost. Vertical projection in a resisting medium: Reflections on observations of Mersenne. American Mathematical Monthly, 121(6):499–505, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.06.499.

Carter:2014:TDK

[615] J. Scott Carter, Daniel S. Silver, and Susan G. Williams. Three dimensions of knot coloring. American Mathematical Monthly, 121(6):506–514, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.06.506.

Kwong:2014:APS

[616] Harris Kwong. An alternate proof of Sury’s Fibonacci–Lucas relation. American Mathematical Monthly, 121(6):514, June 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.06.514.

Pujol:2014:HNF

[617] Jose Pujol. On Hamilton’s nearly-forgotten early work on the relation between rotations and quaternions and on the composition of rotations. American Mathematical Monthly, 121(6):515–522, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.06.515.

Sury:2014:HAH

[618] B. Sury. A heuristic argument for Hua’s identity using geometric series. American Mathematical Monthly, 121(6):522, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.06.522.

Wright:2014:LIC

[619] Tommy Wright. Lagrange’s identity and Congressional apportionment. American Mathematical Monthly, 121(6):523–528, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.06.523.

Wu:2014:PR

[620] Chai Wah Wu. Pigeonholes and repunits. American Mathematical Monthly, 121(6):529–533, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.121.06.529.

Ferland:2014:RCP

[621] Kevin K. Ferland. Record crossword puzzles. American Mathematical Monthly, 121(6):534–536, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.06.534.

Spivey:2014:CPA

[622] Michael Z. Spivey. A combinatorial proof for the alternating convolution of the central binomial coefficients. American Mathematical Monthly, 121(6):537–540, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.121.06.537.

Nikolov:2014:CGL

[623] Nikolai Nikolov and Blagovest Sendov. A converse of the Gauss–Lucas theorem. American Mathematical Monthly, 121(6):541–544, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.06.541.

Nigussie:2014:FYS

[624] Yared Nigussie. Finding your seat versus tossing a coin. American Mathematical Monthly, 121(6):545–546, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.06.545.

Badertscher:2014:SDP

[625] Erich Badertscher. A simple direct proof of Marden’s theorem. American Mathematical Monthly, 121(6):547–548, June 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.06.547.

Anonymous:2014:PSf

[626] Anonymous. Problems and solutions. American Mathematical Monthly, 121(6):549–556, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.06.549.

Swenson:2014:BRB

[627] James A. Swenson. Book review: Which Numbers Are Real?, By Michael Henle. Mathematical Association of America, Washington, DC, 2012, x + 219 pp., ISBN 978-0-88385-777-9, $53.00. American Mathematical Monthly, 121(6):557–560, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.06.557.

Anonymous:2014:BMf

[628] Anonymous. Back matter. American Mathematical Monthly, 121(6):??, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.06.bm.

Anonymous:2014:FMg

[629] Anonymous. Front matter. American Mathematical Monthly, 121 (7):561–562, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.07.fm.

Brugalle:2014:BTG

[630] Erwan Brugallé and Kristin Shaw. A bit of tropical geometry. American Mathematical Monthly, 121(7):563–589, August 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.07.563.

Brafman:2014:AAP

[631] Marcelo Brafman. Another algebraic Pythagorean proof. American Mathematical Monthly, 121(7):589, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.121.07.589.

Brown:2014:FQS

[632] Bryan Brown, Michael Dairyko, Stephan Ramon Garcia, Bob Lutz, and Michael Someck. Four quotient set gems. American Mathematical Monthly, 121(7):590–598, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.07.590.

deSa:2014:BFN

[633] Vinícius G. Pereira de Sá and Celina M. H. de Figueiredo. Blindfriendly von Neumann’s heads or tails. American Mathematical Monthly, 121(7):600–609, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.07.600.

Katz:2014:ZSU

[634] Richard Katz, Mike Krebs, and Anthony Shaheen. Zero sums on unit square vertex sets and plane colorings. American Mathematical Monthly, 121(7):610–618, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.07.610.

Zagier:2014:RTJ

[635] D. Zagier. A refinement of a theorem of J. E. Littlewood. American Mathematical Monthly, 121(7):618, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.07.618.

Bhatia:2014:PLM

[636] Rajendra Bhatia and Rajesh Sharma. Positive linear maps and spreads of matrices. American Mathematical Monthly, 121(7):619–624, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.07.619.

Chang:2014:TAP

[637] Wen D. Chang and Russell A. Gordon. Trisecting angles in Pythagorean triangles. American Mathematical Monthly, 121(7):625–631, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.07.625.

Marques:2014:YAD

[638] Diego Marques. Yet another direct proof of the uncountability of the transcendental numbers. American Mathematical Monthly, 121(7):631, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL https://www.jstor.org/stable/10.4169/amer.
math.monthly.121.07.631.

Pinchasi:2014:PCO

[639] Rom Pinchasi. Points covered an odd number of times by translates. American Mathematical Monthly, 121(7):632–636, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.07.632.

Geck:2014:CGE

[640] Meinolf Geck. On the characterization of Galois extensions. American Mathematical Monthly, 121(7):637–639, August 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.07.637.

Moreno:2014:SRS

[641] Samuel G. Moreno and Esther M. García-Caballero. Simpson’s rule and sums of powers. American Mathematical Monthly, 121(7):639, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.07.639.

Dreher:2014:CIC

[642] F. Dreher and T. Samuel. Continuous images of Cantor’s ternary set. American Mathematical Monthly, 121(7):640–643, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.07.640.

Beezer:2014:EEF

[643] Robert A. Beezer. Extended echelon form and four subspaces. American Mathematical Monthly, 121(7):644–647, August 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.07.644.

Anonymous:2014:PSg

[644] Anonymous. Problems and solutions. American Mathematical Monthly, 121(7):648–657, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.07.648.

Barr:2014:BRB

[645] Michael Barr. Book review: The Outer Limits of Reason, Noson Yanofsky. MIT Press, Cambridge, MA and London, England, 2013, xiv + 403 pp., ISBN 978-0-262-01935-4, $29.95. American Mathematical Monthly, 121(7):658–660, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.07.658.

Anonymous:2014:BMg

[646] Anonymous. Back matter. American Mathematical Monthly, 121(7): ??, August 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.07.bm.

Anonymous:2014:FMh

[647] Anonymous. Front matter. American Mathematical Monthly, 121 (8):659–662, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.08.fm.

Klosinski:2014:SFW

[648] Leonard F. Klosinski, Gerald L. Alexanderson, and Mark Krusemeyer. The Seventy-Fourth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 121(8):663–673, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.08.663.

Diaconis:2014:CGT

[649] Persi Diaconis, Xuancheng Shao, and Kannan Soundararajan. Carries, group theory, and additive combinatorics. American Mathematical Monthly, 121(8):674–688, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.08.674.

Deleglise:2014:MAL

[650] Marc Deléglise and Andrew Markoe. On the maximum arc length of monotonic functions. American Mathematical Monthly, 121(8):689–699, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.08.689.

Borwein:2014:DNS

[651] David Borwein and Jonathan Borwein. Deriving new sinc results from old. American Mathematical Monthly, 121(8):700–705, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://docserver.carma.newcastle.edu.au/1502/;   https:
//www.jstor.org/stable/10.4169/amer.math.monthly.121.08.700.

Joswig:2014:FTL

[652] Michael Joswig and Günter M. Ziegler. Foldable triangulations of lattice polygons. American Mathematical Monthly, 121(8):706–710, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.08.706.

Novak:2014:PRW

[653] Jonathan Novak. Pólya’s random walk theorem. American Mathematical Monthly, 121(8):711–716, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.08.711.

Schramm:2014:SAS

[654] Michael Schramm, John Troutman, and Daniel Waterman. Segmentally alternating series. American Mathematical Monthly, 121(8):717– 722, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.08.717.

Molokach:2014:LCP

[655] John Molokach. Law of cosines — a proof without words. American Mathematical Monthly, 121(8):722, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.121.08.722.

Holmes:2014:CBC

[656] Mark H. Holmes. Connections between cubic splines and quadrature rules. American Mathematical Monthly, 121(8):723–726, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.08.723.

Osborne:2014:HS

[657] M. Scott Osborne. Hausdorffization and such. American Mathematical Monthly, 121(8):727–733, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.08.727.

Davis:2014:BCI

[658] Donald M. Davis. Binomial coefficients involving infinite powers of primes. American Mathematical Monthly, 121(8):734–737, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.08.734.

Anonymous:2014:PSh

[659] Anonymous. Problems and solutions. American Mathematical Monthly, 121(8):738–745, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Loft:2014:BRB

[660] Brian M. Loft. Book review: The Cult of Pythagorus: Math and Myths, by Alberto A. Martínez. University of Pittsburgh Press, Pittsburgh, PA, 2012, xxii + 264 pp., ISBN 978-0-8229-4418-8, $27.95. American Mathematical Monthly, 121(8):746–748, October 2014. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.121.08.746.

Anonymous:2014:BMh

[661] Anonymous. Back matter. American Mathematical Monthly, 121(8): ??, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.08.bm.

Anonymous:2014:FMi

[662] Anonymous. Front matter. American Mathematical Monthly, 121 (9):747–750, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.09.fm.

Anonymous:2014:FMj

[663] Anonymous. Front matter. American Mathematical Monthly, 121 (9):747–750, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.09.fm.

Adler:2014:LGE

[664] Fred Adler, Elizabeth S. Allman, and Lisette G. de Pillis. A letter from the Guest Editors. American Mathematical Monthly, 121 (9):751–753, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.09.751.

Potts:2014:MAT

[665] Jonathan R. Potts and Mark A. Lewis. A mathematical approach to territorial pattern formation. American Mathematical Monthly, 121 (9):754–770, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.09.754.

Steel:2014:TEL

[666] Mike Steel. Tracing evolutionary links between species. American Mathematical Monthly, 121(9):771–792, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.121.09.771.

Edwards:2014:DGN

[667] Roderick Edwards and Leon Glass. Dynamics in genetic networks. American Mathematical Monthly, 121(9):793–809, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.09.793.

Nguyen:2014:CFA

[668] Hoa Nguyen, Ricardo Cortez, and Lisa Fauci. Computing flows around microorganisms: Slender-body theory and beyond. American Mathematical Monthly, 121(9):810–823, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.09.810.

Terman:2014:MN

[669] David Terman. Mathematical neuroscience. American Mathematical Monthly, 121(9):824–839, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.09.824.

Jackson:2014:MOU

[670] Trachette Jackson, Natalia Komarova, and Kristin Swanson. Mathematical oncology: Using mathematics to enable cancer discoveries. American Mathematical Monthly, 121(9):840–856, November 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.09.840.

Heyer:2014:SBN

[671] Laurie J. Heyer and Jeffrey L. Poet. Synthetic biology: A new frontier. American Mathematical Monthly, 121(9):857–867, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
121.09.857.

Anonymous:2014:BMi

[672] Anonymous. Back matter. American Mathematical Monthly, 121(9):??, November 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.09.bm.

Anonymous:2014:BMj

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.09.bm.

Anonymous:2014:FMk

[674] Anonymous. Front matter. American Mathematical Monthly, 121 (10):867–870, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.10.fm.

Anonymous:2014:FMl

[675] Anonymous. Front matter. American Mathematical Monthly, 121 (10):867–870, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.10.fm.

Giblin:2014:EE

[676] Peter J. Giblin and J. Paul Warder. Evolving evolutoids. American Mathematical Monthly, 121(10):871–889, December 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.10.871.

Moree:2014:NSC

[677] Pieter Moree. Numerical semigroups, cyclotomic polynomials, and Bernoulli numbers. American Mathematical Monthly, 121(10):890–902, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.10.890.

Ni:2014:PTT

[678] Lei Ni. A Perron-type theorem on the principal eigenvalue of nonsymmetric elliptic operators. American Mathematical Monthly, 121 (10):903–908, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.10.903.

Clark:2014:ACO

[679] Pete L. Clark and Niels J. Diepeveen. Absolute convergence in ordered fields. American Mathematical Monthly, 121(10):909–916, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.10.909.

Trench:2014:UCS

[680] William F. Trench. An unnoticed consequence of Szegő’s distribution theorem. American Mathematical Monthly, 121(10):917–921, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.10.917.

Button:2014:CIG

[681] Jack Button, Maurice Chiodo, and Mariano Zeron-Medina Laris. Coset intersection graphs for groups. American Mathematical Monthly, 121 (10):922–926, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.10.922.

Northshield:2014:CDC

[682] Sam Northshield. Complex Descartes circle theorem. American Mathematical Monthly, 121(10):927–931, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.121.10.927.

Dey:2014:LAP

[683] Pallab Kanti Dey and R. Thangadurai. The length of an arithmetic progression represented by a binary quadratic form. American Mathematical Monthly, 121(10):932–936, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.10.932.

Hampton:2014:CCT

[684] Marshall Hampton. Cosines and Cayley, triangles and tetrahedra. American Mathematical Monthly, 121(10):937–941, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.121.10.937.

Acker:2014:NST

[685] Felipe Acker. A note on the spectral theorem in the finite-dimensional real case. American Mathematical Monthly, 121(10):942–945, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.10.942.

Anonymous:2014:PSi

[686] Anonymous. Problems and solutions. American Mathematical Monthly, 121(10):946–953, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.10.946.

Anonymous:2014:PSj

[687] Anonymous. Problems and solutions. American Mathematical Monthly, 121(10):946–953, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.10.946.

Walsh:2014:BRB

[688] Genevieve S. Walsh. Book review: Mostly Surfaces, By Richard Evan Schwartz. American Mathematical Society, Providence, RI, 2010, xiii + 314 pp., ISBN 978-0-8218-5368-9, $47.00. American Mathematical Monthly, 121(10):954–956, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.10.954.

Anonymous:2014:EEb

[689] Anonymous. Editor’s endnotes. American Mathematical Monthly, 121 (10):957–959, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.121.10.957.

Anonymous:2014:MR

[690] Anonymous. Monthly referees for 2014. American Mathematical Monthly, 121(10):960–962, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.121.10.960.

Sondow:2014:BYA

[691] Jonathan Sondow. From the Monthly over 100 years ago . . .. American Mathematical Monthly, 121(10):963, December 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.10.963.

Sondow:2014:MYA

[692] Jonathan Sondow. From the Monthly over 100 years ago. . .. American Mathematical Monthly, 121(10):963, December 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.10.963.

Nimbran:2014:OMP

[693] Amrik Singh Nimbran. One more proof of the irrationality of \sqrt{2}. American Mathematical Monthly, 121(10):964, December 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.121.10.964.

Anonymous:2014:BMk

[694] Anonymous. Back matter. American Mathematical Monthly, 121(10):??, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.10.bm.

Anonymous:2014:BMl

[695] Anonymous. Back matter. American Mathematical Monthly, 121(10):??, December 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.121.10.bm.

Anonymous:2015:FMa

[696] Anonymous. Front matter. American Mathematical Monthly, 122(1): 1–2, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.01.fm.

Chapman:2015:LEC

[697] Scott T. Chapman. A letter from the Editor: 2015 is the centennial year of the MAA. American Mathematical Monthly, 122(1):3, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.01.3.

Zitarelli:2015:FYM

[698] David E. Zitarelli. The first 100 years of the MAA. American Mathematical Monthly, 122(1):4–23, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.01.4.

Swenton:2015:SPF

[699] Frank Swenton. Simultaneous proof of the first fundamental theorem of calculus and integrability of continuous functions. American Mathematical Monthly, 122(1):23, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.01.23.

Andreescu:2015:NP

[700] Titu Andreescu, Vladimir Georgiev, Oleg Mushkarov, Vladimir Georgiev, and Oleg Mushkarov. Napoleon polygons. American Mathematical Monthly, 122(1):24–29, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.01.24.

Williams:2015:HRR

[701] Kenneth S. Williams. Historical remark on Ramanujan’s tau function. American Mathematical Monthly, 122(1):30–35, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.01.30.

Curgus:2015:SSM

[702] Branko Ćurgus, Robert I. Jewett, and Robert I. Jewett. Somewhat stochastic matrices. American Mathematical Monthly, 122(1):36–42, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.01.36.

Hardy:2015:STI

[703] Michael Hardy. Stereographic trigonometric identities. American Mathematical Monthly, 122(1):43–47, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.01.43.

Lin:2015:IMP

[704] Xianzu Lin. Infinitely many primes in the arithmetic progression kn − 1. American Mathematical Monthly, 122(1):48–51, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.01.48.

Kayll:2015:DME

[705] P. Mark Kayll. Deranged matchings: Enumeration by integration!! American Mathematical Monthly, 122(1):51, January 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.01.51.

Motta:2015:PTB

[706] Francis C. Motta, Patrick D. Shipman, and Bethany Springer. A point of tangency between combinatorics and differential geometry. American Mathematical Monthly, 122(1):52–55, January 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.01.52.

Giv:2015:ACW

[707] Hossein Hosseini Giv. The axiom of choice, well-ordering, and wellclassification. American Mathematical Monthly, 122(1):56–59, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.01.56.

Ehrenhorg:2015:UMT

[708] R. Ehrenhorg. An unbiased marriage theorem. American Mathematical Monthly, 122(1):59, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.01.59.

Ojeda:2015:KSR

[709] Ignacio Ojeda. Kronecker square roots and the block vec matrix. American Mathematical Monthly, 122(1):60–64, January 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.01.60.

Panzone:2015:FR

[710] Pablo A. Panzone. On a formula of S. Ramanujan. American Mathematical Monthly, 122(1):65–69, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.01.65.

Letzter:2015:CAF

[711] Edward S. Letzter. Continuity is an adjoint functor. American Mathematical Monthly, 122(1):70–74, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.01.70.

Anonymous:2015:PSa

[712] Anonymous. Problems and solutions. American Mathematical Monthly, 122(1):75–82, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.01.75.

Schaefer:2015:BRB

[713] Edward F. Schaefer. Book review: The Mathematics of Encryption: An Elementary Introduction, Reviewed work(s): The Mathematics of Encryption: An Elementary Introduction. By Margaret Cozzens and Steven J. Miller. American Mathematical Society, Providence, RI, 2013, xviii + 332 pp., ISBN 978-0-8218-8321-1, $49.00. American Mathematical Monthly, 122(1):83–88, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.01.83.

Anonymous:2015:BMa

[714] Anonymous. Back matter. American Mathematical Monthly, 122(1): ??, January 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.01.bm.

Anonymous:2015:FMb

[715] Anonymous. Front matter. American Mathematical Monthly, 122(2):87– 90, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.02.fm.

Nievergelt:2015:BAE

[716] Yves Nievergelt. Backward analysis explained with monthly payments. American Mathematical Monthly, 122(2):91–120, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.02.91.

ONeill:2015:HDY

[717] Christopher O’Neill, Roberto Pelayo, and Roberto Pelayo. How do you measure primality? American Mathematical Monthly, 122(2):121–137, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.02.121.

Shukurov:2015:PSN

[718] Aydin Shukurov. The power system is never a basis in the space of continuous functions. American Mathematical Monthly, 122(2):137, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.02.137.

delaTorre:2015:NPS

[719] Carlos A. Catalá de la Torre. A new proof for a sharp van der Corput’s lemma. American Mathematical Monthly, 122(2):138–142, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.02.138.

Kocak:2015:DAR

[720] S. Koçak. Dihedral angles of the regular n-simplex. American Mathematical Monthly, 122(2):143, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.02.143.

Akopyan:2015:GC

[721] Arseniy V. Akopyan. Geometry of the cardioid. American Mathematical Monthly, 122(2):144–150, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.02.144.

Brafman:2015:AAP

[722] Marcelo Brafman. Another algebraic Pythagorean proof. American Mathematical Monthly, 122(2):150, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.122.02.150.

Molinari:2015:EPI

[723] G. Molinari. Elementary prime identities. American Mathematical Monthly, 122(2):151–154, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.02.151.

Grater:2015:EBI

[724] Joachim Gräter, Karl-Joachim Wirths, and Karl-Joachim Wirths. On elementary bounds for ∑_{k=n}^{∞} k^{−s}. American Mathematical Monthly, 122(2):155–158, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.02.155.

Pollack:2015:EGE

[725] Paul Pollack. An easy generalization of Euler’s theorem on the series of prime reciprocals. American Mathematical Monthly, 122(2):159–163, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.02.159.

Drucker:2015:CPT

[726] Daniel Drucker. A comprehensive Pythagorean theorem for all dimensions. American Mathematical Monthly, 122(2):164–168, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.02.164.

Moreno:2015:GPMa

[727] Samuel G. Moreno and Esther M. García-Caballero. A geometric proof of Morrie’s Law. American Mathematical Monthly, 122(2):168, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.02.168.

Li:2015:TEP

[728] Bao Qin Li. Two elementary properties of entire functions. American Mathematical Monthly, 122(2):169–172, February 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.02.169.

Dai:2015:FBC

[729] Xiongping Dai. From the first Borel–Cantelli lemma to Poincaré’s recurrence theorem. American Mathematical Monthly, 122(2):173–174, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.02.173.

Anonymous:2015:PSb

[730] Anonymous. Problems and solutions. American Mathematical Monthly, 122(2):175–182, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.02.175.

Folland:2015:BRBa

[731] Gerald B. Folland. Book review: Hidden Harmony—Geometric Fantasies: The Rise of Complex Function Theory, By Umberto Bottazzini and Jeremy Gray. Springer, New York, 2013, xvii + 848 pp., ISBN 978-1-4614-5724-4, $189.00 (hardcover), ISBN 978-1-4614-5725-1, $149.00 (eBook). American Mathematical Monthly, 122(2):183–188, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.02.183.

Anonymous:2015:BMb

[732] Anonymous. Back matter. American Mathematical Monthly, 122(2): ??, February 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.02.bm.

Anonymous:2015:FMc

[733] Anonymous. Front matter. American Mathematical Monthly, 122 (3):189–190, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.03.fm.

Quinn:2015:YGG

[734] Jennifer J. Quinn. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2015 to W. James Lewis for distinguished service to mathematics. American Mathematical Monthly, 122(3):191–194, March 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.03.191.

Borwein:2015:PPB

[735] Jonathan M. Borwein and Scott T. Chapman. I prefer pi: a brief history and anthology of articles in the American Mathematical Monthly. American Mathematical Monthly, 122(3):195–216, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/1615/;
 https://www.jstor.org/stable/10.4169;   https://www.jstor.
org/stable/10.4169/amer.math.monthly.122.03.195. See addenda

[845].

Peifer:2015:MDO

[736] David Peifer. Max Dehn and the origins of topology and infinite group theory. American Mathematical Monthly, 122(3):217–233, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.03.217.

Moreno:2015:GPMb

[737] Samuel G. Moreno and Esther M. García-Caballero. A geometric proof of Morrie’s Law. American Mathematical Monthly, 122(3):233, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.03.233.

Eising:2015:SAG

[738] Jaap Eising, David Radcliffe, and Jaap Top. A simple answer to Gelfand’s question. American Mathematical Monthly, 122(3):234–245, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.03.234.

DeLeon:2015:UUC

[739] Doreen De Leon. Using undetermined coefficients to solve certain classes of variable-coefficient equations. American Mathematical Monthly, 122 (3):246–255, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.03.246.

Beck:2015:VGF

[740] Matthias Beck and Neville Robbins. Variations on a generating-function theme: Enumerating compositions with parts avoiding an arithmetic sequence. American Mathematical Monthly, 122(3):256–263, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.03.256.

Petrunin:2015:AMP

[741] Anton Petrunin. Area minimizing polyhedral surfaces are saddle. American Mathematical Monthly, 122(3):264–267, March 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.03.264.

Bloch:2015:ODM

[742] William Goldbloom Bloch. Open discontinuous maps from R^{n} onto R^{n}. American Mathematical Monthly, 122(3):268–271, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.03.268.

Khan:2015:AMH

[743] Mizan R. Khan and Richard Magner. An application of modular hyperbolas to quadratic residues. American Mathematical Monthly, 122(3):272–274, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.03.272.

Tarnauceanu:2015:FGC

[744] Marius Tarnauceanu. Finite groups with a certain number of cyclic subgroups. American Mathematical Monthly, 122(3):275–276, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.03.275.

Konvalina:2015:PMA

[745] John Konvalina. Powers of matrices and algebraic centrosymmetry. American Mathematical Monthly, 122(3):277–279, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.03.277.

Horak:2015:PCC

[746] Peter Horak. Partitioning R^{n} into connected components. American Mathematical Monthly, 122(3):280–283, March 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.03.280.

Gaitanas:2015:EFP

[747] Konstantinos Gaitanas. An explicit formula for the prime counting function which is valid infinitely often. American Mathematical Monthly, 122(3):283, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.03.283.

Anonymous:2015:PSc

[748] Anonymous. Problems and solutions. American Mathematical Monthly, 122(3):284–291, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.03.284.

Hersh:2015:BRB

[749] Reuben Hersh. Book review: How Humans Learn to Think Mathematically. Exploring the Three Worlds of Mathematics. American Mathematical Monthly, 122(3):292–296, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.03.292.

Anonymous:2015:BMc

[750] Anonymous. Back matter. American Mathematical Monthly, 122(3): ??, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.03.bm.

Anonymous:2015:FMd

[751] Anonymous. Front matter. American Mathematical Monthly, 122(4): 297–298, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.04.fm.

Montgomery:2015:TBP

[752] Richard Montgomery. The three-body problem and the shape sphere. American Mathematical Monthly, 122(4):299–321, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.04.299.

Gilbert:2015:FR

[753] Rebekah Ann Gilbert. A fine rediscovery. American Mathematical Monthly, 122(4):322–331, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.122.04.322.

Cufi:2015:CP

[754] Juliá Cufí, Agustí Reventó, and Carlos J. RodrÍz. Curvature for polygons. American Mathematical Monthly, 122(4):332–337, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.04.332.

Dilcher:2015:PMR

[755] Karl Dilcher and Larry Ericksen. The polynomials of Mahler and roots of unity. American Mathematical Monthly, 122(4):338–353, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.04.338.

Santos:2015:NMS

[756] Altino F. Santos. A new metric on spherical f-tilings. American Mathematical Monthly, 122(4):354–361, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.04.354.

Fitzpatrick:2015:ACF

[757] Patrick M. Fitzpatrick and Brian R. Hunt. Absolute continuity of a function and uniform integrability of its divided differences. American Mathematical Monthly, 122(4):362–366, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.04.362.

Ponomarenko:2015:BCS

[758] Vadim Ponomarenko. A base conversion surprise. American Mathematical Monthly, 122(4):366 402–366, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.04.366.

Adamaszek:2015:FNS

[759] Michal Adamaszek. Face numbers of down-sets. American Mathematical Monthly, 122(4):367–370, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.04.367.

Huylebrouck:2015:GWF

[760] Dirk Huylebrouck. Generalizing Wallis’ formula. American Mathematical Monthly, 122(4):371–372, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.04.371.

Kilic:2015:NGP

[761] Mehmet Kilic. A new geometric proof for Morley’s Theorem. American Mathematical Monthly, 122(4):373–376, April 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.04.373.

Katsuura:2015:NIS

[762] Hidefumi Katsuura. A new infinite series representation of ln k. American Mathematical Monthly, 122(4):376, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.122.04.376.

Arnold:2015:CET

[763] Maxim Arnold and Vadim Zharnitsky. Cyclic evasion in the three bug problem. American Mathematical Monthly, 122(4):377–380, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.04.377.

Baralic:2015:SPB

[764] orde Baralić. A short proof of the Bradley Theorem. American Mathematical Monthly, 122(4):381–385, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.04.381.

Schmuland:2015:LRP

[765] Byron Schmuland. Linear recurrences via probability. American Mathematical Monthly, 122(4):386–389, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.04.386.

Anonymous:2015:PSd

[766] Anonymous. Problems and solutions. American Mathematical Monthly, 122(4):390–397, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.04.390.

Traves:2015:BRB

[767] Will Traves. Book review: Perspectives in Projective Geometry. A Guided Tour Through Real and Complex Geometry. American Mathematical Monthly, 122(4):398–402, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.122.04.398.

Anonymous:2015:BMd

[768] Anonymous. Back matter. American Mathematical Monthly, 122(4): ??, April 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.04.bm.

Anonymous:2015:FMe

[769] Anonymous. Front matter. American Mathematical Monthly, 122(5): 405–406, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.5.fm.

Saari:2015:MDM

[770] Donald G. Saari. Mathematics and the “dark matter” puzzle. American Mathematical Monthly, 122(5):407–423, May 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.5.407.

Chin:2015:PTT

[771] Alex Chin, Gary Gordon, Kellie MacPhee, and Charles Vincent. Pick a tree — any tree. American Mathematical Monthly, 122(5):424–432, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.5.424.

Aebi:2015:GWT

[772] Christian Aebi and Grant Cairns. Generalizations of Wilson’s Theorem for double-, hyper-, sub- and superfactorials. American Mathematical Monthly, 122(5):433–443, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.5.433.

Ciaurri:2015:SCK

[773] Óscar Ciaurri, Luis M. Navas, Francisco J. Ruiz, and Juan L. Varona. A simple computation of ζ(2k). American Mathematical Monthly, 122(5): 444–451, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.5.444.

Heo:2015:NPP

[774] Nam Gu Heo. A new proof of the Pythagorean theorem. American Mathematical Monthly, 122(5):451, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.5.451.

Boyle:2015:URC

[775] Jeffrey A. Boyle. Using rolling circles to generate caustic envelopes resulting from reflected light. American Mathematical Monthly, 122(5): 452–466, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.5.452.

Northshield:2015:OLP

[776] Sam Northshield. A one-line proof of the infinitude of primes. American Mathematical Monthly, 122(5):466, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.5.466.

Izmestiev:2015:PCQ

[777] Ivan Izmestiev. A porism for cyclic quadrilaterals, butterfly theorems, and hyperbolic geometry. American Mathematical Monthly, 122(5):467– 475, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.5.467.

Bras-Amoros:2015:SLE

[778] Maria Bras-Amorós and Marta Pujol. Side lengths of equiangular polygons (as seen by a coding theorist). American Mathematical Monthly, 122(5):476–478, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.5.476.

Pinelis:2015:RBF

[779] Iosif Pinelis. Relationships between the first four moments. American Mathematical Monthly, 122(5):479–481, May 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.5.479.

Januszewski:2015:UBS

[780] Janusz Januszewski. An upper bound for some coloring problem. American Mathematical Monthly, 122(5):481, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.122.5.481.

Geck:2015:ERS

[781] Meinolf Geck. Eigenvalues of real symmetric matrices. American Mathematical Monthly, 122(5):482–483, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.5.482.

Veklych:2015:OFP

[782] Bogdan Veklych. A one-formula proof of the nonvanishing of L-functions of real characters at 1. American Mathematical Monthly, 122(5):484–485, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.5.484.

Castro:2015:VCH

[783] Jesús Jerónimo Castro. The volume of the convex hull of a body and its homothetic copies. American Mathematical Monthly, 122(5):486–489, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.5.486.

Steele:2015:EMM

[784] J. Michael Steele. Explaining a mysterious maximal inequality — and a path to the law of large numbers. American Mathematical Monthly, 122(5):490–494, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.5.490.

Hardy:2015:UID

[785] Michael Hardy. Using infinitesimals to differentiate secant and tangent. American Mathematical Monthly, 122(5):494, May 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.5.494.

Besenyei:2015:BAG

[786] Ádám Besenyei. The Brocard angle and a geometrical gem from Dmitriev and Dynkin. American Mathematical Monthly, 122(5):495–499, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.5.495.

Anonymous:2015:PSe

[787] Anonymous. Problems and solutions. American Mathematical Monthly, 122(5):500–507, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.5.500.

Folland:2015:BRBb

[788] Gerald B. Folland. Book review: A History in Sum: 150 Years of Mathematics at Harvard (1825–1975), By Steve Nadis and Shing-Tung Yau. Harvard University Press, Cambridge, 2013, xviii + 249 pp., ISBN 978-0-674-72500-3, $39.95. American Mathematical Monthly, 122(5):508–510, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.5.508.

Anonymous:2015:EEa

[789] Anonymous. Editor’s endnotes. American Mathematical Monthly, 122(5):511–512, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.5.511.

Anonymous:2015:BMe

[790] Anonymous. Back matter. American Mathematical Monthly, 122(5):??, May 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.5.bm.

Anonymous:2015:FMf

[791] Anonymous. Front matter. American Mathematical Monthly, 122(6): 513–514, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.6.fm.

Epple:2015:CFS

[792] Dennis D. A. Epple and Jason Siefken. Collapse: A Fibonacci and Sturmian game. American Mathematical Monthly, 122(6):515–527, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.6.515.

Williams:2015:FIT

[793] Kenneth S. Williams. A “four integers” theorem and a “five integers” theorem. American Mathematical Monthly, 122(6):528–536, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.6.528.

Halbeisen:2015:SPP

[794] Lorenz Halbeisen and Norbert Hungerbühler. A simple proof of Poncelet’s theorem (on the occasion of its bicentennial). American Mathematical Monthly, 122(6):537–551, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.6.537.

Salgarkar:2015:ADP

[795] Aaditya Salgarkar and Rishi Mehta. Alternate derivation for the parity of the n^{th} triangular number. American Mathematical Monthly, 122(6):551, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.6.551.

Werner:2015:CNF

[796] Nicholas J. Werner. Covering numbers of finite rings. American Mathematical Monthly, 122(6):552–566, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.6.552.

Rooin:2015:GPB

[797] J. Rooin and A. Alikhani. A geometric proof of the Beppo–Levi inequality. American Mathematical Monthly, 122(6):566, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.6.566.

Haynes:2015:DAC

[798] Alan Haynes and Sara Munday. Diophantine approximation and coloring. American Mathematical Monthly, 122(6):567–580, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.6.567.

Tandra:2015:PRT

[799] Haryono Tandra. A proof of the replacement theorem by the notion of maximal elements. American Mathematical Monthly, 122(6):580, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.6.580.

Murakami:2015:PPS

[800] Satomi Murakami and Hiroki Ohwa. Periodic points of some discontinuous mappings. American Mathematical Monthly, 122(6):581–586, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.6.581.

Ponomarenko:2015:AFX

[801] Vadim Ponomarenko. Asymptotic formula for (1 + 1/x)^{x}, revisited. American Mathematical Monthly, 122(6):587, June 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.6.587.

Barnard:2015:PQR

[802] Virgil Barnard. A proof of quadratic reciprocity. American Mathematical Monthly, 122(6):588–592, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.6.588.

Daners:2015:UCC

[803] Daniel Daners. Uniform continuity of continuous functions on compact metric spaces. American Mathematical Monthly, 122(6):592, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.6.592.

Pinelis:2015:HCS

[804] Iosif Pinelis. On the Hölder and Cauchy–Schwarz inequalities. American Mathematical Monthly, 122(6):593–595, June 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.6.593.

Gascon:2015:APR

[805] José Gascón. Another proof that the real numbers R are uncountable. American Mathematical Monthly, 122(6):596–597, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.6.596.

Tandra:2015:SMV

[806] Haryono Tandra. A simple modified version for Ferguson’s proof of the extreme value theorem. American Mathematical Monthly, 122(6):598, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.6.598.

Merca:2015:AFF

[807] Mircea Merca. An alternative to Faulhaber’s formula. American Mathematical Monthly, 122(6):599–601, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.6.599.

Drammatico:2015:NPE

[808] Tommaso Drammatico. A new proof of the equality of mixed second partial derivatives. American Mathematical Monthly, 122(6):602–603, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.6.602.

Anonymous:2015:PSf

[809] Anonymous. Problems and solutions. American Mathematical Monthly, 122(6):604–612, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.6.604.

Cohen:2015:BRM

[810] Marion Cohen. Book review: Mathematicians on Creativity. Edited by Peter B. Borwein, Peter Liljedahl, and Helen Zhai, The Mathematical Association of America, Washington DC, 2014, xviii + 199 pp, ISBN 978-0-88385-574-4, $30.00. American Mathematical Monthly, 122 (6):613–616, June/July 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.6.613.

Anonymous:2015:BMf

[811] Anonymous. Back matter. American Mathematical Monthly, 122(6):??, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.6.bm.

Anonymous:2015:FMg

[812] Anonymous. Front matter. American Mathematical Monthly, 122 (7):616–618, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.7.fm.

Raman-Sundstrom:2015:PHC

[813] Manya Raman-Sundström. A pedagogical history of compactness. American Mathematical Monthly, 122(7):619–635, August 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.7.619.

Pomerance:2015:DMB

[814] Carl Pomerance. Divisors of the middle binomial coefficient. American Mathematical Monthly, 122(7):636–644, August 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.7.636.

Ashby:2015:IPR

[815] Neil Ashby. Isosynchronous paths on a rotating surface. American Mathematical Monthly, 122(7):645–659, August 2015. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.122.7.645.

Beck:2015:PFS

[816] Matthias Beck, Ana Berrizbeitia, Michael Dairyko, Claudia Rodriguez, Amanda Ruiz, and Schuyler Veeneman. Parking functions, Shi arrangements, and mixed graphs. American Mathematical Monthly, 122(7):660– 673, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.7.660.

Barrientos:2015:FEE

[817] Pablo G. Barrientos. A family of eventually expanding piecewise linear maps of the interval. American Mathematical Monthly, 122(7):674– 680, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.7.674.

Cowen:2015:DER

[818] Robert Cowen. Deleting edges from Ramsey minimal examples. American Mathematical Monthly, 122(7):681–683, August 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.7.681.

Marques:2015:NFL

[819] Diego Marques. A new Fibonacci–Lucas relation. American Mathematical Monthly, 122(7):683, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.7.683.

Tuenter:2015:DBS

[820] Hans J. H. Tuenter. A double binomial sum: Putnam 35-a4 revisited. American Mathematical Monthly, 122(7):684–689, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.7.684.

Hirschhorn:2015:WPC

[821] Michael D. Hirschhorn. Wallis’s product and the central binomial coefficient. American Mathematical Monthly, 122(7):689, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.7.689.

McClure:2015:ETC

[822] James E. McClure and Alec McGail. An elementary treatment of the construction of the free product of groups. American Mathematical Monthly, 122(7):690–692, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.7.690.

Ciaurri:2015:EPE

[823] Óscar Ciaurri. Euler’s product expansion for the sine: An elementary proof. American Mathematical Monthly, 122(7):693–695, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.7.693.

Omey:2015:XOC

[824] Edward Omey. On Xiang’s observations concerning the Cauchy–Schwarz inequality. American Mathematical Monthly, 122(7):696–698, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.7.696.

Moreno:2015:DUS

[825] Samuel G. Moreno. The dramatic and ultimate shortening of a doctoral dissertation in mathematics. American Mathematical Monthly, 122(7): 699, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.7.699.

Anonymous:2015:PSg

[826] Anonymous. Problems and solutions. American Mathematical Monthly, 122(7):700–707, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.7.700.

Malkevitch:2015:BRB

[827] Joseph Malkevitch. Book review: The Mathematical Coloring Book. American Mathematical Monthly, 122(7):708–712, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.7.708.

Anonymous:2015:BMg

[828] Anonymous. Back matter. American Mathematical Monthly, 122(7): ??, August 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.7.bm.

Anonymous:2015:FMh

[829] Anonymous. Front matter. American Mathematical Monthly, 122 (8):713–714, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.8.fm.

Klosinski:2015:SFW

[830] Leonard F. Klosinski, Gerald L. Alexanderson, and Mark Krusemeyer. The Seventy-Fifth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 122(8):715–725, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.8.715.

Cassidy:2015:ERB

[831] California Jack Cassidy. Early round bluffing in poker. American Mathematical Monthly, 122(8):726–744, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.122.8.726.

Apostol:2015:VSA

[832] Tom M. Apostol and Mamikon A. Mnatsakanian. Volume/surface area relations for n-dimensional spheres, pseudospheres, and catenoids. American Mathematical Monthly, 122(8):745–756, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.8.745.

Pollack:2015:BBN

[833] P. Pollack and J. Vandehey. Besicovitch, bisection, and the normality of 0.(1)(4)(9)(16)(25). American Mathematical Monthly, 122(8):757– 765, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.8.757.

Hou:2015:RFR

[834] Shui-Hung Hou and Edwin Hou. On a recursion formula related to confluent Vandermonde. American Mathematical Monthly, 122(8):766–772, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.8.766.

Dzul-Kifli:2015:DCD

[835] Syahida Che Dzul-Kifli and Chris Good. On Devaney chaos and dense periodic points: Period 3 and higher implies chaos. American Mathematical Monthly, 122(8):773–780, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.8.773.

Mortici:2015:FFM

[836] Cristinel Mortici. Funny forms of the Mean Value Theorem. American Mathematical Monthly, 122(8):780, October 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.8.780.

Tokieda:2015:VPC

[837] Tadashi Tokieda. A viscosity proof of the Cauchy–Schwarz inequality. American Mathematical Monthly, 122(8):781, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.8.781.

Kania:2015:SPF

[838] Tomasz Kania. A short proof of the fact that the matrix trace is the expectation of the numerical values. American Mathematical Monthly, 122(8):782–783, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.8.782.

Alpoge:2015:VWP

[839] Levent Alpoge. van der Waerden and the primes. American Mathematical Monthly, 122(8):784–785, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.8.784.

Cerf:2015:EIS

[840] Raphaël Cerf and Matthias Gorny. An exponential inequality for symmetric random variables. American Mathematical Monthly, 122(8):786– 789, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.8.786.

Nathanson:2015:PMR

[841] Melvyn B. Nathanson. Pairs of matrices in GL_{2}(R_{≥0}) that freely generate. American Mathematical Monthly, 122(8):790–792, October 2015.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.8.790.

Lengvarszky:2015:PI

[842] Zsolt Lengvárszky. Pythagoras by integral. American Mathematical Monthly, 122(8):792, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.8.792.

Titkos:2015:SPL

[843] Tamás Titkos. A simple proof of the Lebesgue decomposition theorem. American Mathematical Monthly, 122(8):793–794, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.8.793.

Tandra:2015:NPC

[844] Haryono Tandra. A new proof of the change of variable theorem for the Riemann integral. American Mathematical Monthly, 122(8):795– 799, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.8.795. See corrigendum [1045].

Borwein:2015:PPA

[845] Jonathan M. Borwein and Scott T. Chapman. I prefer pi: Addenda. American Mathematical Monthly, 122(8):800, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.

122.8.800. See [735]. The addenda correct formulas (4) and (12), references [34] and [74], and add a new reference.

Anonymous:2015:PSh

[846] Anonymous. Problems and solutions. American Mathematical Monthly, 122(8):801–808, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.8.801.

Tunno:2015:BRR

[847] Ferebee Tunno. Book review: Reviews. American Mathematical Monthly, 122(8):809–812, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.8.809.

Titkos:2015:HME

[848] Tamás Titkos. How to make equivalent measures? American Mathematical Monthly, 122(8):812, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.8.812.

Anonymous:2015:BMh

[849] Anonymous. Back matter. American Mathematical Monthly, 122(8): ??, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.8.bm.

Anonymous:2015:FMi

[850] Anonymous. Front matter. American Mathematical Monthly, 122 (9):813–814, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.9.fm.

Lu:2015:SS

[851] Zhiqin Lu and Julie Rowlett. The sound of symmetry. American Mathematical Monthly, 122(9):815–835, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.122.9.815.

Hajnal:2015:CPC

[852] Péter Hajnal, László I. Szabó, and Vilmos Totik. Convex polygons and common transversals. American Mathematical Monthly, 122(9):836–844, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.9.836.

Andreescu:2015:CNP

[853] Titu Andreescu, Vladimir Georgiev, and Oleg Mushkarov. Correction to ”Napoleon Polygons”. American Mathematical Monthly, 122(9):844, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.9.844.

Ren:2015:CBF

[854] Qingchun Ren, Jürgen Richter-Gebert, and Bernd Sturmfels. Cayley– Bacharach formulas. American Mathematical Monthly, 122(9):845–854, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.9.845.

Hasfura-Buenaga:2015:KDA

[855] J. Roberto Hasfura-Buenaga. Kronecker’s Diophantine approximation and the asymptotics of solutions of difference equations. American Mathematical Monthly, 122(9):855–861, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.122.9.855.

Nitecki:2015:CSS

[856] Zbigniew Nitecki. Cantorvals and subsum sets of null sequences. American Mathematical Monthly, 122(9):862–870, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.9.862.

Stanton:2015:SCU

[857] Robert O. Stanton. Sets closed under the division algorithm. American Mathematical Monthly, 122(9):871–879, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.9.871.

Andrews:2015:CNA

[858] George E. Andrews, Aviezri S. Fraenkel, and James A. Sellers. Characterizing the number of m-ary partitions modulo m. American Mathematical Monthly, 122(9):880–885, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.9.880.

Osler:2015:PNR

[859] Thomas J. Osler. A product of nested radicals for the AGM. American Mathematical Monthly, 122(9):886–887, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.9.886.

delaLlave:2015:CSI

[860] Rafael de la Llave. Computing some integrals by computing derivatives. American Mathematical Monthly, 122(9):888–889, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.9.888.

Ryabogin:2015:LNS

[861] Dmitry Ryabogin. A lemma of Nakajima and Süss on convex bodies. American Mathematical Monthly, 122(9):890–892, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.9.890.

Lampakis:2015:NPE

[862] Elias Lampakis. A new proof of Euler’s inequality. American Mathematical Monthly, 122(9):892, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.9.892.

Hoang:2015:LCT

[863] Nguyen S. Hoang. A limit comparison test for general series. American Mathematical Monthly, 122(9):893–896, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.122.9.893.

Daneshpajouh:2015:NPR

[864] Hamid Reza Daneshpajouh. A new proof of ℵ_{0}-resolvability of a metric space without an isolated point. American Mathematical Monthly, 122(9):897–898, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.9.897.

Anonymous:2015:PSi

[865] Anonymous. Problems and solutions. American Mathematical Monthly, 122(9):899–906, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.9.899.

Hartshorne:2015:BRR

[866] Robin Hartshorne. Book review: Reviews. American Mathematical Monthly, 122(9):907–911, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.9.907.

Henle:2015:BRR

[867] Michael Henle. Book review: Reviews. American Mathematical Monthly, 122(9):912–915, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.9.912.

Anonymous:2015:SO

[868] Anonymous. Statement of ownership. American Mathematical Monthly, 122(9):916, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.9.916.

Anonymous:2015:BMi

[869] Anonymous. Back matter. American Mathematical Monthly, 122(9):??, November 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.9.bm.

Anonymous:2015:FMj

[870] Anonymous. Front matter. American Mathematical Monthly, 122 (10):917–918, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.10.fm.

Shen:2015:EMC

[871] Zhaiming Shen and Lasse Rempe-Gillen. The exponential map is chaotic: An invitation to transcendental dynamics. American Mathematical Monthly, 122(10):919–940, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.10.919.

Kuczmarski:2015:HAN

[872] Fred Kuczmarski and James Kuczmarski. Hanging around in nonuniform fields. American Mathematical Monthly, 122(10):941–957, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL https://www.jstor.org/stable/10.4169/amer.
math.monthly.122.10.941.

Nowicki:2015:SDL

[873] Andrzej Nowicki. Strong divisibility and LCM-sequences. American Mathematical Monthly, 122(10):958–966, December 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.10.958.

Castro:2015:CCW

[874] Jesús Jerónimo Castro. On convex curves which have many inscribed triangles of maximum area. American Mathematical Monthly, 122 (10):967–971, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.10.967.

Cioaba:2015:GPP

[875] Sebastian M. Cioaba and Peter J. Cameron. A graph partition problem. American Mathematical Monthly, 122(10):972–982, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.10.972.

Gaspar:2015:ATO

[876] Jaime Gaspar. All triangles at once. American Mathematical Monthly, 122(10):982, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.10.982.

Bezhanishvili:2015:WDL

[877] Guram Bezhanishvili and David Pengelley. What does “less than or equal” really mean? American Mathematical Monthly, 122(10):983–989, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.10.983.

Tandra:2015:SPU

[878] Haryono Tandra. A simple proof of the uniform continuity of realvalued continuous functions on compact intervals. American Mathematical Monthly, 122(10):989, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.10.989.

Kilic:2015:PUG

[879] Mehmet Kiliç and Sahin Koçak. On products of uniquely geodesic spaces. American Mathematical Monthly, 122(10):990–994, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.122.10.990.

Hajja:2015:EPT

[880] Sadi Abu-Saymeh, Mowaffaq Hajja, and Hellmuth Stachel. Equicevian points of a triangle. American Mathematical Monthly, 122(10): 995–1000, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.10.995.

Peng:2015:GLT

[881] Weimin Peng. A generalized Liouville theorem for entire functions. American Mathematical Monthly, 122(10):1001–1002, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.10.1001.

Yamamori:2015:YAP

[882] Atsushi Yamamori. Yet another proof of Poincaré’s theorem. American Mathematical Monthly, 122(10):1003–1004, December 2015. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.10.1003.

Lange:2015:PRW

[883] Kenneth Lange. Polya’s random walk theorem revisited. American Mathematical Monthly, 122(10):1005–1007, December 2015. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.122.10.1005.

Cereceda:2015:NIP

[884] José Luis Cereceda. Newton’s interpolation polynomial for the sums of powers of integers. American Mathematical Monthly, 122(10):1007, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.10.1007.

Zhu:2015:SFP

[885] Jun-Ming Zhu. A semi-finite proof of Jacobi’s triple product identity. American Mathematical Monthly, 122(10):1008–1009, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
122.10.1008.

Anonymous:2015:PSj

[886] Anonymous. Problems and solutions. American Mathematical Monthly, 122(10):1010–1017, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.122.10.1010.

Harris:2015:BRR

[887] Michael Harris. Book review: Reviews. American Mathematical Monthly, 122(10):1018–1022, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.122.10.1018.

Anonymous:2015:EEb

[888] Anonymous. Editor’s endnotes. American Mathematical Monthly, 122 (10):1023–1025, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.122.10.1023.

Anonymous:2015:MR

[889] Anonymous. Monthly referees for 2015. American Mathematical Monthly, 122(10):1026–1028, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.122.10.1026.

Anonymous:2015:BMj

[890] Anonymous. Back matter. American Mathematical Monthly, 122(10):??, December 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.10.bm.

Anonymous:2016:FMa

[891] Anonymous. Front matter. American Mathematical Monthly, 123(1): 1–2, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.1.fm.

Neidinger:2016:FBG

[892] Richard D. Neidinger. A fair-bold gambling function is simply singular. American Mathematical Monthly, 123(1):3–18, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.1.3.

Kassel:2016:LRS

[893] Adrien Kassel and David B. Wilson. The looping rate and sandpile density of planar graphs. American Mathematical Monthly, 123(1):19– 39, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.1.19.

Cid:2016:IPS

[894] J. Ángel Cid and Rodrigo López Pouso. Integration by parts and by substitution unified with applications to Green’s theorem and uniqueness for ODEs. American Mathematical Monthly, 123(1):40–52, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.1.40.

Murty:2016:GET

[895] M. Ram Murty and Chester Weatherby. A generalization of Euler’s theorem for ζ(2k). American Mathematical Monthly, 123(1):53–65, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.1.53.

Shivarajkumar:2016:BSC

[896] Shivarajkumar. Beyond Schur’s conjecture. American Mathematical Monthly, 123(1):66–70, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.1.66.

Shitov:2016:DCG

[897] Yaroslav Shitov. The diameter of chess grows exponentially. American Mathematical Monthly, 123(1):71–77, January 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.1.71.

Velleman:2016:MG

[898] Daniel J. Velleman. Monthly gems. American Mathematical Monthly, 123 (1):77, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.1.77.

Slattery:2016:PMG

[899] Michael C. Slattery. On a property motivated by groups with a specified number of subgroups. American Mathematical Monthly, 123(1): 78–81, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.1.78.

Viro:2016:PCI

[900] Oleg Viro. Pictorial calculus for isometries. American Mathematical Monthly, 123(1):82–87, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.1.82.

Gaitanas:2016:PSR

[901] Konstantinos Gaitanas. A partial solution of representing 1 as a sum of n distinct unit fractions. American Mathematical Monthly, 123(1):87, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.1.87.

Radcliffe:2016:FSE

[902] David G. Radcliffe. A function that is surjective on every interval. American Mathematical Monthly, 123(1):88–89, January 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.1.88.

Osler:2016:ILC

[903] Thomas Osler. Iterations for the lemniscate constant resembling the Archimedean algorithm for pi. American Mathematical Monthly, 123(1): 90–93, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.1.90.

Kataria:2016:PPM

[904] Kuldeep Kumar Kataria. A probabilistic proof of the multinomial theorem. American Mathematical Monthly, 123(1):94–96, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.1.94.

Li:2016:SAU

[905] Xiaoxue Li. A synthetic approach to a USAMO problem. American Mathematical Monthly, 123(1):96, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.123.1.96.

Anonymous:2016:PSa

[906] Anonymous. Problems and solutions. American Mathematical Monthly, 123(1):97–105, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.1.97.

Glass:2016:BRR

[907] Darren B. Glass. Book review: Reviews. American Mathematical Monthly, 123(1):106–112, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.1.106.

Sandor:2016:NLM

[908] József Sándor. A note on the logarithmic mean. American Mathematical Monthly, 123(1):112, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.1.112.

Anonymous:2016:BMa

[909] Anonymous. Back matter. American Mathematical Monthly, 123(1): ??, January 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.1.bm.

Anonymous:2016:FMb

[910] Anonymous. Front matter. American Mathematical Monthly, 123 (2):113–114, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.2.fm.

Motta:2016:OTT

[911] Francis C. Motta, Patrick D. Shipman, and Bethany D. Springer. Optimally topologically transitive orbits in discrete dynamical systems. American Mathematical Monthly, 123(2):115–135, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.2.115.

Morrison:2016:CYP

[912] Ian Morrison and David Swinarski. Can you play a fair game of craps with a loaded pair of dice? American Mathematical Monthly, 123 (2):136–148, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.2.136.

Kent:2016:GSC

[913] Deborah A. Kent and David J. Muraki. A geometric solution of a cubic by Omar Khayyam … in which colored diagrams are used instead of letters for the greater ease of learners. American Mathematical Monthly, 123(2):149–160, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.2.149.

Suarez-Alvarez:2016:CER

[914] Mariano Suárez-Álvarez. Cyclic extensions are radical. American Mathematical Monthly, 123(2):160, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.2.160.

Hiller:2016:OFU

[915] Josh Hiller. Old friends in unexpected places: Pascal (and other) matrices in GL_{n}(C). American Mathematical Monthly, 123(2):161–167, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.2.161.

Brillhart:2016:UCS

[916] John Brillhart, Richard Blecksmith, and Mike Decaro. Using conic sections to factor integers. American Mathematical Monthly, 123(2):168– 174, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.2.168.

Spivey:2016:PPB

[917] Michael Z. Spivey. Probabilistic proofs of a binomial identity, its inverse, and generalizations. American Mathematical Monthly, 123(2):175–180, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.2.175.

Dekking:2016:HKS

[918] E. W. Dekking and F. M. Dekking. Helge von Koch’s snowflake curve revisited. American Mathematical Monthly, 123(2):181–184, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.2.181.

Baker:2016:SCD

[919] Jonathan Baker. Strong convexity does not imply radial unboundedness. American Mathematical Monthly, 123(2):185–188, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.2.185.

Sun:2016:STP

[920] Xin-Jie Sun. On sums of three pentagonal numbers. American Mathematical Monthly, 123(2):189–191, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.123.2.189.

Lutgen:2016:CPA

[921] Jeffrey Lutgen. A combinatorial proof of an alternating convolution identity for multichoose numbers. American Mathematical Monthly, 123(2):192–196, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.2.192.

Anonymous:2016:PSb

[922] Anonymous. Problems and solutions. American Mathematical Monthly, 123(2):197–204, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.2.197.

Nunemacher:2016:BRR

[923] Jeffrey Nunemacher. Book review: Reviews. American Mathematical Monthly, 123(2):205–208, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.2.205.

Anonymous:2016:BMb

[924] Anonymous. Back matter. American Mathematical Monthly, 123(2): ??, February 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.2.bm.

Anonymous:2016:FMc

[925] Anonymous. Front matter. American Mathematical Monthly, 123 (3):209–210, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.3.fm.

Quinn:2016:YGG

[926] Jennifer J. Quinn. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2016 to George Berzsenyi for distinguished service to mathematics. American Mathematical Monthly, 123(3):211–213, March 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.3.211.

Parshall:2016:MPR

[927] Karen Hunger Parshall. Mathematics and the politics of race: The case of William Claytor (Ph.D., University of Pennsylvania, 1933). American Mathematical Monthly, 123(3):214–240, March 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.3.214.

Boyd:2016:CTE

[928] John P. Boyd. Correcting three errors in Kantorovich & Krylov’s approximate methods of higher analysis. American Mathematical Monthly, 123(3):241–257, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.3.241.

Bekker:2016:MLN

[929] Boris M. Bekker, Oleg A. Ivanov, and Alexander S. Merkurjev. Milnor’s lemma, Newton’s method, and continued fractions. American Mathematical Monthly, 123(3):258–266, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.3.258.

Kimberling:2016:SBS

[930] Clark Kimberling and Kenneth B. Stolarsky. Slow Beatty sequences, devious convergence, and partitional divergence. American Mathematical Monthly, 123(3):267–273, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.3.267.

Butler:2016:IPS

[931] Steve Butler, Ron Graham, and Richard Stong. Inserting plus signs and adding. American Mathematical Monthly, 123(3):274–279, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.3.274.

Mertens:2016:EMU

[932] Tom Mertens and Jeff Weeks. Electromagnetism in a multiconnected universe. American Mathematical Monthly, 123(3):280–286, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.3.280.

Ovall:2016:LME

[933] Jeffrey S. Ovall. The Laplacian and mean and extreme values. American Mathematical Monthly, 123(3):287–291, March 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.3.287.

McGrath:2016:SJB

[934] Peter McGrath. On the smooth Jordan Brouwer separation theorem. American Mathematical Monthly, 123(3):292–295, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.3.292.

Ramirez:2016:ICG

[935] José L. Ramírez. An identity of Carlitz and its generalization. American Mathematical Monthly, 123(3):295, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.123.3.295.

Anonymous:2016:PSc

[936] Anonymous. Problems and solutions. American Mathematical Monthly, 123(3):296–303, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.3.296.

Jackson:2016:BRR

[937] Craig H. Jackson. Book review: Reviews. American Mathematical Monthly, 123(3):304–308, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.3.304.

Anonymous:2016:BMc

[938] Anonymous. Back matter. American Mathematical Monthly, 123(3): ??, March 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.3.bm.

Anonymous:2016:FMd

[939] Anonymous. Front matter. American Mathematical Monthly, 123(4): 309–310, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.4.fm.

Cahen:2016:WYS

[940] Paul-Jean Cahen and Jean-Luc Chabert. What you should know about integer-valued polynomials. American Mathematical Monthly, 123(4): 311–337, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.4.311.

Johnson:2016:RPT

[941] Keith Johnson and Kira Scheibelhut. Rational polynomials that take integer values at the Fibonacci numbers. American Mathematical Monthly, 123(4):338–346, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.4.338.

Brookfield:2016:FF

[942] Gary Brookfield. Factoring forms. American Mathematical Monthly, 123(4):347–362, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.4.347.

Cairns:2016:CLP

[943] Grant Cairns, Yuri Nikolayevsky, and Gavin Rossiter. Conewise linear periodic maps of the plane with integer coefficients. American Mathematical Monthly, 123(4):363–375, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.4.363.

Ayad:2016:WDG

[944] Mohamed Ayad, Omar Kihel, and Jesse Larone. When does a given polynomial with integer coefficients divide another? American Mathematical Monthly, 123(4):376–381, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.4.376.

Calkin:2016:WMC

[945] Neil J. Calkin, Beth Novick, and Hayato Ushijima-Mwesigwa. What Moser could have asked: Counting Hamilton cycles in tournaments. American Mathematical Monthly, 123(4):382–386, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.4.382.

Luzia:2016:PPT

[946] Nuno Luzia. A proof of the Pythagorean Theorem after Descartes. American Mathematical Monthly, 123(4):386, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.123.4.386.

Costin:2016:SGR

[947] Ovidiu Costin, Neil Falkner, and Jeffery D. McNeal. Some generalizations of the Riemann–Lebesgue lemma. American Mathematical Monthly, 123(4):387–391, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.4.387.

Landweber:2016:FDC

[948] Peter S. Landweber, Emanuel A. Lazar, and Neel Patel. On fiber diameters of continuous maps. American Mathematical Monthly, 123(4): 392–397, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.4.392.

Lopez-Aguayo:2016:IM

[949] Daniel López-Aguayo. Irrationality via maximality. American Mathematical Monthly, 123(4):397, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.4.397.

Bernstein:2016:LSS

[950] Dennis S. Bernstein, Khaled F. Aljanaideh, and Arthur E. Frazho. Laurent series and l^{p} sequences. American Mathematical Monthly, 123(4): 398, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.4.398.

Anonymous:2016:PSd

[951] Anonymous. Problems and solutions. American Mathematical Monthly, 123(4):399–406, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.4.399.

Garrity:2016:BRR

[952] Thomas Garrity. Book review: Reviews. American Mathematical Monthly, 123(4):407–412, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.4.407.

Anonymous:2016:BMd

[953] Anonymous. Back matter. American Mathematical Monthly, 123(4): ??, April 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.4.bm.

Anonymous:2016:FMe

[954] Anonymous. Front matter. American Mathematical Monthly, 123(5): 413–414, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.5.fm.

Eilers:2016:LCP

[955] Søren Eilers. The LEGO counting problem. American Mathematical Monthly, 123(5):415–426, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.5.415.

Boas:2016:MF

[956] Harold P. Boas. Mocposite functions. American Mathematical Monthly, 123(5):427–438, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.5.427.

Apostol:2016:NLS

[957] Tom M. Apostol and Mamikon A. Mnatsakanian. A new look at surfaces of constant curvature. American Mathematical Monthly, 123(5):439–447, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.5.439.

Cruz-Sampedro:2016:HRC

[958] Jaime Cruz-Sampedro and Margarita Tetlalmatzi-Montiel. Hardy’s reduction for a class of Liouville integrals of elementary functions. American Mathematical Monthly, 123(5):448–470, May 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.5.448.

Josevich:2016:AAP

[959] Piotr Josevich. An alternative approach to the product rule. American Mathematical Monthly, 123(5):470, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.5.470.

Mullen:2016:FSM

[960] G. L. Mullen, D. Panario, and D. Thomson. Fast and simple modular interpolation using factorial representation. American Mathematical Monthly, 123(5):471–480, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.5.471.

Pinelis:2016:MSR

[961] Iosif Pinelis. On measurable semigroups in R. American Mathematical Monthly, 123(5):481, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.5.481.

Sheydvasser:2016:CUS

[962] Arseniy Sheydvasser. A corrigendum to unreasonable slightness. American Mathematical Monthly, 123(5):482–485, May 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.5.482. See

[200].

Dilcher:2016:ESL

[963] Karl Dilcher and Christophe Vignat. Euler and the strong law of small numbers. American Mathematical Monthly, 123(5):486–490, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.5.486.

Hsu:2016:RNP

[964] Tim Hsu and San José. Rational nonaxis points on the unit circle have irrational angles. American Mathematical Monthly, 123(5):490, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.5.490.

Pinelis:2016:EAD

[965] Iosif Pinelis. Explicit additive decomposition of norms on R^{2}. American Mathematical Monthly, 123(5):491–496, May 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.5.491.

Gaul:2016:PAP

[966] Michael Gaul and Fred Kuczmarski. Parabolas and Archimedes’ 2/3property. American Mathematical Monthly, 123(5):497–501, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.5.497.

McGrath:2016:ESP

[967] Peter McGrath. An extremely short proof of the hairy ball theorem. American Mathematical Monthly, 123(5):502–503, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.5.502.

Anonymous:2016:PSe

[968] Anonymous. Problems and solutions. American Mathematical Monthly, 123(5):504–511, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.5.504.

Alexanderson:2016:BRR

[969] Gerald L. Alexanderson. Book review: Reviews. American Mathematical Monthly, 123(5):512–514, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.5.512.

Anonymous:2016:EEN

[970] Anonymous. Editor’s end notes. American Mathematical Monthly, 123(5):515–516, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.5.515.

Anonymous:2016:BMe

[971] Anonymous. Back matter. American Mathematical Monthly, 123(5):??, May 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.5.bm.

Anonymous:2016:FMf

[972] Anonymous. Front matter. American Mathematical Monthly, 123(6): 517–518, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.6.fm.

Johnson:2016:A

[973] Jennifer M. Johnson and János Kollár. Arcology. American Mathematical Monthly, 123(6):519–541, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.6.519.

Butler:2016:MFH

[974] Steve Butler, Persi Diaconis, and Ron Graham. The mathematics of the flip and horseshoe shuffles. American Mathematical Monthly, 123(6):542–556, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.6.542.

Angelis:2016:WC

[975] Valerio De Angelis and Dominic Marcello. Wilf’s conjecture. American Mathematical Monthly, 123(6):557–573, June 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.6.557.

Tojo:2016:HAP

[976] F. Adrián F. Tojo. A hyperbolic analog of the phasor addition formula. American Mathematical Monthly, 123(6):574–582, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.6.574.

Panzone:2016:TPD

[977] Pablo A. Panzone. Tiling the plane with different hexagons and triangles. American Mathematical Monthly, 123(6):583–591, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.6.583.

Kenter:2016:HWY

[978] Franklin H. J. Kenter. How to win your betting pool with Jensen’s inequality and the law of large numbers. American Mathematical Monthly, 123(6):592–596, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.6.592.

Gurel:2016:OPS

[979] Erhan Gürel. On the occurrence of perfect squares among values of certain polynomial products. American Mathematical Monthly, 123(6): 597–599, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.6.597.

Swan:2016:ENV

[980] Hunter Swan. Expected number of vertices of a hypercube slice. American Mathematical Monthly, 123(6):600–604, June 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.6.600.

McCauley:2016:PUL

[981] Peter McCauley. Proofs without using limits of both calculus power rules. American Mathematical Monthly, 123(6):604, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.6.604.

Khanh:2016:QLP

[982] Pham Duy Khanh and Marc Lassonde. Quasiconvex linear perturbations and convexity. American Mathematical Monthly, 123(6):605–608, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.6.605.

Amir:2016:SNE

[983] Ariel Amir, Mikhail Lemeshko, and Tadashi Tokieda. Surprises in numerical expressions of physical constants. American Mathematical Monthly, 123(6):609–612, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.6.609.

Anonymous:2016:PSf

[984] Anonymous. Problems and solutions. American Mathematical Monthly, 123(6):613–620, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.6.613.

Hogben:2016:BRR

[985] Leslie Hogben and Mark Hunacek. Book review: Reviews. American Mathematical Monthly, 123(6):621–624, June 2016. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.123.6.621.

Anonymous:2016:BMf

[986] Anonymous. Back matter. American Mathematical Monthly, 123(6):??, June 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.6.bm.

Anonymous:2016:FMg

[987] Anonymous. Front matter. American Mathematical Monthly, 123 (7):625–626, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.7.fm.

Florescu:2016:RRW

[988] Laura Florescu, Lionel Levine, and Yuval Peres. The range of a rotor walk. American Mathematical Monthly, 123(7):627–642, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.7.627.

Zalcman:2016:TTT

[989] Lawrence Zalcman. A tale of three theorems. American Mathematical Monthly, 123(7):643–656, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.7.643.

Balazs:2016:ENN

[990] Márton Balázs and Áron Folly. An electric network for nonreversible Markov chains. American Mathematical Monthly, 123(7):657–682, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.7.657.

Coll:2016:CAP

[991] Vincent E. Coll and Jeff Dodd. A characteristic averaging property of the catenary. American Mathematical Monthly, 123(7):683–688, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.7.683.

Ivanov:2016:SCT

[992] Dennis Ivanov and Serge Tabachnikov. The six circles theorem revisited. American Mathematical Monthly, 123(7):689–698, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.7.689.

Avron:2016:CFP

[993] Arnon Avron and Nachum Dershowitz. Cayley’s formula: A page from the book. American Mathematical Monthly, 123(7):699–700, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.7.699.

Hardy:2016:TSI

[994] Michael Hardy. On tangents and secants of infinite sums. American Mathematical Monthly, 123(7):701–703, August 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.7.701.

Kasman:2016:FMD

[995] Alex Kasman. Factorization of a matrix differential operator using functions in its kernel. American Mathematical Monthly, 123(7):704–709, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.7.704.

Sittinger:2016:CUT

[996] Brian D. Sittinger. Computing ζ(2m) by using telescoping sums. American Mathematical Monthly, 123(7):710–715, August 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.7.710.

Yao:2016:GIF

[997] Yuanyuan Yao and Wenxia Li. Generating iterated function systems for the Vicsek snowflake and the Koch curve. American Mathematical Monthly, 123(7):716–721, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.7.716.

Anonymous:2016:PSg

[998] Anonymous. Problems and solutions. American Mathematical Monthly, 123(7):722–730, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.7.722.

Richey:2016:BR

[999] Matthew Richey. Book review. American Mathematical Monthly, 123(7):731–736, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.7.731.

Anonymous:2016:BMg

[1000] Anonymous. Back matter. American Mathematical Monthly, 123(7): ??, August 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.7.bm.

Anonymous:2016:FMh

[1001] Anonymous. Front matter. American Mathematical Monthly, 123 (8):737–738, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.8.fm.

Klosinski:2016:SSW

[1002] Leonard F. Klosinski, Gerald L. Alexanderson, and Mark Krusemeyer. The Seventy-Sixth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 123(8):739–752, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.8.739.

Lagarias:2016:EKP

[1003] Jeffrey C. Lagarias. Erdős, Klarner, and the 3x + 1 problem. American Mathematical Monthly, 123(8):753–776, October 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.8.753.

Bajpai:2016:APF

[1004] Dvij Bajpai and Daniel J. Velleman. Anonymity in predicting the future. American Mathematical Monthly, 123(8):777–788, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.8.777.

Caldero:2016:ETF

[1005] Philippe Caldero and Jérôme Germoni. Equilateral triangles and the Fano plane. American Mathematical Monthly, 123(8):789–801, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.8.789.

Michael:2016:AGT

[1006] T. S. Michael and Val Pinciu. The art gallery theorem, revisited. American Mathematical Monthly, 123(8):802–807, October 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.8.802.

Mitchell:2016:ANA

[1007] Lon Mitchell, Michael A. Jones, and Brittany Shelton. Abelian and non-Abelian triangle mysteries. American Mathematical Monthly, 123 (8):808–813, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.8.808.

Kong:2016:AAP

[1008] Yong Kong and Shaowei Zhang. The adventitious angles problem: The lonely fractional derived angle. American Mathematical Monthly, 123(8):814–816, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.8.814.

Jeronimo-Castro:2016:MIC

[1009] J. Jerónimo-Castro and C. Yee-Romero. Maximal isoptic chords of convex curves. American Mathematical Monthly, 123(8):817–820, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.8.817.

deLaGrandville:2016:ADE

[1010] Olivier de La Grandville. An alternate derivation of the Euler–Poisson equation. American Mathematical Monthly, 123(8):821–824, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.8.821.

Grzesik:2016:YAA

[1011] J. A. Grzesik. Yet another analytic proof of Napoleon’s Theorem. American Mathematical Monthly, 123(8):824, October 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.8.824.

Farkas:2016:VBL

[1012] Bálint Farkas and Sven-Ake Wegner. Variations on Barbalat’s Lemma. American Mathematical Monthly, 123(8):825–830, October 2016. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.8.825.

Anonymous:2016:PSh

[1013] Anonymous. Problems and solutions. American Mathematical Monthly, 123(8):831–839, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.8.831.

Kidwell:2016:BR

[1014] Mark Kidwell. Book review. American Mathematical Monthly, 123 (8):840–843, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.8.840.

Mutuk:2016:PPC

[1015] Halil Mutuk. A physical proof for Cauchy’s inequality. American Mathematical Monthly, 123(8):843, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.8.843.

Anonymous:2016:BMh

[1016] Anonymous. Back matter. American Mathematical Monthly, 123(8): ??, October 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.8.bm.

Anonymous:2016:FMi

[1017] Anonymous. Front matter. American Mathematical Monthly, 123 (9):845–846, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.9.fm.

Chapman:2016:LEJ

[1018] Scott T. Chapman. A letter from the Editor: Jonathan M. Borwein (1951–2016). American Mathematical Monthly, 123(9):847–848, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.9.847.

Baginski:2016:FEG

[1019] Paul Baginski, Ryan Rodriguez, George J. Schaeffer, and Yiwei She. Finding elements with given factorization lengths and multiplicities. American Mathematical Monthly, 123(9):849–870, November 2016. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.9.849.

Ramshaw:2016:SHE

[1020] Lyle Ramshaw. Stråhle’s equal-tempered triumph. American Mathematical Monthly, 123(9):871–883, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.9.871.

Cortissoz:2016:NHF

[1021] Jean C. Cortissoz. A note on harmonic functions on surfaces. American Mathematical Monthly, 123(9):884–893, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.9.884.

Ninjbat:2016:AMM

[1022] Uuganbaatar Ninjbat. Another memorable Mongolian Olympiad problem: A two-colored rectangular lattice. American Mathematical Monthly, 123(9):894–899, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.9.894.

Kramer:2016:MFP

[1023] Josh Brown Kramer, Jonathan Cutler, and A. J. Radcliffe. The multistep friendship paradox. American Mathematical Monthly, 123 (9):900–908, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.9.900.

Eastman:2016:VSR

[1024] Sean Eastman, Selwyn Hollis, Kawee Numpacharoen, and Jared Schlieper. The volume of the spatial region corresponding to n×n correlation matrices. American Mathematical Monthly, 123(9):909–918, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.9.909.

Martinjak:2016:TES

[1025] Ivica Martinjak. Two extensions of the Sury’s identity. American Mathematical Monthly, 123(9):919, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.9.919.

Schomburg:2016:POD

[1026] Bernd Schomburg. Primitivity and orbit decomposition. American Mathematical Monthly, 123(9):920–923, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.123.9.920.

Dudek:2016:SME

[1027] Adrian W. Dudek. On the success of mishandling Euclid’s lemma. American Mathematical Monthly, 123(9):924–927, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.9.924.

Hoffman:2016:SPG

[1028] Alan J. Hoffman and Chai Wah Wu. A simple proof of a generalized Cauchy–Binet theorem. American Mathematical Monthly, 123 (9):928–930, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.9.928.

Anghel:2016:OBR

[1029] Nicolae Anghel. Optimal bounds for the roots of polynomials. American Mathematical Monthly, 123(9):931–935, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.9.931.

Bevelacqua:2016:FNE

[1030] Anthony J. Bevelacqua. A family of non-Euclidean PIDs. American Mathematical Monthly, 123(9):936–939, November 2016. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.123.9.936.

Kataria:2016:APB

[1031] Kuldeep Kumar Kataria. An alternate proof of the binomial theorem. American Mathematical Monthly, 123(9):940, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.123.9.940.

Anonymous:2016:PSi

[1032] Anonymous. Problems and solutions. American Mathematical Monthly, 123(9):941–948, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.9.941.

Folland:2016:BRR

[1033] Gerald B. Folland. Book review. American Mathematical Monthly, 123 (9):949–952, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.9.949.

Anonymous:2016:BMi

[1034] Anonymous. Back matter. American Mathematical Monthly, 123(9):??, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.9.bm.

Anonymous:2016:FM

[1035] Anonymous. Front matter. American Mathematical Monthly, 123 (10):953–954, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.10.fm.

Chapman:2016:LEL

[1036] Scott T. Chapman. A letter from the Editor: The long and winding road. American Mathematical Monthly, 123(10):955–958, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.10.955.

Bradshaw:2016:SDN

[1037] Franco Bradshaw. Single digit NFL scores. American Mathematical Monthly, 123(10):959, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.10.959.

Geroldinger:2016:SL

[1038] Alfred Geroldinger. Sets of lengths. American Mathematical Monthly, 123(10):960–988, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.10.960.

Klyachko:2016:BF

[1039] Anton A. Klyachko and Anton N. Vassilyev. Balanced factorizations. American Mathematical Monthly, 123(10):989–1000, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.10.989.

Nathanson:2016:CPF

[1040] Melvyn B. Nathanson. Cantor polynomials and the Fueter–Pólya theorem. American Mathematical Monthly, 123(10):1001–1012, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.10.1001.

Broussard:2016:OVG

[1041] Aaron M. Broussard, Martin E. Malandro, and Abagayle Serreyn. Optimizing the video game multi-jump: Player strategy, AI, and level design. American Mathematical Monthly, 123(10):1013–1032, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.10.1013.

Lagarias:2016:DFF

[1042] Jeffrey C. Lagarias, Takumi Murayama, and D. Harry Richman. Dilated floor functions that commute. American Mathematical Monthly, 123 (10):1033–1038, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.10.1033.

Garcia:2016:QFN

[1043] Stephan Ramon Garcia and Florian Luca. Quotients of Fibonacci numbers. American Mathematical Monthly, 123(10):1039–1044, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.123.10.1039.

Ehrenborg:2016:PCD

[1044] Richard Ehrenborg and Alex Happ. Parking cars of different sizes. American Mathematical Monthly, 123(10):1045–1048, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
123.10.1045.

Anonymous:2016:CNP

[1045] Anonymous. Corrigendum to ”A New Proof of the Change of Variable Theorem for the Riemann Integral”. American Mathematical Monthly, 123(10):1049, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.10.1049. See [844].

Anonymous:2016:PS

[1046] Anonymous. Problems and solutions. American Mathematical Monthly, 123(10):1050–1057, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.10.1050.

Marx:2016:BRF

[1047] Chris A. Marx. Book review: A First Course in the Calculus of Variations. By Mark Kot. Student Mathematical Library Vol 72, American Mathematical Society, Providence, Rhode Island, 2014, x + 298 pp., ISBN 978-1-4704-1495-5, $50. American Mathematical Monthly, 123 (10):1058–1061, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.123.10.1058.

Anonymous:2016:MR

[1048] Anonymous. Monthly referees for 2016. American Mathematical Monthly, 123(10):1062–1066, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.123.10.1062.

Anonymous:2016:BM

[1049] Anonymous. Back matter. American Mathematical Monthly, 123(10):??, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.123.10.bm.

Anonymous:2017:FMa

[1050] Anonymous. Front matter. American Mathematical Monthly, 124(1): 1–2, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.1.fm.

Colley:2017:LE

[1051] Susan Jane Colley. Letter from the Editor. American Mathematical Monthly, 124(1):3, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.1.3.

Blair:2017:NKG

[1052] Ryan Blair, Marion Campisi, Jesse Johnson, Scott A. Taylor, and Maggy Tomova. Neighbors of knots in the Gordian graph. American Mathematical Monthly, 124(1):4–23, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.1.4.

Vanderbei:2017:FS

[1053] Robert J. Vanderbei. The falling slinky. American Mathematical Monthly, 124(1):24–36, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.124.1.24.

Cafure:2017:ICR

[1054] Antonio Cafure and Eda Cesaratto. Irreducibility criteria for reciprocal polynomials and applications. American Mathematical Monthly, 124(1):37–53, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.1.37.

Gallian:2017:SFY

[1055] Joseph A. Gallian. Seventy-five years of the Putnam mathematical competition*. American Mathematical Monthly, 124(1):54–59, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.1.54.

Ustuzhanin:2017:NWP

[1056] Saveliy Ustuzhanin. A novel way to prove the irrationality of \sqrt{2}. American Mathematical Monthly, 124(1):59, January 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.1.59.

Anghel:2017:PPR

[1057] Nicolae Anghel. On a 2015 Putnam problem related to a double recursion. American Mathematical Monthly, 124(1):60–69, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.1.60.

Stewart:2017:WDA

[1058] Ian Stewart. Why do all triangles form a triangle? American Mathematical Monthly, 124(1):70–73, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.1.70.

Lupu:2017:NPF

[1059] Cezar Lupu. A new proof of the Finsler–Hadwiger inequality. American Mathematical Monthly, 124(1):73, January 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.1.73.

Gurtas:2017:CPM

[1060] Yusuf Z. Gürtas. Chebyshev polynomials and the minimal polynomial of cos(2π/n). American Mathematical Monthly, 124(1):74–78, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.1.74.

Shamir:2017:TSU

[1061] Eli Shamir and Benny Sudakov. Two-sided, unbiased version of Hall’s marriage theorem. American Mathematical Monthly, 124(1):79–80, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.1.79.

vanMunster:2017:HH

[1062] Bart van Munster. Hausdorffization and homotopy. American Mathematical Monthly, 124(1):81–82, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.1.81.

Anonymous:2017:PSa

[1063] Anonymous. Problems and solutions. American Mathematical Monthly, 124(1):83–91, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.1.83.

Rosenhouse:2017:R

[1064] Jason Rosenhouse. Reviews. American Mathematical Monthly, 124(1): 92–95, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.1.92.

Anonymous:2017:BMa

[1065] Anonymous. Back matter. American Mathematical Monthly, 124(1): ??, January 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.1.bm.

Anonymous:2017:FMb

[1066] Anonymous. Front matter. American Mathematical Monthly, 124(2): 98, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.2.fm.

Crannell:2017:IS

[1067] Annalisa Crannell, Marc Frantz, and Fumiko Futamura. The image of a square. American Mathematical Monthly, 124(2):99–115, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.2.99.

Stenger:2017:EMM

[1068] Allen Stenger. Experimental math for math monthly problems. American Mathematical Monthly, 124(2):116–131, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.2.116.

Maltenfort:2017:CAS

[1069] Michael Maltenfort. Characterizing additive systems. American Mathematical Monthly, 124(2):132–148, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.124.2.132.

Libbus:2017:RMS

[1070] Bisharah Libbus, Gordon Simons, and Yi-Ching Yao. Rotating multiple sets of labeled points to bring them into close coincidence: A generalized Wahba problem. American Mathematical Monthly, 124 (2):149–160, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.2.149.

Ginebaugh:2017:GIP

[1071] Scott Ginebaugh. Generalized infinite products for powers of e^{1/k}. American Mathematical Monthly, 124(2):161–165, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.2.161.

Aebi:2017:SQR

[1072] Christian Aebi and Grant Cairns. Sums of quadratic residues and nonresidues. American Mathematical Monthly, 124(2):166–169, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.2.166.

Mercer:2017:APT

[1073] Idris Mercer. Another proof that there are infinitely many primes. American Mathematical Monthly, 124(2):169, February 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.2.169.

Panraksa:2017:NAR

[1074] Chatchawan Panraksa and Pornrat Ruengrot. A note on average of roots of unity. American Mathematical Monthly, 124(2):170–174, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.2.170.

Audet:2017:MAE

[1075] Charles Audet. Maximal area of equilateral small polygons. American Mathematical Monthly, 124(2):175–178, February 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.2.175.

Anonymous:2017:PSb

[1076] Anonymous. Problems and solutions. American Mathematical Monthly, 124(2):179–187, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.2.179.

Baker:2017:PM

[1077] Alan Baker. The philosophies of mathematics. American Mathematical Monthly, 124(2):188–192, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.2.188.

Anonymous:2017:BMb

[1078] Anonymous. Back matter. American Mathematical Monthly, 124(2): ??, February 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.2.bm.

Anonymous:2017:FMc

[1079] Anonymous. Front matter. American Mathematical Monthly, 124(3): 194, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.3.fm.

Crannell:2017:YGG

[1080] Annalisa Crannell and Doug Ensley. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2017 to Martha Siegel for Distinguished Service to Mathematics. American Mathematical Monthly, 124(3):195–197, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.3.195.

Clark:2017:ECI

[1081] Pete L. Clark. The Euclidean criterion for irreducibles. American Mathematical Monthly, 124(3):198–216, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.124.3.198.

Bak:2017:ECC

[1082] Joseph Bak and Strashimir G. Popvassilev. The evolution of Cauchy’s closed curve theorem and Newman’s simple proof. American Mathematical Monthly, 124(3):217–231, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.3.217.

Sondow:2017:PPN

[1083] Jonathan Sondow and Kieren MacMillan. Primary pseudoperfect numbers, arithmetic progressions, and the Erdős–Moser equation. American Mathematical Monthly, 124(3):232–240, March 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.3.232.

Momihara:2017:DSC

[1084] Koji Momihara and Masashi Shinohara. Distance sets on circles. American Mathematical Monthly, 124(3):241–254, March 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.3.241.

Weiss:2017:WRS

[1085] Michael Weiss. Where are the rational squares? American Mathematical Monthly, 124(3):255–259, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.3.255.

Mondal:2017:SMI

[1086] Shubhodip Mondal. Surjectivity of maps induced on matrices by polynomials and entire functions. American Mathematical Monthly, 124(3):260– 264, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.3.260.

Huylebrouck:2017:NRC

[1087] Dirk Huylebrouck. A new regular (compound) polyhedron (of infinite Kepler–Poinsot type). American Mathematical Monthly, 124(3):265–268, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL https://www.jstor.org/stable/10.4169/amer.
math.monthly.124.3.265.

Plante:2017:PBV

[1088] Joseph Plante. A proof of Bonnet’s version of the mean value theorem by methods of Cauchy. American Mathematical Monthly, 124(3):269– 273, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.3.269.

Anonymous:2017:PSc

[1089] Anonymous. Problems and solutions. American Mathematical Monthly, 124(3):274–281, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.3.274.

Lanphier:2017:BRS

[1090] Dominic Lanphier. Book review: Summing it Up: From One Plus One to Modern Number Theory. American Mathematical Monthly, 124(3):282–288, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.3.282.

Anonymous:2017:BMc

[1091] Anonymous. Back matter. American Mathematical Monthly, 124(3): ??, March 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.3.bm.

Anonymous:2017:FMd

[1092] Anonymous. Front matter. American Mathematical Monthly, 124(4): 290, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.4.fm.

Becker:2017:BSG

[1093] Paul E. Becker, Martin Derka, Sheridan Houghten, and Jennifer Ulrich. Build a sporadic group in your basement. American Mathematical Monthly, 124(4):291–305, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.4.291.

Boyd:2017:CPS

[1094] John P. Boyd and Houjun Wang. Convergent power series for boundary value problems and eigenproblems with application to atmospheric and oceanic tides. American Mathematical Monthly, 124(4):306–323, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.4.306.

Berger:2017:MGS

[1095] Arno Berger. More grade school triangles. American Mathematical Monthly, 124(4):324–336, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.124.4.324.

Maehara:2017:LT

[1096] Hiroshi Maehara and Horst Martini. On Lexell’s Theorem. American Mathematical Monthly, 124(4):337–344, April 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.4.337.

Aviles:2017:LAA

[1097] Antonio Avilés and Grzegorz Plebanek. A little ado about rectangles. American Mathematical Monthly, 124(4):345–350, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.4.345.

Wooley:2017:SEP

[1098] Trevor D. Wooley. A superpowered Euclidean prime generator. American Mathematical Monthly, 124(4):351–352, April 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.4.351.

McTague:2017:GCD

[1099] Carl McTague. On the greatest common divisor of binomial coefficients. American Mathematical Monthly, 124(4):353–356, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.4.353.

Sawhney:2017:TPG

[1100] Mehtaab S. Sawhney. A telescoping proof of the AM–GM inequality. American Mathematical Monthly, 124(4):356, April 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.4.356.

Speyer:2017:CPT

[1101] David E. Speyer. A counting proof of a theorem of Frobenius. American Mathematical Monthly, 124(4):357–359, April 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.4.357.

Easdale:2017:CST

[1102] Evelyn R. Easdale, Jolene E. Fleming, and Bogdan D. Suceava. Convergence for series with terms defined by a recurrence relation. American Mathematical Monthly, 124(4):360–364, April 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.4.360.

Elsholtz:2017:GCP

[1103] Christian Elsholtz. Golomb’s conjecture on prime gaps. American Mathematical Monthly, 124(4):365–368, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.4.365.

Anonymous:2017:PSd

[1104] Anonymous. Problems and solutions. American Mathematical Monthly, 124(4):369–378, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.4.369.

Cohen:2017:BRM

[1105] Marion Cohen. Book review: Mom the Chemistry Professor: Personal Accounts and Advice from Chemistry Professors Who Are Mothers. American Mathematical Monthly, 124(4):379–383, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.4.379.

Anonymous:2017:BMd

[1106] Anonymous. Back matter. American Mathematical Monthly, 124(4): ??, April 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.4.bm.

Anonymous:2017:FMe

[1107] Anonymous. Front matter. American Mathematical Monthly, 124(5):386, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.5.fm.

Deijfen:2017:FFS

[1108] Maria Deijfen, Alexander E. Holroyd, and James B. Martin. Friendly frogs, stable marriage, and the magic of invariance. American Mathematical Monthly, 124(5):387–402, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.5.387.

Monterde:2017:HEU

[1109] Juan Monterde and David Rochera. Holditch’s ellipse unveiled. American Mathematical Monthly, 124(5):403–421, May 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.5.403.

Marko:2017:SRT

[1110] Frantisek Marko and Semyon Litvinov. On the Steiner–Routh theorem for simplices. American Mathematical Monthly, 124(5):422–435, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.5.422.

Balcerzak:2017:LEN

[1111] Marek Balcerzak, Michal Poplawski, and Julia Wódka. Local extrema and nonopenness points of continuous functions. American Mathematical Monthly, 124(5):436–443, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.5.436.

Farrell:2017:ME

[1112] Jeremiah Farrell and William Johnston. Mathematical evolution. American Mathematical Monthly, 124(5):444, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.124.5.444.

Frenkel:2017:GCD

[1113] Péter E. Frenkel and József Pelikán. On the greatest common divisor of the value of two polynomials. American Mathematical Monthly, 124(5):446–450, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.5.446.

DeAngelis:2017:SDS

[1114] Valerio De Angelis. The Stern diatomic sequence via generalized Chebyshev polynomials. American Mathematical Monthly, 124(5):451–455, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.5.451.

Nathanson:2017:SPR

[1115] Melvyn B. Nathanson. Subalgebras of a polynomial ring that are not finitely generated. American Mathematical Monthly, 124(5):456–458, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.5.456.

Bogosel:2017:GPS

[1116] Beniamin Bogosel. A geometric proof of the Siebeck–Marden theorem. American Mathematical Monthly, 124(5):459–463, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.5.459.

Tandra:2017:MDP

[1117] Haryono Tandra. A more direct proof of the extreme value theorem. American Mathematical Monthly, 124(5):464, May 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.5.464.

Anonymous:2017:PSe

[1118] Anonymous. Problems and solutions. American Mathematical Monthly, 124(5):465–474, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.5.465.

Hersh:2017:BRE

[1119] Reuben Hersh. Book review: Elements of Mathematics: From Euclid to Gödel. American Mathematical Monthly, 124(5):475–478, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.5.475.

Anonymous:2017:EEa

[1120] Anonymous. Editor’s endnotes. American Mathematical Monthly, 124 (5):480, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.5.480.

Anonymous:2017:BMe

[1121] Anonymous. Back matter. American Mathematical Monthly, 124(5):??, May 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.5.bm.

Anonymous:2017:FMf

[1122] Anonymous. Front matter. American Mathematical Monthly, 124 (6):481–482, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.6.fm.

Su:2017:MHF

[1123] Francis Edward Su. Mathematics for human flourishing. American Mathematical Monthly, 124(6):483–493, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.

jstor.org/stable/10.4169/amer.math.monthly.124.6.483.

Kupavskii:2017:TPP

[1124] Andrey Kupavskii and János Pach. From Tarski’s plank problem to simultaneous approximation. American Mathematical Monthly, 124 (6):494–505, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.6.494.

Aldous:2017:INE

[1125] David Aldous and Weijian Han. Introducing Nash equilibria via an online casual game that people actually play. American Mathematical Monthly, 124(6):506–517, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.6.506.

Cairns:2017:ECS

[1126] Grant Cairns. Equitable candy sharing. American Mathematical Monthly, 124(6):518–526, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.6.518.

Holland:2017:LBS

[1127] Finbarr Holland. On the limiting behavior of solutions of linear difference equations. American Mathematical Monthly, 124(6):527–534, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.6.527.

Carter:2017:FWM

[1128] Paul Carter and David Lowry-Duda. On functions whose mean value abscissas are midpoints, with connections to harmonic functions. American Mathematical Monthly, 124(6):535–542, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.6.535.

Glass:2017:KFA

[1129] Darren B. Glass. Klein four actions on graphs and sets. American Mathematical Monthly, 124(6):543–547, June/July 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.6.543.

Patrakeev:2017:DET

[1130] Mikhail A. Patrakeev. Dissection of an equilateral triangle into five congruent open sets. American Mathematical Monthly, 124(6):547, June/ July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.6.547.

Steinlein:2017:FLT

[1131] Heinrich Steinlein. Fermat’s Little Theorem and Gauss congruence: Matrix versions and cyclic permutations. American Mathematical Monthly, 124(6):548–553, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.6.548.

Merca:2017:RZF

[1132] Mircea Merca. The Riemann zeta function with even arguments as sums over integer partitions. American Mathematical Monthly, 124 (6):554–557, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.6.554.

Jones:2017:EDN

[1133] Lee K. Jones. An elementary derivation of the numerical integration bounds in beginning calculus. American Mathematical Monthly, 124 (6):558–561, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.6.558.

Mestrovic:2017:VSP

[1134] Romeo Mestrović. A very short proof of the infinitude of primes. American Mathematical Monthly, 124(6):562, June/July 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.6.562.

Anonymous:2017:PSf

[1135] Anonymous. Problems and solutions. American Mathematical Monthly, 124(6):563–571, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.6.563.

Nunemacher:2017:BRV

[1136] Jeffrey Nunemacher. Book review: Vector Calculus, Linear Algebra, and Differential Forms: A Unified Approach (5th Edition). American Mathematical Monthly, 124(6):572–575, June/July 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.6.572.

Anonymous:2017:BMf

[1137] Anonymous. Back matter. American Mathematical Monthly, 124(6):??, June/July 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.6.bm.

Anonymous:2017:FMg

[1138] Anonymous. Front matter. American Mathematical Monthly, 124(7): 577–578, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.7.fm.

Winkler:2017:SBC

[1139] Peter Winkler. The Sleeping Beauty controversy. American Mathematical Monthly, 124(7):579–587, August/September 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.7.579.

Akopyan:2017:LCP

[1140] Arseniy Akopyan and Vladislav Vysotsky. On the lengths of curves passing through boundary points of a planar convex shape. American Mathematical Monthly, 124(7):588–596, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.7.588.

Apagodu:2017:UFD

[1141] Moa Apagodu and Doron Zeilberger. Using the “freshman’s dream” to prove combinatorial congruences. American Mathematical Monthly, 124(7):597–608, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.7.597.

West:2017:FSE

[1142] Philip P. West and Brian D. Sittinger. A further stroll into the Eisenstein primes. American Mathematical Monthly, 124(7):609–620, August/ September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.7.609.

Bowers:2017:MRE

[1143] John C. Bowers and Philip L. Bowers. A Menger redux: Embedding metric spaces isometrically in Euclidean space. American Mathematical Monthly, 124(7):621–636, August/September 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.7.621.

Levin:2017:CWG

[1144] David A. Levin and Yuval Peres. Counting walks and graph homomorphisms via Markov chains and importance sampling. American Mathematical Monthly, 124(7):637–641, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.7.637.

Anonymous:2017:OMM

[1145] Anonymous. The other Moore method. American Mathematical Monthly, 124(7):641, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.124.7.641.

Axler:2017:FRP

[1146] Christian Axler and Thomas Leßmann. On the first k-Ramanujan prime. American Mathematical Monthly, 124(7):642–646, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.7.642.

Nimbran:2017:IPS

[1147] Amrik Singh Nimbran. An infinite product for the square root of an integer. American Mathematical Monthly, 124(7):647–650, August/ September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.7.647.

Kania:2017:BCU

[1148] Tomasz Kania. On bases that are closed under multiplication. American Mathematical Monthly, 124(7):651–653, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.7.651.

Ciaurri:2017:MGA

[1149] Óscar Ciaurri. Maria Gaetana Agnesi meets Johannes Kepler. American Mathematical Monthly, 124(7):654–658, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.7.654.

Singh:2017:APB

[1150] Jitender Singh. Another proof of the binomial theorem. American Mathematical Monthly, 124(7):658, August/September 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.7.658.

Anonymous:2017:PSg

[1151] Anonymous. Problems and solutions. American Mathematical Monthly, 124(7):659–666, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.7.659.

Jongsma:2017:BRS

[1152] Calvin Jongsma. Book review: Sourcebook in the Mathematics of Medieval Europe and North Africa. American Mathematical Monthly, 124 (7):667–672, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.124.7.667.

Anonymous:2017:BMg

[1153] Anonymous. Back matter. American Mathematical Monthly, 124(7): ??, August/September 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.7.bm.

Anonymous:2017:FMh

[1154] Anonymous. Front matter. American Mathematical Monthly, 124 (8):673–674, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.8.fm.

Klosinski:2017:SSW

[1155] Leonard F. Klosinski, Gerald L. Alexanderson, and Mark Krusemeyer. The Seventy-Seventh William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 124(8):675–687, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.8.675.

Basu:2017:GFP

[1156] Soham Basu and Daniel J. Velleman. On Gauss’s first proof of the fundamental theorem of algebra. American Mathematical Monthly, 124(8):688–694, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.8.688.

Kellner:2017:PSD

[1157] Bernd C. Kellner and Jonathan Sondow. Power-sum denominators. American Mathematical Monthly, 124(8):695–709, October 2017. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.8.695.

Parlier:2017:DDF

[1158] Hugo Parlier and Samuel Zappa. Distances in domino flip graphs. American Mathematical Monthly, 124(8):710–722, October 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.8.710.

Garcia:2017:LIS

[1159] Stephan Ramon Garcia. Linearly independent spanning sets. American Mathematical Monthly, 124(8):722, October 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.8.722.

Abel:2017:RCO

[1160] Ulrich Abel, Linda Beukemann, and Vitaliy Kushnirevych. Rolling curves: An old proof of the roulette lemma. American Mathematical Monthly, 124(8):723–736, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.8.723.

Platt:2017:STS

[1161] David J. Platt and Timothy S. Trudgian. On the sum of two squares and at most two powers of 2. American Mathematical Monthly, 124 (8):737–740, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.8.737.

Marklof:2017:TGT

[1162] Jens Marklof and Andreas Strömbergsson. The three gap theorem and the space of lattices. American Mathematical Monthly, 124(8):741–745, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.8.741.

Kannai:2017:OOP

[1163] Zoltán Kánnai. A one-to-one popcorn function. American Mathematical Monthly, 124(8):746–748, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.8.746.

Jukna:2017:MCS

[1164] Stasys Jukna. Minkowski complexity of sets: An easy lower bound. American Mathematical Monthly, 124(8):749–753, October 2017. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.8.749.

Anonymous:2017:PSh

[1165] Anonymous. Problems and solutions. American Mathematical Monthly, 124(8):754–762, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.8.754.

Anonymous:2017:PMC

[1166] Anonymous. Problems for metagrobologists: A collection of puzzles with real mathematical, logical, or scientific content; one hundred prisoners and a light bulb. American Mathematical Monthly, 124(8):763–768, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.8.763.

Anonymous:2017:BMh

[1167] Anonymous. Back matter. American Mathematical Monthly, 124(8): ??, October 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.8.bm.

Anonymous:2017:FMi

[1168] Anonymous. Front matter. American Mathematical Monthly, 124 (9):769–770, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.9.fm.

Chapman:2017:GE

[1169] Scott T. Chapman and Joseph A. Gallian. From the Guest Editors. American Mathematical Monthly, 124(9):771, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.9.771.

Anonymous:2017:AC

[1170] Anonymous. About the cover. American Mathematical Monthly, 124(9): 772, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.9.772.

Cahen:2017:BEW

[1171] Paul-Jean Cahen, Jean-Luc Chabert, and Kiran S. Kedlaya. Bhargava’s early work: The Genesis of P -orderings. American Mathematical Monthly, 124(9):773–790, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.9.773.

Adams:2017:TKF

[1172] Colin Adams. Turning knots into flowers and related undergraduate research. American Mathematical Monthly, 124(9):791–806, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.9.791.

Silva:2017:MLN

[1173] Cesar E. Silva. On mixing-like notions in infinite measure. American Mathematical Monthly, 124(9):807–825, November 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.9.807.

Chu:2017:DAR

[1174] Adrian Chun Pong Chu. Dihedral angle of the regular n-simplex via Menelaus’ Theorem. American Mathematical Monthly, 124(9):826, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.9.826.

Zhao:2017:ERG

[1175] Yufei Zhao. Extremal regular graphs: Independent sets and graph homomorphisms. American Mathematical Monthly, 124(9):827–843, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.9.827.

Nath:2017:ATC

[1176] Rishi Nath. Advances in the theory of cores and simultaneous core partitions. American Mathematical Monthly, 124(9):844–861, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL https://www.jstor.org/stable/10.4169/amer.math.monthly.
124.9.844.

Kaplan:2017:CNS

[1177] Nathan Kaplan. Counting numerical semigroups. American Mathematical Monthly, 124(9):862–875, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.9.862.

Castillo-Chavez:2017:SDR

[1178] Carlos Castillo-Chavez, Christopher Kribs, and Benjamin Morin. Student-driven research at the Mathematical and Theoretical Biology Institute. American Mathematical Monthly, 124(9):876–892, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.9.876.

Anonymous:2017:BMi

[1179] Anonymous. Back matter. American Mathematical Monthly, 124(9):??, November 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.9.bm.

Anonymous:2017:FM

[1180] Anonymous. Front matter. American Mathematical Monthly, 124 (10):893–894, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.10.fm.

Frankl:2017:DCN

[1181] Peter Frankl, Amram Meir, and János Pach. Decomposition of a cube into nearly equal smaller cubes. American Mathematical Monthly, 124 (10):895–904, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.10.895.

Barnsley:2017:SSP

[1182] Michael Barnsley and Andrew Vince. Self-similar polygonal tiling. American Mathematical Monthly, 124(10):905–921, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https:

//www.jstor.org/stable/10.4169/amer.math.monthly.124.10.905.

Tsukerman:2017:BMT

[1183] Emmanuel Tsukerman and Ellen Veomett. Brunn–Minkowski theory and Cauchy’s surface area formula. American Mathematical Monthly, 124(10):922–929, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.10.922.

Bugeaud:2017:ARR

[1184] Yann Bugeaud, Andrej Dujella, Tomislav Pejković, and Bruno Salvy. Absolute real root separation. American Mathematical Monthly, 124 (10):930–936, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.10.930.

Kreh:2017:LV

[1185] Martin Kreh. “Lights out” and variants. American Mathematical Monthly, 124(10):937–950, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.10.937.

Granville:2017:SAP

[1186] Andrew Granville. Squares in arithmetic progressions and infinitely many primes. American Mathematical Monthly, 124(10):951–954, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.10.951.

Cooper:2017:TBF

[1187] Joshua Cooper and Anton Swifton. Throwing a ball as far as possible, revisited. American Mathematical Monthly, 124(10):955–959, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.math.
monthly.124.10.955.

Chebolu:2017:HMU

[1188] Sunil K. Chebolu and Keir Lockridge. How many units can a commutative ring have? American Mathematical Monthly, 124(10):960–965, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.10.960.

Szigeti:2017:CEP

[1189] Jenő Szigeti and Leon vanWyk. A constructive elementary proof of the Skolem–Noether theorem for matrix algebras. American Mathematical Monthly, 124(10):966–968, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.10.966.

Cerruti:2017:IPF

[1190] Umberto Cerruti and Nadir Murru. If the primes are finite, then all of them divide the number one. American Mathematical Monthly, 124(10):969, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.10.969.

Anonymous:2017:PS

[1191] Anonymous. Problems and solutions. American Mathematical Monthly, 124(10):970–978, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.10.970.

Anonymous:2017:BRM

[1192] Anonymous. Book review: From Music to Mathematics: Exploring the Connections, Gareth E. Roberts. Johns Hopkins University Press, Baltimore, MD, 2016. Price: $49.95 ISBN: 978-1-4214-1918-3. American Mathematical Monthly, 124(10):979–982, December 2017. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://

www.jstor.org/stable/10.4169/amer.math.monthly.124.10.979.

Anonymous:2017:EEb

[1193] Anonymous. Editor’s endnotes. American Mathematical Monthly, 124 (10):983–985, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/stable/

10.4169/amer.math.monthly.124.10.983.

Anonymous:2017:MR

[1194] Anonymous. Monthly referees for 2017. American Mathematical Monthly, 124(10):986–988, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/10.4169/amer.math.monthly.124.10.986.

Anonymous:2017:BM

[1195] Anonymous. Back matter. American Mathematical Monthly, 124(10):??, December 2017. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.124.10.bm.

Rice:2018:PPP

[1196] Adrian Rice. Partnership, partition, and proof: The path to the Hardy– Ramanujan partition formula. American Mathematical Monthly, 125(1): 3–15, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Boas:2018:CRS

[1197] Harold P. Boas. Cauchy’s residue sore thumb. American Mathematical Monthly, 125(1):16–28, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:2018:YAM

[1198] Anonymous. 100 years ago this month in The American Mathematical Monthly, edited by Vadim Ponomarenko. American Mathematical Monthly, 125(1):28, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mazur:2018:AVP

[1199] Kristen Mazur, Mutiara Sondjaja, Matthew Wright, and Carolyn Yarnall. Approval voting in product societies. American Mathematical Monthly, 125(1):29–43, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Treeby:2018:FTP

[1200] David Treeby. Further thoughts on a paradoxical tower. American Mathematical Monthly, 125(1):44–60, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Meyer:2018:RMU

[1201] Carl D. Meyer. Rank my update, please. American Mathematical Monthly, 125(1):61–64, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shukla:2018:SPC

[1202] Alok Shukla. A short proof of Cayley’s tree formula. American Mathematical Monthly, 125(1):65–68, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Clifford:2018:PNR

[1203] John Clifford, Kelly Jabbusch, and Michael Lachance. Preimages of the numerical range. American Mathematical Monthly, 125(1):69–74, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Marco:2018:AEC

[1204] Giuseppe De Marco. An almost everywhere constant function surjective on every interval. American Mathematical Monthly, 125(1):75–76, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Titkos:2018:SPF

[1205] Tamás Titkos. The singular part as fixed point. American Mathematical Monthly, 125(1):77–80, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Singh:2018:NPM

[1206] Jitender Singh. A noninductive proof of de Moivre’s formula. American Mathematical Monthly, 125(1):80, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSa

[1207] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(1):81–89, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dietz:2018:MLC

[1208] Geoffrey Dietz. The mathematics Lover’s companion: Masterpieces for everyone. American Mathematical Monthly, 125(1):90–95, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berndt:2018:CPG

[1209] Bruce C. Berndt, Sun Kim, and Alexandru Zaharescu. The circle problem of Gauss and the divisor problem of Dirichlet — still unsolved. American Mathematical Monthly, 125(2):99–114, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maltenfort:2018:PF

[1210] Michael Maltenfort. Pascal functions. American Mathematical Monthly, 125(2):115–129, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dawson:2018:NER

[1211] C. Bryan Dawson. A new extension of the Riemann integral. American Mathematical Monthly, 125(2):130–140, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kulczycki:2018:ESS

[1212] Marcin Kulczycki, Dominik Kwietniak, and Jian Li. Entropy of subordinate shift spaces. American Mathematical Monthly, 125(2):141–148, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brooks:2018:RCN

[1213] Cara D. Brooks and Alberto A. Condori. A resolvent criterion for normality. American Mathematical Monthly, 125(2):149–156, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sondow:2018:LSG

[1214] Jonathan Sondow. Lorentz Stöer’s Geometria et Perspectiva. American Mathematical Monthly, 125(2):157, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Melman:2018:OPM

[1215] Aaron Melman. Optimality of a polynomial multiplier. American Mathematical Monthly, 125(2):158–163, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pinchasi:2018:TUS

[1216] Rom Pinchasi. A theorem on unit segments on the real line. American Mathematical Monthly, 125(2):164–168, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Howard:2018:FEC

[1217] Ralph Howard, Virginia Johnson, and George F. McNulty. A functional equation characterization of Archimedean ordered fields. American Mathematical Monthly, 125(2):169–172, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stuart:2018:IIS

[1218] Christopher Stuart. An inequality involving sin(n). American Mathematical Monthly, 125(2):173–174, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chu:2018:PFR

[1219] Wenchang Chu. Partial fractions and Riesz’ interpolation formula for trigonometric polynomials. American Mathematical Monthly, 125(2): 175–178, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:2018:PS

[1220] Anonymous. Problems and solutions. American Mathematical Monthly, 125(2):179–187, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cowen:2018:BGM

[1221] Robert Cowen. Review: A Beginner’s Guide to Mathematical Logic, By Raymond M. Smullyan. Dover, Mineola, NY, 2014. vi + 284 pp. store.doverpublications.com/0486492370.html ISBN: 978-0-486-49237-7, $19.95. A Beginner’s Further Guide to Mathematical Logic. By Raymond M. Smullyan. World Scientific, Hackensack, NJ, 2017. x + 278 pp. worldscientific.com/worldscibooks/10.1142/9865 ISBN: 978-981-4725-72-9, $24.00 (paperback). American Mathematical Monthly, 125(2): 188–192, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Crannell:2018:YGG

[1222] Annalisa Crannell. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2018 to David Bressoud for distinguished service to mathematics. American Mathematical Monthly, 125(3):195–197, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Vandervelde:2018:WNR

[1223] Sam Vandervelde. The Worpitzky numbers revisited. American Mathematical Monthly, 125(3):198–206, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gordon:2018:ICP

[1224] Russell A. Gordon and José Carlos Santos. An interesting construction problem. American Mathematical Monthly, 125(3):207–221, 2018. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Plaza:2018:HSD

[1225] Ángel Plaza. The harmonic series diverges. American Mathematical Monthly, 125(3):222, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Strang:2018:MFM

[1226] Gilbert Strang. Multiplying and factoring matrices. American Mathematical Monthly, 125(3):223–230, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Salwinski:2018:CBC

[1227] David Salwinski. The continuous binomial coefficient: an elementary approach. American Mathematical Monthly, 125(3):231–244, 2018. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gilula:2018:CSR

[1228] Maxim Gilula. A class of simple rearrangements of the alternating harmonic series. American Mathematical Monthly, 125(3):245–256, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Woods:2018:AGP

[1229] Kevin Woods. AdaM and GrahaM play the stock market. American Mathematical Monthly, 125(3):257, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goldmakher:2018:RLF

[1230] Leo Goldmakher and Paul Pollack. Refinements of Lagrange’s Four-Square Theorem. American Mathematical Monthly, 125(3):258–263, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shiu:2018:FTG

[1231] Peter Shiu. A footnote to the Three Gaps Theorem. American Mathematical Monthly, 125(3):264–266, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leahy:2018:MAS

[1232] Andrew Leahy. The method of Archimedes in the Seventeenth Century. American Mathematical Monthly, 125(3):267–272, 2018. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mazur:2018:IVT

[1233] Marcin Mazur. An inequality for the volume of a tetrahedron. American Mathematical Monthly, 125(3):273–275, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSb

[1234] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(3):276–283, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Glass:2018:JS

[1235] Darren Glass. The joy of SET. American Mathematical Monthly, 125(3): 284–288, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Broaddus:2018:SCS

[1236] David Broaddus, Stephen Lovett, Dawson Miller, and Caitlin Smith. Seeing curvature on specular surfaces. American Mathematical Monthly, 125 (4):291–305, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Carter:2018:MCI

[1237] Larry Carter and Stan Wagon. The Mensa Correctional Institute. American Mathematical Monthly, 125(4):306–319, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bras-Amoros:2018:DAR

[1238] Maria Bras-Amorós. A decoding approach to Reed–Solomon codes from their definition. American Mathematical Monthly, 125(4):320–338, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ho:2018:CDU

[1239] Chungwu Ho and Seth Zimmerman. On certain dense, uncountable subsets of the real line. American Mathematical Monthly, 125(4):339–346, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leher:2018:AS

[1240] Eli Leher. is not 1.41421356237 or anything of the sort. American Mathematical Monthly, 125(4):346, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Keller:2018:IDH

[1241] Matthias Keller, Yehuda Pinchover, and Felix Pogorzelski. An improved discrete Hardy inequality. American Mathematical Monthly, 125(4):347– 350, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lubeck:2018:SHS

[1242] Brian Lubeck and Vadim Ponomarenko. Subsums of the harmonic series. American Mathematical Monthly, 125(4):351–355, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pandey:2018:DDN

[1243] Vaibhav Pandey, Sagar Shrivastava, and Balasubramanian Sury. A Dedekind domain with nontrivial class group. American Mathematical Monthly, 125(4):356–359, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Villarino:2018:EAS

[1244] Mark B. Villarino. The error in an alternating series. American Mathematical Monthly, 125(4):360–364, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Adell:2018:BIM

[1245] José A. Adell and Alberto Lekuona. Binomial identities and moments of random variables. American Mathematical Monthly, 125(4):365–369, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSc

[1246] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(4):370–378, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taalman:2018:RVM

[1247] Laura Taalman. Review: Visualizing Mathematics With 3D Printing by Henry Segerman and Pasta by Design by George Legendre. American Mathematical Monthly, 125(4):379–384, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Buhler:2018:MND

[1248] Joe Buhler, Ron Graham, and Al Hales. Maximally nontransitive dice. American Mathematical Monthly, 125(5):387–399, 2018. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:2018:GFM

[1249] Jonathan M. Borwein and Robert M. Corless. Gamma and factorial in the Monthly. American Mathematical Monthly, 125(5):400–424, May 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lemos:2018:SCP

[1250] Abílio Lemos and Pedro H. A. de Oliveira. Suggested corrections for “A Principal Ideal Domain That Is Not a Euclidean Domain”. American Mathematical Monthly, 125(5):425, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

delaFuente:2018:VML

[1251] Daniel de la Fuente, Rafael Ramírez-Uclés, and Juan F. Ruiz-Hidalgo. A visual model of the Lorentz triangular moduli space. American Mathematical Monthly, 125(5):426–432, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

daSilva:2018:OTG

[1252] João Lita da Silva. On one type of generalized Vandermonde determinants. American Mathematical Monthly, 125(5):433–442, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lynd:2018:SGP

[1253] Chris D. Lynd and James Wright Sharpe. Sequences generated by powers of the k-th-order Fibonacci recurrence relation. American Mathematical Monthly, 125(5):443–446, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Munagi:2018:SPI

[1254] Augustine O. Munagi. Set partitions with isolated singletons. American Mathematical Monthly, 125(5):447–452, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Smirnov:2018:GTT

[1255] A. G. Smirnov. A Gronwall-type trigonometric inequality. American Mathematical Monthly, 125(5):453–456, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Azouzi:2018:CPS

[1256] Youssef Azouzi. Cluster points of slow sequences. American Mathematical Monthly, 125(5):457–461, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Curtin:2018:ASP

[1257] Eugene Curtin. Another short proof of the Hairy Ball Theorem. American Mathematical Monthly, 125(5):462–463, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rodriguez-Quintana:2018:BCF

[1258] Iago Rodríguez-Quintana. A beautiful continued fraction for π. American Mathematical Monthly, 125(5):463, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jakhar:2018:ICP

[1259] Anuj Jakhar. An irreducibility criterion for polynomials with integer coefficients. American Mathematical Monthly, 125(5):464–465, 2018. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSd

[1260] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(5):466–475, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ash:2018:PNR

[1261] Avner Ash. Review: Prime Numbers and the Riemann Hypothesis by Barry Mazur and William Stein. American Mathematical Monthly, 125 (5):476–480, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Granville:2018:UDS

[1262] Andrew Granville. Using dynamical systems to construct infinitely many primes. American Mathematical Monthly, 125(6):483–496, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chai:2018:RPA

[1263] Ying Chai, Liping Yuan, and Tudor Zamfirescu. Rupert property of Archimedean solids. American Mathematical Monthly, 125(6):497–504, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Huber:2018:CR

[1264] Greg Huber, Kay Pechenick Shultz, and John E. Wetzel. The n-cube is Rupert. American Mathematical Monthly, 125(6):505–512, 2018. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Villarino:2018:HPH

[1265] Mark B. Villarino, William Gasarch, and Kenneth W. Regan. Hilbert’s proof of his irreducibility theorem. American Mathematical Monthly, 125 (6):513–530, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

deCastro:2018:CBC

[1266] Philip de Castro, Desiree Domini, Tom Edgar, Devon M. Johnson, Steven Klee, and Ranjani Sundaresan. Counting binomial coefficients divisible by a prime power. American Mathematical Monthly, 125(6):531–540, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McGrath:2018:NRH

[1267] Peter McGrath. A note regarding Hopf’s Umlaufsatz. American Mathematical Monthly, 125(6):541–544, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chin:2018:FRO

[1268] Angelina Y. M. Chin. Finite rings of odd order with few nilpotent and idempotent elements. American Mathematical Monthly, 125(6):545–548, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Asgarli:2018:NPW

[1269] Shamil Asgarli. A new proof of Warning’s Second Theorem. American Mathematical Monthly, 125(6):549–553, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Singh:2018:SPL

[1270] Jitender Singh. A short proof that Lebesgue outer measure of an interval is its length. American Mathematical Monthly, 125(6):553, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Coudene:2018:SIC

[1271] Yves Coudène. A strange inequality concerning alternating series. American Mathematical Monthly, 125(6):554–557, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pause:2018:BPS

[1272] Kapil R. Shenvi Pause. Basel Problem: a solution motivated by the power of a point. American Mathematical Monthly, 125(6):558–560, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Moreno:2018:CBD

[1273] Samuel G. Moreno and Esther M. García-Caballero. A characterization of Bernoulli differential equations. American Mathematical Monthly, 125 (6):561, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSe

[1274] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(6):562–570, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rosenhouse:2018:R

[1275] Jason Rosenhouse. Reviews. American Mathematical Monthly, 125(6): 571–576, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morrison:2018:SDF

[1276] Ian Morrison. Sacks of dice with fair totals. American Mathematical Monthly, 125(7):579–592, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goins:2018:LPV

[1277] Edray H. Goins, Pamela E. Harris, Bethany Kubik, and Aba Mbirika. Lattice point visibility on generalized lines of sight. American Mathematical Monthly, 125(7):593–601, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gard:2018:WGC

[1278] Andrew Gard. The wild goose chase problem. American Mathematical Monthly, 125(7):602–611, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Al-Afifi:2018:ADG

[1279] Ghada Al-Afifi, Mowaffaq Hajja, and Ahmad Hamdan. Another ndimensional generalization of Pompeiu’s Theorem. American Mathematical Monthly, 125(7):612–622, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nathanson:2018:EPK

[1280] Melvyn B. Nathanson. An elementary proof for the Krull dimension of a polynomial ring. American Mathematical Monthly, 125(7):623–637, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berele:2018:NSP

[1281] Allan Berele and Stefan Catoiu. Nonuniqueness of sixpartite points. American Mathematical Monthly, 125(7):638–642, 2018. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chapman:2018:WHW

[1282] Scott T. Chapman. What happens when the division algorithm “almost” works. American Mathematical Monthly, 125(7):643–647, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stoica:2018:WMF

[1283] George Stoica. When must a flat function be identically 0? American Mathematical Monthly, 125(7):648–649, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gencev:2018:PIB

[1284] Marian Gencev. On a proof of the inequality between the arithmetic and geometric means. American Mathematical Monthly, 125(7):650–652, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Reinholz:2018:PFL

[1285] Daniel L. Reinholz. Peer feedback for learning mathematics. American Mathematical Monthly, 125(7):653–658, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Trevino:2018:SPS

[1286] Enrique Treviño. A short proof of a sum of powers formula. American Mathematical Monthly, 125(7):659, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSf

[1287] Gerald A. Edgar, Daniel Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(7):660–668, 2018. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Meadows:2018:R

[1288] Alex Meadows. Reviews. American Mathematical Monthly, 125(7):669– 672, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Krusemeyer:2018:SEW

[1289] Mark Krusemeyer and Daniel H. Ullman. The Seventy-Eighth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 125(8):675–688, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chan:2018:RRT

[1290] Stephanie Chan. Rational right triangles of a given area. American Mathematical Monthly, 125(8):689–703, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pearson:2018:VFK

[1291] John Pearson. Viète’s formula, Knar’s formula, and the geometry of the gamma function. American Mathematical Monthly, 125(8):704–714, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Talvila:2018:YKY

[1292] Erik Talvila. You know you’re a mathematician if . . .. American Mathematical Monthly, 125(8):714, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Morales:2018:WPL

[1293] Alejandro H. Morales, Igor Pak, and Greta Panova. Why is pi less than twice phi? American Mathematical Monthly, 125(8):715–723, 2018. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Veerman:2018:SBB

[1294] J. J. P. Veerman. Social balance and the Bernoulli equation. American Mathematical Monthly, 125(8):724–732, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shioji:2018:SPU

[1295] Naoki Shioji. Simple proofs of the uniform convexity of L^{p} and the Riesz representation theorem for L^{p}. American Mathematical Monthly, 125(8): 733–738, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ciesielski:2018:MC

[1296] Krzysztof Chris Ciesielski. Monsters in calculus. American Mathematical Monthly, 125(8):739–744, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thakkar:2018:IPU

[1297] Aalok Thakkar. Infinitude of primes using formal languages. American Mathematical Monthly, 125(8):745–749, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Patterson:2018:SCS

[1298] T. N. L. Patterson. Sines, cosines, square roots, and binary bits. American Mathematical Monthly, 125(8):750–754, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Stander:2018:SSW

[1299] Mark Stander. The smallest solution to AB = C, where A, B, and C are square matrices composed of distinct primes. American Mathematical Monthly, 125(8):754, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSg

[1300] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(8):755–763, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Merris:2018:R

[1301] Russell Merris. Reviews. American Mathematical Monthly, 125(8):764– 768, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brams:2018:CRM

[1302] Steven J. Brams, Mehmet S. Ismail, D. Marc Kilgour, and Walter Stromquist. Catch-up: a rule that makes service sports more competitive. American Mathematical Monthly, 125(9):771–796, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Williams:2018:EYW

[1303] Kenneth S. Williams. Everything you wanted to know about ax^{2} +by^{2} + cz^{2}+dt^{2} but were afraid to ask. American Mathematical Monthly, 125(9): 797–810, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andreescu:2018:QFC

[1304] Titu Andreescu and Oleg Mushkarov. Quadratic forms, Chebyshev polynomials, and geometric inequalities. American Mathematical Monthly, 125(9):811–819, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tointon:2018:MCT

[1305] Matthew C. H. Tointon. The Mahler conjecture in two dimensions via the probabilistic method. American Mathematical Monthly, 125(9):820–828, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sallows:2018:SDC

[1306] Lee Sallows. A self-descriptive crossword puzzle. American Mathematical Monthly, 125(9):829, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Baryshnikov:2018:STD

[1307] Yuliy Baryshnikov and Robert Ghrist. Stokes’s Theorem, data, and the polar ice caps. American Mathematical Monthly, 125(9):830–834, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Harutyunyan:2018:WCI

[1308] Davit Harutyunyan. When the Cauchy inequality becomes a formula. American Mathematical Monthly, 125(9):835–838, 2018. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ribeiro:2018:APF

[1309] Pedro Ribeiro. Another proof of the famous formula for the zeta function at positive even integers. American Mathematical Monthly, 125(9): 839–841, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leonetti:2018:CCF

[1310] Paolo Leonetti. A characterization of convex functions. American Mathematical Monthly, 125(9):842–844, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Plante:2018:CEA

[1311] Joseph Plante. Cubic equations from an analytic point of view. American Mathematical Monthly, 125(9):845–849, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Park:2018:DPH

[1312] Bryan Park. On the divergence of the prime harmonic series. American Mathematical Monthly, 125(9):850, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSh

[1313] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(9):851–859, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cohen:2018:R

[1314] Marion Cohen. Reviews. American Mathematical Monthly, 125(9):860– 864, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sallows:2018:SSD

[1315] Lee Sallows. Solution to the self-descriptive crossword puzzle. American Mathematical Monthly, 125(9):864, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Giansiracusa:2018:GC

[1316] Noah Giansiracusa and Cameron Ricciardi. Geometry in the courtroom. American Mathematical Monthly, 125(10):867–877, 2018. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alexeev:2018:ITG

[1317] Boris Alexeev and Dustin G. Mixon. An impossibility theorem for gerrymandering. American Mathematical Monthly, 125(10):878–884, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ding:2018:LLC

[1318] Pisheng Ding. On the lengths of level curves. American Mathematical Monthly, 125(10):885–900, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Clausing:2018:DM

[1319] Achim Clausing. Ducci matrices. American Mathematical Monthly, 125 (10):901–921, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zheng:2018:GMP

[1320] Chen Zheng and Zhijun Zhou. A generalization of a Monthly problem. American Mathematical Monthly, 125(10):922, 2018. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hocutt:2018:EDN

[1321] J. A. Hocutt and P. L. Robinson. Everywhere differentiable, nowhere continuous functions. American Mathematical Monthly, 125(10):923–928, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Merca:2018:PIR

[1322] Mircea Merca and Maxie D. Schmidt. A partition identity related to Stanley’s Theorem. American Mathematical Monthly, 125(10):929–933, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bary-Soroker:2018:RPD

[1323] Lior Bary-Soroker and Ofir Gorodetsky. Roots of polynomials and the derangement problem. American Mathematical Monthly, 125(10):934–938, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Torne:2018:LTM

[1324] José Antonio Torné and Juan Luis Varona. The Lindemann Theorem for matrices. American Mathematical Monthly, 125(10):939–940, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Liu:2018:SSM

[1325] Peng Liu and Shibo Liu. On the surjectivity of smooth maps into Euclidean spaces and the fundamental theorem of algebra. American Mathematical Monthly, 125(10):941–943, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2018:PSi

[1326] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 125(10):944–952, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Guralnick:2018:R

[1327] Robert M. Guralnick and Gary M. Seitz. Reviews. American Mathematical Monthly, 125(10):953–955, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Editor:2018:EE

[1328] Susan Jane Colley Editor. Editor’s endnotes. American Mathematical Monthly, 125(10):956–957, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:2018:MR

[1329] Anonymous. Monthly referees for 2018. American Mathematical Monthly, 125(10):958–960, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Colley:2019:LE

[1330] Susan Jane Colley. Letter from the Editor. American Mathematical Monthly, 126(1):3, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Athreya:2019:CSA

[1331] Jayadev S. Athreya, Bruce Reznick, and Jeremy T. Tyson. Cantor set arithmetic. American Mathematical Monthly, 126(1):4–17, 2019. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Frick:2019:ARH

[1332] Florian Frick, Kelsey Houston-Edwards, and Frédéric Meunier. Achieving rental harmony with a secretive roommate. American Mathematical Monthly, 126(1):18–32, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sanchez-Reyes:2019:JMR

[1333] Javier Sánchez-Reyes. The Joukowsky map reveals the cubic equation. American Mathematical Monthly, 126(1):33–40, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bossaller:2019:AIM

[1334] Daniel P. Bossaller and Sergio R. López-Permouth. On the associativity of infinite matrix multiplication. American Mathematical Monthly, 126 (1):41–52, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pompe:2019:CLN

[1335] Waldemar Pompe. Concurrent lines in the Napoleon-style configurations. American Mathematical Monthly, 126(1):53–59, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sambale:2019:PSN

[1336] Benjamin Sambale. Pseudo Sylow numbers. American Mathematical Monthly, 126(1):60–65, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cocke:2019:NEM

[1337] William Cocke and Geetha Venkataraman. On the number of elements of maximal order in a group. American Mathematical Monthly, 126(1): 66–69, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fridman:2019:PRC

[1338] Dylan Fridman, Juli Garbulsky, Bruno Glecer, James Grime, and Massi Tron Florentin. A prime-representing constant. American Mathematical Monthly, 126(1):70–73, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Caballero:2019:CHP

[1339] José Manuel Rodríguez Caballero. A characterization of the hypotenuses of primitive Pythagorean triangles using partitions into consecutive parts. American Mathematical Monthly, 126(1):74–77, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ponomarenko:2019:HJ

[1340] Vadim Ponomarenko. Help with the jumble. American Mathematical Monthly, 126(1):77, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Donnelly:2019:CF

[1341] John Donnelly. On a conjecture of Fickett. American Mathematical Monthly, 126(1):78–80, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kuhapatanakul:2019:GFL

[1342] K. Kuhapatanakul and K. Thongsing. Generalizations of the Fibonacci– Lucas relations. American Mathematical Monthly, 126(1):81, 2019. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSa

[1343] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 126(1):82–90, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

by:2019:R

[1344] Tullia Dymarz Reviewed by. Reviews. American Mathematical Monthly, 126(1):91–96, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gorkin:2019:FTT

[1345] Pamela Gorkin. Four theorems with their foci on ellipses. American Mathematical Monthly, 126(2):99–111, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sadhukhan:2019:SCP

[1346] Arpan Sadhukhan. A simple construction to prove Mycielski’s theorem. American Mathematical Monthly, 126(2):112, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cantarella:2019:RTP

[1347] Jason Cantarella, Tom Needham, Clayton Shonkwiler, and Gavin Stewart. Random triangles and polygons in the plane. American Mathematical Monthly, 126(2):113–134, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hajja:2019:NQA

[1348] Mowaffaq Hajja and Jonathan Sondow. Newton quadrilaterals, the associated cubic equations, and their rational solutions. American Mathematical Monthly, 126(2):135–150, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jeronimo-Castro:2019:CEB

[1349] J. Jerónimo-Castro and C. Yee-Romero. Characterizations of the Euclidean ball by intersections with planes and slabs. American Mathematical Monthly, 126(2):151–157, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Arnold:2019:MPT

[1350] Maxim Arnold and Anatoly Eydelzon. On matrix Pythagorean triples. American Mathematical Monthly, 126(2):158–160, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Athreya:2019:TVI

[1351] Jayadev S. Athreya and David Aulicino. A trajectory from a vertex to itself on the dodecahedron. American Mathematical Monthly, 126(2): 161–162, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Leonetti:2019:CRS

[1352] Paolo Leonetti. Convergence rates of subseries. American Mathematical Monthly, 126(2):163–167, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sury:2019:RTA

[1353] B. Sury. A ring-theoretic approach to bound the totient function. American Mathematical Monthly, 126(2):167, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Nguyen:2019:NCI

[1354] Nicholas Phat Nguyen. A note on cyclotomic integers. American Mathematical Monthly, 126(2):168–172, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andrica:2019:SNS

[1355] Dorin Andrica and Vlad Crisan. The smallest nontrivial solution to x_{k} ≡ 1(modn) and related sequences. American Mathematical Monthly, 126 (2):173–178, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Donnelly:2019:GAT

[1356] Robert G. Donnelly and Molly W. Dunkum. The group action tri-cycle theorem. American Mathematical Monthly, 126(2):179, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSb

[1357] Edited by Gerald A. Edgar, Daniel H. Ullman, Douglas B. West, and with the collaboration of Paul Bracken. Problems and solutions. American Mathematical Monthly, 126(2):180–188, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berndt:2019:R

[1358] Reviewed by Bruce Berndt. Reviews. American Mathematical Monthly, 126(2):189–192, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Peeples:2019:YGG

[1359] Joanne Peeples, James Álvarez, Ruth Charney, John Harris, Jim Lewis, and Nancy Neudauer. Yueh-Gin Gung and Dr. Charles Y. Hu Award for 2019 to Philip Uri Treisman for distinguished service to mathematics. American Mathematical Monthly, 126(3):195–198, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ullman:2019:DB

[1360] Daniel H. Ullman and Daniel J. Velleman. Differences of bijections. American Mathematical Monthly, 126(3):199–216, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Allouche:2019:TBI

[1361] J.-P. Allouche. Two binomial identities of ruehr revisited. American Mathematical Monthly, 126(3):217–225, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Soundararajan:2019:EZP

[1362] K. Soundararajan. Equidistribution of zeros of polynomials. American Mathematical Monthly, 126(3):226–236, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Perdomo:2019:RTM

[1363] Oscar M. Perdomo. The round Taylor method. American Mathematical Monthly, 126(3):237–251, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ramesh:2019:NGT

[1364] V. P. Ramesh, R. Thangadurai, and R. Thatchaayini. A note on Gauss’s Theorem on primitive roots. American Mathematical Monthly, 126(3): 252–254, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Yap:2019:TC

[1365] Shirley Llamado Yap. A Theodorus construction. American Mathematical Monthly, 126(3):255–259, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mawhin:2019:SPH

[1366] Jean Mawhin. Simple proofs of the Hadamard and Poincaré–Miranda theorems using the Brouwer Fixed Point Theorem. American Mathematical Monthly, 126(3):260–263, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Falkner:2019:HPH

[1367] Neil Falkner. Hahn’s proof of the Hahn decomposition theorem, and related matters. American Mathematical Monthly, 126(3):264–268, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Forgacs:2019:NIS

[1368] Tamás Forgács, Jack Luong, and Joshua Williamson. A note on infinite series with recursively defined terms. American Mathematical Monthly, 126(3):269–274, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rounds:2019:ELS

[1369] Bill Rounds. Euclid’s Lemma and the square root of 2. American Mathematical Monthly, 126(3):274, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gordon:2019:CSH

[1370] Russell A. Gordon. Comments on “Subsums of the Harmonic Series”. American Mathematical Monthly, 126(3):275–279, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSc

[1371] Edited by Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 126(3):280– 288, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Seacrest:2019:PSP

[1372] Deborah E. Seacrest and Tyler Seacrest. The prisoner shouting puzzle and variations. American Mathematical Monthly, 126(4):291–305, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taylor:2019:PAE

[1373] Alan D. Taylor and Stan Wagon. A paradox arising from the elimination of a paradox. American Mathematical Monthly, 126(4):306–318, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Holland:2019:CSI

[1374] Finbarr Holland and Roger Smyth. Cardioids and self-inversive cubic polynomials. American Mathematical Monthly, 126(4):319–329, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chapman:2019:WCN

[1375] Scott T. Chapman. So what is class number 2? American Mathematical Monthly, 126(4):330–339, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mahmoudi:2019:GIS

[1376] M. G. Mahmoudi. A geometric interpretation of the simplicity of SO(3). American Mathematical Monthly, 126(4):340–350, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Demers:2019:SNC

[1377] Simon Demers. The smallest number that cannot be collapsed using fewer than 6 plus signs is 3003344034004. American Mathematical Monthly, 126 (4):351, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ozawa:2019:TMP

[1378] Makoto Ozawa and Scott A. Taylor. Two more proofs that the Kinoshita Graph is knotted. American Mathematical Monthly, 126(4):352– 357, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Korus:2019:NPM

[1379] Péter Kórus. Notes on projectile motion. American Mathematical Monthly, 126(4):358–360, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Glasser:2019:NBR

[1380] M. L. Glasser. A note on Beukers’s and related double integrals. American Mathematical Monthly, 126(4):361–363, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jukna:2019:CFC

[1381] Stasys Jukna. Coin flipping cannot shorten arithmetic computations. American Mathematical Monthly, 126(4):364–366, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berendonk:2019:NTT

[1382] Stephan Berendonk. A Napoleonic Theorem for trapezoids. American Mathematical Monthly, 126(4):367–369, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSd

[1383] Edited by Gerald A. Edgar, Daniel H. Ullman, Douglas B. West, and with the collaboration of Paul Bracken. Problems and solutions. American Mathematical Monthly, 126(4):370–378, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Beineke:2019:R

[1384] Lowell Beineke. Reviews. American Mathematical Monthly, 126(4):379– 384, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Torrence:2019:PBF

[1385] Bruce Torrence. Passing the buck and firing Fibonacci: Adventures with the stochastic abacus. American Mathematical Monthly, 126(5):387–399, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fomin:2019:MIU

[1386] Dmitri V. Fomin. Is the multiset of n integers uniquely determined by the multiset of its s-sums? American Mathematical Monthly, 126(5): 400–417, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Antonelli:2019:SLB

[1387] Tim Antonelli. A surprising link between integer partitions and Euler’s number e. American Mathematical Monthly, 126(5):418–429, 2019. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Munch:2019:CTE

[1388] Niels Juul Munch. A Chebyshev theorem for ellipses in the complex plane. American Mathematical Monthly, 126(5):430–436, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:2019:HJ

[1389] Anonymous. The hardest jumbles. American Mathematical Monthly, 126 (5):436, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cameron:2019:DSV

[1390] Thomas R. Cameron. The determinant from signed volume to the Laplace expansion. American Mathematical Monthly, 126(5):437–447, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Schneegans:2019:HC

[1391] Olivier Schneegans. How close to 10 is π^{2} ? American Mathematical Monthly, 126(5):448, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Herzog:2019:KTF

[1392] Gerd Herzog and Peer Chr. Kunstmann. Korovkin’s theorem for functionals and limits for box integrals. American Mathematical Monthly, 126 (5):449–454, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ponomarenko:2019:YAM

[1393] Edited by Vadim Ponomarenko. 100 years ago this month in The American Mathematical Monthly. American Mathematical Monthly, 126(5): 454, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Andruszkiewicz:2019:CCS

[1394] R. R. Andruszkiewicz and K. Pryszczepko. A curious characterization of subrings of the field Q and applications. American Mathematical Monthly, 126(5):455–458, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mortini:2019:SPC

[1395] Raymond Mortini. A simpler proof of Cohen’s Factorization Theorem. American Mathematical Monthly, 126(5):459–463, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Blaszczyk:2019:SPS

[1396] Aleksander Błaszczyk. A simple proof of Sierpiński’s Theorem. American Mathematical Monthly, 126(5):464–466, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Plaza:2019:TPP

[1397] Angel Plaza. k-th power of a partial sum. American Mathematical Monthly, 126(5):467, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSe

[1398] Edited by Gerald A. Edgar, Daniel H. Ullman, Douglas B. West, and with the collaboration of Paul Bracken. Problems and solutions. American Mathematical Monthly, 126(5):468–476, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ratliff:2019:R

[1399] Reviewed by Thomas Ratliff. Reviews. American Mathematical Monthly, 126(5):477–479, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:2019:SHJ

[1400] Anonymous. Solution to the hardest jumbles. American Mathematical Monthly, 126(5):479, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Colley:2019:EEa

[1401] Susan Jane Colley. Editor’s endnotes. American Mathematical Monthly, 126(5):480, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Adams:2019:KRK

[1402] Colin Adams, Allison Henrich, Kate Kearney, and Nicholas Scoville. Knots related by knotoids. American Mathematical Monthly, 126(6):

483–490, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Garcia:2019:EBE

[1403] Ronaldo Garcia. Elliptic billiards and ellipses associated to the 3-periodic orbits. American Mathematical Monthly, 126(6):491–504, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kolba:2019:NIS

[1404] Tiffany Kolba, Anthony Coniglio, Sarah Sparks, and Daniel Weithers. Noise-induced stabilization of perturbed Hamiltonian systems. American Mathematical Monthly, 126(6):505–518, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Santiago:2019:GTA

[1405] José Hernández Santiago. The group-theoretic analog of Kuratowski’s closure-complement theorem. American Mathematical Monthly, 126(6): 519–526, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Brooks:2019:IFS

[1406] Jennifer Brooks. An interesting family of symmetric polynomials. American Mathematical Monthly, 126(6):527–540, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Holdener:2019:PDP

[1407] Judy Holdener and Emily Rachfal. Perfect and deficient perfect numbers. American Mathematical Monthly, 126(6):541–546, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Treeby:2019:R

[1408] David Treeby. Relativity. American Mathematical Monthly, 126(6):546, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ciesielski:2019:BCR

[1409] K. C. Ciesielski, M. E. Martínez-Gómez, and J. B. Seoane-Sepúlveda. “Big” continuous restrictions of arbitrary functions. American Mathematical Monthly, 126(6):547–552, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fraser:2019:AAP

[1410] Jonathan M. Fraser. Almost arithmetic progressions in the primes and other large sets. American Mathematical Monthly, 126(6):553–558, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Swan:2019:NES

[1411] Hunter Swan. Nonnegative eigenvectors of symmetric matrices. American Mathematical Monthly, 126(6):559–560, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gelinas:2019:IRD

[1412] Jacques Gélinas and Alexander G. Smirnov. Integral representation for derivatives of cos(x). American Mathematical Monthly, 126(6):561, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSf

[1413] Edited by Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 126(6):562– 570, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lanphier:2019:R

[1414] Dominic Lanphier. Reviews. American Mathematical Monthly, 126(6): 571–576, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Little:2019:MLT

[1415] John B. Little. The many lives of the twisted cubic. American Mathematical Monthly, 126(7):579–592, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tapp:2019:MPG

[1416] Kristopher Tapp. Measuring political gerrymandering. American Mathematical Monthly, 126(7):593–609, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Papadopoulos:2019:TTM

[1417] Athanase Papadopoulos. Three theorems of menelaus. American Mathematical Monthly, 126(7):610–619, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:2019:PRH

[1418] Anonymous. The Paul R. Halmos-Lester R. Ford Awards for 2018. American Mathematical Monthly, 126(7):619, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grinshpan:2019:LAP

[1419] Anatolii Grinshpan and Hugo J. Woerdeman. A linear-algebraic proof of Hilbert’s Ternary Quartic Theorem. American Mathematical Monthly, 126(7):620–627, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Maltenfort:2019:DSF

[1420] Michael Maltenfort. Decimal selection functions. American Mathematical Monthly, 126(7):628–634, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Younsi:2019:PCC

[1421] Malik Younsi. Peano curves in complex analysis. American Mathematical Monthly, 126(7):635–640, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jameson:2019:LMR

[1422] Graham Jameson and Peter R. Mercer. The logarithmic mean revisited. American Mathematical Monthly, 126(7):641–645, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chamberland:2019:AFP

[1423] Marc Chamberland and Eugene A. Herman. Arctangent formulas and pi. American Mathematical Monthly, 126(7):646–650, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Baoulina:2019:IRC

[1424] Ioulia N. Baoulina. Integer ratios of consecutive alternating power sums. American Mathematical Monthly, 126(7):651–654, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lonke:2019:SCV

[1425] Yossi Lonke. Shadows of cube vertices. American Mathematical Monthly, 126(7):655–656, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Aebi:2019:SPS

[1426] Christian Aebi. 53 is the smallest prime that is not the sum or difference of powers of the first two prime numbers. American Mathematical Monthly, 126(7):657, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSg

[1427] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 126(7):658–666, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Crisman:2019:R

[1428] Karl-Dieter Crisman. Reviews. American Mathematical Monthly, 126(7): 667–672, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Krusemeyer:2019:SNW

[1429] Mark Krusemeyer and Daniel H. Ullman. The Seventy-Ninth William Lowell Putnam Mathematical Competition. American Mathematical Monthly, 126(8):675–687, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pomerance:2019:PSC

[1430] Carl Pomerance and Chris Spicer. Proof of the Sheldon Conjecture. American Mathematical Monthly, 126(8):688–698, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lauritzen:2019:EAI

[1431] Niels Lauritzen and Jesper Funch Thomsen. A Euclidean algorithm for integer matrices. American Mathematical Monthly, 126(8):699, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hegselmann:2019:CFO

[1432] Rainer Hegselmann and Ulrich Krause. Consensus and fragmentation of opinions with a focus on bounded confidence. American Mathematical Monthly, 126(8):700–716, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Talvila:2019:FTI

[1433] Erik Talvila. Fourier transform inversion using an elementary differential equation and a contour integral. American Mathematical Monthly, 126 (8):717–727, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Luca:2019:PSC

[1434] Florian Luca and Pantelimon Stanica. Perfect squares as concatenation of consecutive integers. American Mathematical Monthly, 126(8): 728–734, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

East:2019:ITG

[1435] James East and Ron Niles. Integer triangles of given perimeter: a new approach via group theory. American Mathematical Monthly, 126(8): 735–739, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ellingsen:2019:FLH

[1436] Harold W. Ellingsen Jr. A fresh look at highly composite numbers. American Mathematical Monthly, 126(8):740–741, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knox:2019:GRI

[1437] Fiachra Knox, Bojan Mohar, and David R. Wood. A Golden ratio inequality for vertex degrees of graphs. American Mathematical Monthly, 126(8):742–747, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kutzman:2019:CSS

[1438] Misha Kutzman and Vadim Ponomarenko. Count to six in six languages. American Mathematical Monthly, 126(8):747, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Goyal:2019:PAF

[1439] Ashish Goyal, Hemant Kalra, and Deepak Gumber. On the probability that an automorphism fixes a group element. American Mathematical Monthly, 126(8):748–753, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

deCamargo:2019:SAP

[1440] André Pierro de Camargo. A straightforward approach to the product rule. American Mathematical Monthly, 126(8):754, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSh

[1441] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 126(8):755–763, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bremser:2019:R

[1442] Priscilla Bremser. Reviews. American Mathematical Monthly, 126(8): 764–768, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ponton:2019:CWT

[1443] Lionel Ponton. The Calkin–Wilf tree of a quadratic surd. American Mathematical Monthly, 126(9):771–785, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mabry:2019:FNI

[1444] Rick Mabry. Fibonacci numbers, integer compositions, and nets of antiprisms. American Mathematical Monthly, 126(9):786–801, 2019. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Shiu:2019:ELE

[1445] Peter Shiu. On Erdős’s last equation. American Mathematical Monthly, 126(9):802–808, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jameson:2019:OVE

[1446] G. J. O. Jameson. Operator-valued extensions of matrix-norm inequalities. American Mathematical Monthly, 126(9):809–815, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ekstrom:2019:DRL

[1447] Fredrik Ekström. Dimension of random limsup sets. American Mathematical Monthly, 126(9):816–824, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zimmerman:2019:PRL

[1448] Seth Zimmerman and Chungwu Ho. Partitioning the real line into an uncountable collection of everywhere uncountably dense sets. American Mathematical Monthly, 126(9):825–834, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Slavik:2019:UVH

[1449] Antonín Slavík. Unbiased version of Hall’s Marriage Theorem in matrix form. American Mathematical Monthly, 126(9):835–837, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mendel:2019:SPJ

[1450] Manor Mendel. A simple proof of the Johnson–Lindenstrauss extension theorem. American Mathematical Monthly, 126(9):838–840, 2019. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Trevino:2019:PPN

[1451] Enrique Treviño. Probabilistic proof that 1 + 2 + ⋯ + n = n(n + 1)/2. American Mathematical Monthly, 126(9):840, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Embacher:2019:NSH

[1452] Franz Embacher and Hans Humenberger. A note on the Stammler hyperbola. American Mathematical Monthly, 126(9):841–844, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gaber:2019:IOT

[1453] Iris Gaber, Arieh Lev, and Romina Zigdon. Insights and observations on teaching the Intermediate Value Theorem. American Mathematical Monthly, 126(9):845–849, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSi

[1454] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 126(9):850–858, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Braun:2019:BR

[1455] Benjamin Braun, Santiago de León, Kyle Franz, and Sami Sultan. Baseb representations. American Mathematical Monthly, 126(9):859, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Scoville:2019:R

[1456] Nicholas Scoville. Reviews. American Mathematical Monthly, 126(9): 860–864, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Henze:2019:AMP

[1457] Norbert Henze and Günter Last. Absent-minded passengers. American Mathematical Monthly, 126(10):867–875, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bergelson:2019:HSP

[1458] Vitaly Bergelson and Joseph Vandehey. A hot spot proof of the Generalized Wall Theorem. American Mathematical Monthly, 126(10):876–890, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bellemare:2019:PCR

[1459] Hugues Bellemare, Yves Langlois, and Thomas Ransford. A positivity conjecture related to the Riemann zeta function. American Mathematical Monthly, 126(10):891–904, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gerencser:2019:TAT

[1460] Balázs Gerencsér and Viktor Harangi. Too acute to be true: The story of acute sets. American Mathematical Monthly, 126(10):905–914, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Huber:2019:HBM

[1461] Mark Huber. Halving the bounds for the Markov, Chebyshev, and Chernoff inequalities using smoothing. American Mathematical Monthly, 126 (10):915–927, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Veklych:2019:MPL

[1462] Bogdan Veklych. A minimalist proof of the Law of Quadratic Reciprocity. American Mathematical Monthly, 126(10):928, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lavau:2019:TTR

[1463] Gérard Lavau. The truncated tetrahedron is Rupert. American Mathematical Monthly, 126(10):929–932, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Geck:2019:EPE

[1464] Meinolf Geck. Eigenvalues and polynomial equations. American Mathematical Monthly, 126(10):933–935, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ysern:2019:ASE

[1465] Bernardo de la Calle Ysern. Asymptotically stable equilibria of gradient systems. American Mathematical Monthly, 126(10):936–939, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sury:2019:LBL

[1466] B. Sury. Lower bound for the least common multiple. American Mathematical Monthly, 126(10):940–942, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Huppe:2019:PV

[1467] François Huppé and Vadim Ponomarenko. Primes in (0, n] vs. (n, m]. American Mathematical Monthly, 126(10):942, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jakhar:2019:UIR

[1468] Anuj Jakhar. A useful irreducibility result for integer polynomials. American Mathematical Monthly, 126(10):943–944, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lou:2019:RWM

[1469] Hongwei Lou. Remark on “When Must a Flat Function be Identically 0?”. American Mathematical Monthly, 126(10):945, 2019. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Edgar:2019:PSj

[1470] Gerald A. Edgar, Daniel H. Ullman, and Douglas B. West. Problems and solutions. American Mathematical Monthly, 126(10):946–954, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

King:2019:R

[1471] Dylan A. King, Thomas C. Martinez, Steven J. Miller, and Chenyang Sun. Reviews. American Mathematical Monthly, 126(10):955–958, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cohen:2019:E

[1472] Marion Deutsche Cohen. End. American Mathematical Monthly, 126 (10):958, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Colley:2019:EEb

[1473] Susan Jane Colley. Editor’s endnotes. American Mathematical Monthly, 126(10):959–960, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Anonymous:2019:MR

[1474] Anonymous. Monthly referees for 2019. American Mathematical Monthly, 126(10):i–iv, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Isaacs:2001:GCS

[1475] I. Martin Isaacs. Geometry for college students, volume 8 of Pure and applied undergraduate texts. American Mathematical Society, Providence, RI, USA, 2001. ISBN 0-8218-4794-5. 222 pp. LCCN ????

Kenschaft:2005:CPS

[1476] Patricia C. Kenschaft. Change is possible: stories of women and minorities in mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3748-6. ix + 212 pp. LCCN QA10.5 .K46 2005.

URL http://www.loc.gov/catdir/toc/fy0608/2005048105.html.

Cover:2006:EIT

[1477] Thomas M. Cover and Joy A. Thomas. Elements of Information Theory. Wiley-Interscience, New York, NY, USA, second edition, 2006. ISBN 0-471-74881-1 (e-book), 0-471-24195-4 (print), 0-471-74882-X. xxiii + 748 pp. LCCN Q360 .C68 2006. URL http://www.loc.

gov/catdir/enhancements/fy0624/2005047799-d.html;   http://
www.loc.gov/catdir/enhancements/fy0624/2005047799-t.html;
http://www.loc.gov/catdir/enhancements/fy0826/2005047799-b.
html.

Fuks:2007:MOT

[1478] Dmitry B. Fuchs and Serge Tabachnikov. Mathematical omnibus: thirty lectures on classic mathematics. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4316-8. xv + 463 pp. LCCN QA37.3 .F83 2007.

Luenberger:2008:LNP

[1479] David G. Luenberger and Yinyu Ye. Linear and Nonlinear Programming. International Series in Operations Research and Management Science Ser. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2008. ISBN 0-387-74502-5. xiii + 547 pp. LCCN QA402.5.

Bergmann:2009:JMD

[1480] Birgit Bergmann and Moritz Epple. Jüdische Mathematiker in der deutschsprachigen akademischen Kultur. (German) [Jewish Mathematicians in German-speaking academic culture]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

2009.  ISBN 3-540-69250-9, 3-540-69252-5 (e-book).  236 pp.

LCCN QA28 .J83 2009. URL http://link.springer.com/book/10.
1007/978-3-540-69252-2/page/1;   http://www.loc.gov/catdir/
enhancements/fy1109/2009378585-d.html;   http://www.loc.gov/
catdir/enhancements/fy1109/2009378585-t.html. See also English

translation [1488].

Kung:2009:CRW

[1481] Joseph P. S. Kung, Gian-Carlo Rota, and Catherine H. Yan. Combinatorics: the Rota way. Cambridge mathematical library. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-511-50473-X (e-book), 1-139-12961-9 (e-book), 0-521-88389-X (hardcover), 0-521-73794-X (paperback), 0-511-80389-3 (e-book). xii + 396 pp. LCCN QA164 .K86 2009.

URL http://www.loc.gov/catdir/toc/fy0904/2008037803.html.

Nillsen:2010:RRD

[1482] Rodney Victor Nillsen. Randomness and recurrence in dynamical systems: a real analysis approach, volume 31 of The Carus mathematical monographs. Mathematical Association of America, Washington, DC, USA, 2010. ISBN 0-88385-043-5. xviii + 357 pp. LCCN QA614.8 .N55 2010. URL http://www.loc.gov/

catdir/enhancements/fy1212/2010924993-d.html;   http://www.
loc.gov/catdir/enhancements/fy1212/2010924993-t.html.

Sandholm:2010:PGE

[1483] William H. Sandholm. Population games and evolutionary dynamics. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-19587-9. xxv + 589 pp. LCCN ????

Hersh:2011:LHM

[1484] Reuben Hersh and Vera John-Steiner. Loving + hating mathematics: challenging the myths of mathematical life. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14247-5 (hardcover), 1-4008-3611-5 (e-book). x + 416 pp. LCCN QA28 .H48 2011.

Lozano-Robledo:2011:ECM

[1485] Alvaro Lozano-Robledo. Elliptic curves, modular forms, and their Lfunctions, volume 58 of Student mathematical library; IAS/Park City mathematical subseries. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-8218-5242-6 (paperback). xiv + 195 pp. LCCN QA567.2.E44 L69 2011.

Sandholm:2011:PGE

[1486] William H. Sandholm. Population games and evolutionary dynamics. Economic learning and social evolution. MIT Press, Cambridge, MA, USA, 2011. ISBN 0-262-19587-9. xxv + 589 pp. LCCN HB144 .S27 2011.

Schwartz:2011:MS

[1487] Richard Evan Schwartz. Mostly surfaces, volume 60 of Student mathematical library. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-8218-5368-6 (paperback). xiii + 314 pp. LCCN QA571 .S385 2011.

Bergmann:2012:TTJ

[1488] Birgit Bergmann, Moritz Epple, and Ruti Ungar. Transcending tradition: Jewish mathematicians in German-speaking academic culture. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-22463-6 (hardcover), 3-642-22464-4 (e-book). 289 pp. LCCN QA28 .T73 2012. English translation by Susanne Bernhart, Staci von Boeckmann, Nicole Gentz, and Stefani Ross of [1480].

Cook:2012:PTS

[1489] William Cook. In pursuit of the traveling salesman: mathematics at the limits of computation. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15270-5. xiii + 228 pp. LCCN QA164 .C69 2012.

Gulick:2012:ECF

[1490] Denny Gulick. Encounters with chaos and fractals. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, second edition, 2012. ISBN 1-58488-517-3 (hardcover). xvi + 371 pp. LCCN Q172.5.C45 G85 2012.

Henle:2012:WNR

[1491] Michael Henle. Which numbers are real? Classroom resource materials. Mathematical Association of America, Washington, DC, USA, 2012. ISBN 0-88385-777-4 (hardcover), 1-61444-107-3 (e-book). x + 219 pp. LCCN QA255 .H46 2012. URL http://www.loc.

gov/catdir/enhancements/fy1212/2012937493-b.html;   http://
www.loc.gov/catdir/enhancements/fy1212/2012937493-d.html;
http://www.loc.gov/catdir/enhancements/fy1212/2012937493-t.
html.

Martinez:2012:CPM

[1492] Alberto A. Martinez. The cult of Pythagoras: math and myths. University of Pittsburgh Press, Pittsburgh, PA, USA, 2012. ISBN 0-8229-4418-9 (hardcover). 263 pp. LCCN QA21 .M335 2012.

Sklar:2012:MPC

[1493] Jessica K. Sklar and Elizabeth Sherr Sklar, editors. Mathematics in popular culture: essays on appearances in film, fiction, games, television and other media. McFarland, Jefferson, NC, USA, 2012. ISBN 0-7864-4978-0 (softcover). viii + 345 pp. LCCN P96.M395 M38 2012. Foreword by Keith Devlin.

Tapp:2012:SME

[1494] Kristopher Tapp. Symmetry: a mathematical exploration. Springer Science+Business Media, LLC, New York, NY, USA, 2012. ISBN 1-4614-0298-0, 1-4614-0299-9 (e-book). xiv + 215 pp. LCCN QA174.7.S96 T36 2012.

Bottazzini:2013:HHG

[1495] U. (Umberto) Bottazzini and Jeremy Gray. Hidden harmony — geometric fantasies: the rise of complex function theory. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-5724-6. xvii + 848 pp. LCCN QA331.7 .B67 2013.

Cozzens:2013:MEE

[1496] Margaret B. Cozzens and Steven J. Miller. The mathematics of encryption: an elementary introduction, volume 29 of Mathematical world. American Mathematical Society, Providence, RI, USA, 2013. ISBN 0-8218-8321-6 (paperback). xvii + 332 pp. LCCN QA268 .C697 2013.

Schneps:2013:MTH

[1497] Leila Schneps and Coralie Colmez. Math on trial: how numbers get used and abused in the courtroom. Basic Books, New York, NY, USA, 2013. ISBN 0-465-03292-3 (hardcover), 0-465-03794-1 (e-book). xi + 256 pp. LCCN K2290.S73 S36 2013.

Yanofsky:2013:OLR

[1498] Noson S. Yanofsky. The outer limits of reason: what science, mathematics, and logic cannot tell us. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-01935-3 (hardcover), 0-262-31676-5 (e-book). xiv + 403 pp. LCCN Q175.32.K45 Y36 2013.

Kot:2014:FCC

[1499] Mark Kot. A first course in the calculus of variations, volume 72 of Student mathematical library. American Mathematical Society, Providence, RI, USA, 2014. ISBN 1-4704-1495-3. x + 298 pp. LCCN QA315 .K744 2014.

Schleicher:2011:IMC

[1500] Dierk Schleicher and Malte Lackmann, editors. An Invitation to Mathematics: From Competitions to Research. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-19532-6 (soft cover), 3-642-19533-4 (e-book). xiv + 220 pp. LCCN ????